### v1

# On the Limit of Language Models as Planning Formalizers

Abstract
Large Language Models have been shown to fail to create executable and verifiable plans in grounded environments. An emerging line of work shows success in using LLM as a formalizer to generate a formal representation (e.g., PDDL) of the planning domain, which can be deterministically solved to find a plan. We systematically evaluate this methodology while bridging some major gaps. While previous work only generates a partial PDDL representation given templated and thus unrealistic environment descriptions, we generate the complete representation given descriptions of various naturalness levels. Among an array of observations critical to improve LLMs’ formal planning ability, we note that large enough models can effectively formalize descriptions as PDDL, outperforming those directly generating plans, while being robust to lexical perturbation. As the descriptions become more natural-sounding, we observe a decrease in performance and provide detailed error analysis. [^1] Our code and data can be found at https://github.com/CassieHuang22/llm-as-pddl-formalizer.
On the Limit of Language Models as Planning Formalizers
Cassie Huang  Li Zhang Drexel University {Cassie.Huang}@drexel.edu{Harry.Zhang}@drexel.edu
1 Introduction
Large language models (LLMs) can make informal plans, such as suggesting ideas for parties or giving general advice on immigration. However, most users, let alone automated agents like robots, would not be able to actually execute those plans step-by-step to fruition – either to organize parties or acquire visas – without significant prior knowledge or external help. This inability to make executable plans lies in LLMs’ inability of grounding and formal reasoning Liu et al. (2023b); Pan et al. (2023); Zhang et al. (2023). Cutting-edge research in the community has evaluated LLMs’ ability to make formal plans in grounded environments, such as textual simulations, where all objects and actions represent actualities in the real world. Therefore, any resulting plan that formally involves those objects and actions would be executable and verifiable by nature. Although formal planning has been desirable in the history of AI Weld (1999), recent work found that even state-of-the-art LLMs are unable to generate formal plans Silver et al. (2024); Valmeekam et al. (2024); Stechly et al. (2024).
Figure 1: LLM-as-formalizer uses natural language descriptions to generate the Domain and Problem File in PDDL, then these are given to a planner to find a plan. We explore the effect of natural-ness of the language in the description, by giving the model both templated and natural descriptions. Examples of Domain Descriptions and Problem Descriptions from the Blocksworld Domain are shown. The green text displays what the two examples have in common (listing all possible actions and restrictions) and the red text displays text that is not considered natural. The “Templated” text corresponds to the “Heavily Templated” version discussed in Section 4.
Instead of using the LLM as a planner to generate the plan directly, an alternative line of work uses the LLM as a formalizer. Here, the LLM generates a formal representation of a planning domain, for example in the planning domain definition language (PDDL), based on some natural language descriptions of the environment. This representation can then be fed into a solver to find the plan deterministically (see Figure 1). Previous work achieved great success by showing that LLM-as-formalizer greatly outperforms LLM-as-planner in various domains Lyu et al. (2023); Xie et al. (2023); Liu et al. (2023a); Zhang et al. (2024a); Zuo et al. (2024); Zhang et al. (2024c); Zhu et al. (2024), as LLMs are more capable of information extraction than formal reasoning Zhang et al. (2024b). However, the above work has two major shortcomings. First, to simplify the task and evaluation, most have only attempted to generate a partial PDDL representation while assuming the rest is provided, often unrealistic in real life. Second, the language used to describe the environments is often artificially templated and structured, leading to potential overestimation of models’ ability.
This paper explores the limit of LLM-as-formalizer devoid of the above two simplifications. We use LLMs to generate the entirety of a PDDL representation, including the domain file and the problem file, given a natural-sounding description of the environment and the task (see Figure 1). On one of the most widely used planning datasets, BlocksWorld PIC (1998) and its derivative, MysteryBlocksWorld Valmeekam et al. (2024), we benchmark a suite of LLMs on generating PDDL that is both solvable and correct. As the descriptions in these datasets are templated, we also construct model-generated, human-validated descriptions that are natural-sounding to different levels.
Our work is the first to systematically analyze state-of-the-art LLMs’ ability of the trending methodology of LLM-as-formalizer on the highly challenging task of formal planning. We put forward an array of observations that will benefit future efforts. Discussed in detail in Section 5, our key findings are as follows.
  • On fully-observed environments such as the BlocksWorld domain, larger GPT models can decently generate entire PDDL, while smaller open-source models cannot.
  • When feasible, LLM-as-formalizer greatly outperforms LLM-as-planner.
  • As the environment descriptions sound more human-like, the models are more challenged.
  • The performance of LLM-as-formalizer is robust to lexical perturbation, while that of LLM-as-planner is not.
  • Errors in PDDL generation span syntax and semantics in both the domain and problem.
2 Task: Formal Planning with PDDL
Formal planning (or classical planning) with PDDL involves a domain file ($\mathbb{DF}$) and problem file ($\mathbb{PF}$) (Figure 1). $\mathbb{DF}$ describes general properties in a planning domain that holds true across problems, while $\mathbb{PF}$ describes specific configurations of each problem instance. Concretely, the $\mathbb{DF}$ defines all available actions for a state-based environment as well as predicates that represent the properties of object types. Each action definition contains the name of the action, parameters and semantics. The semantics of an action include the preconditions which describe the necessary world states where the action is valid to execute, and effects which describe how the states change after the action is executed. The $\mathbb{PF}$ defines the involved objects, the initial states, and the goal states. These two files are then given to a deterministic planner which will algorithmically search for a plan. Such a plan is a series of executable, instantiated actions that sequentially transition the world states from initial to goal. In the AI community, classical planning has been deemed an effective approach to solve real-world users’ problems, as the process is precise, explainable, verifiable, and deterministic.
However, formal planning demands a well-crafted pair of $\mathbb{DF}$ and $\mathbb{PF}$. In a real-world planning scenario, an average user would not describe their environment and problem with PDDL, but more likely with a textual description of the domain ($\mathbb{DD}$) and the problem ($\mathbb{PD}$), which can be specific or loose. Hence, we focus on the textual flavor of formal planning: given $\mathbb{DD}$ and $\mathbb{PD}$, the model outputs a successful plan with regard to the $\mathbb{DF}$ and $\mathbb{PF}$ that are withheld from the model.
3 Methodology: LLM-as-Formalizer
Figure 2: Methodologies for using LLMs in Formal Planning. LLM-as-Planner generates the plan directly, while LLM-as-Formalizer translates the $\mathbb{DD}$ and $\mathbb{PD}$ into PDDL. Previous work such as Liu et al. (2023a) use the LLM to generate partial PDDL, while we generate the entire PDDL (dashed arrow).
To tackle the task above, recent work involving LLMs diverged into two methodologies. The first, LLM-as-planner, directly uses LLMs to generate a plan based on the $\mathbb{DD}$ and a $\mathbb{PD}$. The second, LLM-as-formalizer, uses LLMs to recover the $\mathbb{DF}$ and $\mathbb{PF}$, before using a deterministic planner to arrive at the plan (Figure 2). Our work will focus on the second while using the first as a baseline. LLM-as-formalizer is in essence neurosymbolic, where LLMs help define the structured representation that is otherwise prohibitively costly to annotate and brittle to generalize. Existing works in this line demonstrated success but only generated a partial PDDL representation, while assuming the rest, including $\mathbb{PF}$ goals Lyu et al. (2023); Xie et al. (2023), the $\mathbb{PF}$ Liu et al. (2023a); Zhang et al. (2024a); Zuo et al. (2024), the action semantics in the $\mathbb{DF}$ Zhang et al. (2024c); Zhu et al. (2024), and the domain file Wong et al. (2023); Guan et al. (2023). While this simplifies the task and evaluation, the assumption of provided PDDL components is often unrealistic. Therefore, we bridge this gap by asking the LLM to predict the entire PDDL – both the $\mathbb{DF}$ and $\mathbb{PF}$ [^2] It is however minimally necessary to provide the action space, the identifiers and parameters of the actions in $\mathbb{DF}$, so the agent knows what actions are possible.
4 Evaluation: Datasets, Models, Metrics
To evaluate both approaches above, we work with fully-observed textual environments. Here, the provided $\mathbb{DD}$ and $\mathbb{PF}$ contain all necessary information for the model to make a complete plan.
4.1 Data
We consider two datasets, BlocksWorld PIC (1998) and MysteryBlocksWorld Valmeekam et al. (2024), one of the most widely used domains in related work. Here, the model must rearrange stacks of blocks on a table from an initial configuration to a goal configuration using a single arm. While each instance has ground-truth $\mathbb{DF}$ and $\mathbb{PF}$, they are invisible to the models and for evaluation only. Notwithstanding existing datasets, we generate the problem configurations ourselves to maximize control over ablation studies (see Section 5). Problem configurations were generated by randomly sampling the number of blocks between 2 and 15 and number of stacks in the initial state and goal between 1 and the number of blocks. Then, the natural language description $\mathbb{DD}$s and $\mathbb{PD}$s were created using 3 methods for 3 different levels of ‘naturalness’.
Heavily Templated. The $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as MysteryBlocksworld Valmeekam et al. (2024). This description is almost a word-by-word translation of PDDL. For example, for the ‘pick-up’ action, the ground-truth PDDL $\mathbb{DF}$ would be the following:
(:action pick-up :parameters (?b - block) :precondition (and (clear ?b) (on-table ?b) (arm-empty)) :effect (and (not (on-table ?b)) (not (clear ?b)) (not (arm-empty)) (holding ?b)) )
while the Heavily Templated $\mathbb{DD}$ is:
To perform Pickup action, the following facts need to be true: clear block, block on table, arm-empty. Once Pickup action is performed the following facts will be true: holding block. Once Pickup action is performed the following facts will be false: clear block, block on table, arm-empty.
From an application point of view, spelling out all preconditions and effects in terms of the predicates is paradoxical, as it assumes the user already have the algorithmic awareness of PDDL.
Moderately Templated. The $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as the original BlocksWorld dataset, following Valmeekam et al. (2024). For example. the Moderately Templated description of the ‘pick-up’ action is:
I can only pick up or unstack one block at a time. I can only pick up or unstack a block if my hand is empty. I can only pick up a block if the block is clear. A block is clear if the block has no other blocks on top of it and if the block is not picked up.
While more natural-sounding than the Heavily Templated version, the description still explicitly discusses the preconditions and effects as well as predicates like ‘clear’.
Natural. A realistic pair of $\mathbb{DD}$ and $\mathbb{PD}$ should emulate how real-life users would describe the planning domain and problem, such that a human problem solver would understand and have just enough information to find a plan. To create such descriptions, we use a human-in-the-loop, model-assisted data generation approach.
To generate $\mathbb{DD}$, we ask GPT-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$ for BlocksWorld, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique. We next verify the naturalness of the generated text by making there were variations in language throughout all descriptions generated, but was not giving out unnecessary information.
To generate $\mathbb{PD}$, we first algorithmically and randomly permutate BlocksWorld problems, given various total numbers of blocks and stacks. We then provide the model with a symbolic configuration that contains the number blocks, the initial stack configuration and the goal stack configuration. The model then ‘humanize’ the problem by making it sound natural, given a couple of seed exemplars. We manually verify the correctness of the dataset of the non-templated problems by hand by comparing them against the problem configurations. We then verify the naturalness of the $\mathbb{PD}$ by making sure there is variation but no ambiguity in its language.
The robot arm can pick up and move one block at a time from one position to another. It is only able to move the top block from any stack or table, and have only one block held by the robot arm at a time. The main actions available are ’pick up’, …
The above example of the Natural description no longer discusses the preconditions and effects of each actions one by one, but rather focuses on the general rules to the domain. These rules apply to not only ‘pick-up’ but also other actions. Therefore, the $\mathbb{DD}$ can be much more concise, requires less algorithmic awareness, and more realistic.
In total, we construct 111 problems each varying in the number of objects in the unified domain of BlocksWorld. For each of the two Templated descriptions, there is 1 $\mathbb{DD}$ paired with each of 111 $\mathbb{PD}$s. For the Natural description, there are 111 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s. We refer to this dataset as BlocksWorld-111.
Mystery BlocksWorld is a dataset created in Valmeekam et al. (2024) that obfuscates the original BlocksWorld domain by replacing all the names of the types, predicates, actions, and objects with nonsensical words, akin to a wug test (see an example in Appendix A). This dataset as an control group can effectively test whether models create plans via lexical pattern-matching and memorization. The original Mystery BlocksWorld dataset is Heavily Templated. Since the dataset itself is an artificial perturbation, we do not provide more natural descriptions as that would defeat its purpose. We sampled 100 random problems for our experiment, resulting in 1 $\mathbb{DF}$, 1 $\mathbb{DD}$, 100 $\mathbb{PF}$s, and 100 $\mathbb{PD}$s. We refer to this dataset as MysteryBlocksWorld-100.
Data examples can be found in Appendix A.
4.2 Metrics
Following past work Guan et al. (2023); Zhu et al. (2024), the model-predicted plan is validated using VAL Howey et al. (2004) against the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$ provided above, instead of being compared against “ground-truth” plans like some work Lyu et al. (2023); Liu et al. (2023b); Pan et al. (2023) since there could be multiple correct plans. For the LLM-as-formalizer approach, the predicted $\mathbb{DF}$ and $\mathbb{PF}$ are similarly not compared against the ground-truth, as only the eventual plan is validated because there might be more than one way to formalize the planning domain and problem in PDDL.
We evaluate the predicted plans following Zuo et al. (2024): solvability and correctness. Solvability indicates the percentage of plans that were able to be found based on the predicted $\mathbb{DF}$ and $\mathbb{PF}$, regardless of whether it can correctly work with the actual environment (in our case, described by the gold $\mathbb{DF}$ and $\mathbb{PF}$). In contrast, correctness is a subset that indicates the percentage of actually correct plans. Solvability was determined using the planner dual-bfws-ffparser implemented by Muise (2016) and Correctness was evaluated using VAL333nms.kcl.ac.uk/planning/software/val.html.
4.3 Models
For both of the LLM-as-planner and LLM-as-formalizer approach, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b-it, gemma-2-27b-it Team et al. (2024), llama-3.1-8B-Instruct, llama-3.1-70B-Instruct Dubey et al. (2024), gpt-3.5-turbo-0125, gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o1-preview-2024-09-12444platform.openai.com/docs/models o1 is only reported in some experiments due to prohibitive cost.. We query these models using KANI Zhu et al. (2023) with default hyper-parameters. The open-source models are run using 4 RTX A6000 GPUs, averaging about 1062 input and output tokens for the LLM-as-formalizer approach in BlocksWorld-111. To emulate real-life application with minimal user interference, we use zero-shot prompts for all naturalness levels across all datasets (see prompts in Appendix B).
5 Results and Observations
In this section, we display our results as well as perform an in-depth analysis of the strengths and weaknesses of LLMs in formal planning, to understand the impact of the model choice, naturalness of the description, content of the task, and difficulty of the problem.
5.1 Can LLMs formalize?
Models Solvability Correctness
Llama-3.1-8B 0/111 0/111
[…]
Llama-3.1-70B 0/111 0/111
[…]
gemma-2-9b-it 3/111 3/111
gemma-2-9$\dagger$ - 10/111
gemma-2-27b-it 0/111 0/111
gemma-2-27$\dagger$ - 12/111
gpt-3.5-turbo 2/111 1/111
gpt-4o-mini 22/111 3/111
gpt-4$\dagger$ - 7/111
gpt-4o 73/111 66/111
gpt-4$\dagger$ - 37/111
o1-preview 101/111 101/111
o1$\dagger$ - 93/111
Table 1: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on the Natural BlocksWorld-111. Similarly trends hold for the Heavily and Moderately Templated versions (results are shown in Appendix C).
First, we seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, instead of partial components in previous work. Table 1 displays the results on our experiment using the most natural sounding descriptions on the BlockWorld.
These results demonstrate that big enough LLMs can decently generate PDDL on fully observed environments. We can see that for BlocksWorld, the performance for gpt-4o surpasses the performance for gpt-4o-mini, which surpasses the performance for gpt-3.5-turbo. As the model gets larger, the number of syntax errors decreases, indicating larger models’ stronger ability of code generation. This conclusion also holds for the other levels of naturalness (Figure 3).
We also observe that open-sourced models with size up to 70B almost cannot generate PDDL. Llama models cannot able to find a single plan across all three datasets while gemma models show poor though non-zero performance across all three datasets. Detailed analysis of errors will be discussed in Section 5.5.
5.2 Should LLMs formalize?
Between LLM-as-planner and LLM-as-formalizer, which is the preferred methodology for symbolic planning tasks like BlocksWorld? From Table 1, we can see that the best performing gpt-4o is able to generate solvable PDDL 73/111 times, and of those 73 plans, 66 of them are correct. This far surpasses the LLM-as-planner baseline, which only found correct plans 37/111 times. This conclusion holds for the Templated BlocksWorld-111 data (Figure 3). From Figure 3, for MysteryBlocksWorld-100, we can see that LLM-as-formalizer can generate 70/100 correct plans, which far surpassed LLM-as-planner which did not find a single correct plan as the description becomes unorthodox. For gpt-4o-mini, LLM-as-formalizer still outperforms LLM-as-planner for all cases (all naturalness level across two datasets) except Natural BlocksWorld-111.
These results demonstrate that LLMs as formalizers greatly outperforms those as planners in most cases, whenever these LLMs can foramlize PDDL at all. However, these results also show that models that might not be able to formalize (eg. Llama models) but might still be able to plan, though they will have low performance.
5.3 The more natural, the harder?
Figure 3: Performance across different naturalness level of description on BlocksWorld-111.
In this section, we discuss the question of whether using humanized descriptions makes the problem more difficult. Results from Figure 3 show that as the problem sounds more similar to PDDL, and less natural, the performance of all the models improves. Also, from our results in the Mystery BlocksWorld domain, LLM-as-formalizer performs quite well for a few of the models. However, looking at the domain and problem descriptions, they are very templated and sound like PDDL. This suggests that a more natural-sounding domain and problem description is much more challenging than templated, less natural sounding descriptions. One potential explanation is that pattern matching a template back to PDDL is much easier than having to first parse all the predicates and objects from a passage. Another reason is a more-natural sounding description may leave out implicit common-sense. For example, the Natural BlocksWorld-111 does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved. However, models often fail to invoke this knowledge and will leave out the ‘clear’ predicate, leading to unsolvable PDDL or incorrect plans.
5.4 Do LLMs memorize pretraining?
Do LLMs generate plans or formalize PDDL based on what they have memorized in their training data? We determine this by looking at the results on MysteryBlocksWorld-100, a derivative of BlocksWorld where all names are perturbed and nonsensical. From Figure 3, we can see that LLM-as-planner was not able to find a single correct plan using either gpt-4o-mini or gpt-4o. However, gpt-4o-as-formalizer surpassed this baseline with a Correctness score of 70/100. This suggests that LLM-as-formalizer is robust to lexical perturbation, and its success is not due to memorization of the domain which is a part of the pretraining data.
5.5 What kind of errors?
Natural BlocksWorld-111
Models Syntax Error $\mathbb{DF}$ Error $\mathbb{PF}$ Error
gemma-2-9b-it 15/20 20/20 20/20
gemma-2-27b-it 3/20 20/20 14/20
Llama-3.1-8B 20/20 20/20 18/20
Llama-3.1-70B 20/20 20/20 17/20
gpt-3.5-turbo 10/20 20/20 20/20
gpt-4o-mini 2/20 20/20 19/20
gpt-4o 2/20 2/20 18/20
Table 2: Different model errors in generating PDDL for all the natural BlocksWorld-111, manually annotated on a 20–example subset.
MysteryBlocksWorld-100
Models Syntax Error $\mathbb{DF}$ Error $\mathbb{PF}$ Error
gpt-3.5-turbo 16/20 19/20 2/20
gpt-4o-mini 6/20 20/20 1/20
gpt-4o 5/20 16/20 0/20
Table 3: Different errors of gpt-3.5-turbo, gpt-4o-mini and gpt-4o on Heavily Templated MysteryBlocksWorld-100, manually annotated on a 20–example subset.
Natural BlocksWorld-111
Models Wrong Precondition Wrong Effect Missing Predicate Missing Action Missing Parameters
gpt-4o-mini 11/20 18/20 19/20 1/20 2/20
gpt-4o 0/20 2/20 0/20 0/20 0/20
Heavily Templated MysteryBlocksWorld-100
Models Wrong Precondition Wrong Effect Missing Predicate Missing Action Missing Parameters
gpt-4o-mini 14/20 17/20 17/20 0/20 5/20
gpt-4o 13/20 14/20 0/20 0/20 2/20
Table 4: Analysis of errors found in $\mathbb{DF}$ for Natural BlocksWorld-111 and Heavily Templated MysteryBlocksWorld-100 out of 20 randomly sampled instances.
In this section, we discuss the kind of errors in PDDL generation. We perform an error analysis on a random 20 sample subset of problems where a plan was not found, or the found plan was not correct. From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and in the $\mathbb{PF}$. Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect action preconditions and effects, incorrect or missing predicates, and missing or incorrect action parameters. The error analysis can be found for the Natural BlocksWorld-111 in Table 2 and Table 4.
As mentioned before, as the models get larger, the amount of syntax errors in the generated PDDL decreases. Smaller models consistently make mistakes in generating the $\mathbb{DF}$. Interestingly, the most common error made for gpt-4o came from the $\mathbb{PF}$, which is intuitively easier to generate than the $\mathbb{DF}$. Common errors in the $\mathbb{PF}$ included incorrect predicates in the initial state and goal state, possibly because the descriptions are natural rather than templated, so that the implicit information in the $\mathbb{PD}$ may be nontrivial to infer. Regarding the $\mathbb{DF}$, the most common error is an incorrect effect in an action. For example, in the ‘unstack’ action, the model does not make the next block ‘clear’ when the top block has been placed in the hand.
For the open-source models, the most common error is syntax errors. For examples, models repeatedly use the keyword ‘preconditions’ instead of ‘precondition’ which might suggest a lack of grasp of the PDDL language. The smallest gemma model also made many mistakes in syntax, though fewer than the Llama models, However, despite the syntax errors, there are still many semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$, which include missing predicates and incorrect effects in the actions. For the Heavily Templated BlocksWorld-111 results, there is a significant gap between the number of plans that were found, and the number of found that were correct. Upon looking at the generated PDDL, we found that the most common error made was swapping the parameters in the preconditions of the ‘stack’ action, leading to incorrect plans.
For the MysteryBlocksWorld-100 dataset (Table 3), we see a similar trend of the syntax errors decreasing as the models get larger. For all three GPT models analyzed, there are barely any errors in the $\mathbb{PF}$ but rather the most common errors came from the $\mathbb{DF}$. Since the Mystery BlocksWorld domain is a result of lexical perturbation, the task of PDDL generation is akin to symbolic information extraction or translation, devoid of much use of commonsense knowledge. As the descriptions for MysteryBlocksWorld-100 were heavily templated, sounding the most similar to PDDL, all the predicates would be listed out in the $\mathbb{PD}$ and the model would just need to match them to PDDL syntax in the $\mathbb{PF}$. While a similar essence, this is more of a challenge for $\mathbb{DF}$ since the clauses of preconditions and effects are more involved. From Table 4, a similar trend between BlocksWorld-111 and MysteryBlocksWorld-100 also suggests that the LLM-as-formalizer methodology is robust to such perturbation.
6 Related Work
Planning with LLMs There has been a large amount of research using LLMs for planning tasks. Some use LLMs for informal planning, also known as script or procedure learning Zhang et al. (2020); Lyu et al. (2021); Lal et al. (2024). While modern LLMs can make coherent and plausible informal plans, they are ungrounded and so lack executability and verifiability. Work that use LLMs for formal planning in grounded environments generally conclude the inability of such LLMs-as-planners Silver et al. (2024); Valmeekam et al. (2024); Stechly et al. (2024). Follow-up work tackles this shortcoming by using the LLM as a heuristic, not just a planner, such as by proposing candidate plans that are iteratively verified Valmeekam et al. (2023); Kambhampati et al. (2024). While we consider the standard LLM-as-planner as a baseline, our focus is on LLM-as-formalizer, an alternative methodology for the same problem.
LLMs as PDDL Formalizer Here, LLMs do not provide plans but rather generate the a PDDL representation of the domain and problem, which is then run through a solver to find the plan. This methodology has proven successful in a number of recent works, where the LLM generates different parts but not all of the PDDL for simplified evaluation. Zuo et al. (2024); Zhang et al. (2024a); Liu et al. (2023a) use the LLM to predict the entire $\mathbb{PF}$, while Xie et al. (2023); Lyu et al. (2023) predict just the goal for the $\mathbb{PF}$. Some predict the $\mathbb{DF}$, such as Zhang et al. (2024c); Zhu et al. (2024) that generate the action semantics of the $\mathbb{DF}$ and Wong et al. (2023) who also predicts the predicates from a candidate list. Closest to our work is Guan et al. (2023) which predicts the $\mathbb{DF}$ as well as the $\mathbb{PF}$ goal. However, our work of holistically generating PDDL shows that coming up with the initial state in the $\mathbb{PF}$ is non-trivial (Section 5.5). Moreover, we vary the level of naturalness of descriptions in addition to the templated ones, which prove to be more challenging and insightful (Section 5.3).
While the above discussions pertain to LLMs generating PDDL, many work on embodied agents outside the NLP community tackle similar problems with different focus Li et al. (2024).
LLMs as Formalizer Our work hinges on modern LLMs’ ability to generate code Chen et al. (2021). In addition to writing or debugging programs Jiang et al. (2024), LLMs are also used to generate formal, interim representations that are not necessarily PDDL for problem solving. For example, Gao et al. (2023); Lyu et al. (2023); Tang et al. (2024) use the LLM to generate executable Python code for solving symbolic problems. In other work, the generated code may not be executable and is provided to another LLMs to facilitate reasoning Madaan et al. (2022); Zhang et al. (2023).
A table comparing a couple of these works can be seen in Table 7 in the Appendix (Section F).
7 Conclusion
We explore the limit of state-of-the-art LLMs to be used as a formalizer by generating the entire $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL given natural language descriptions of different naturalness levels. We find that the performance depends on a number of factors, including the size of the model and the naturalness of the descriptions. While the LLM-as-formalizer methodology greatly outperforms the LLM-as-planner methodology in domains exemplified by BlocksWorld, we conclude that with zero-shot prompting only large GPT models are currently capable for the task. Therefore, future work should attempt to equip open-source models with similar ability to democratize the ability of making executable plans. We also find that LLM-as-formalizer is robust to lexical perturbation, demonstrating strong performance in long-tail domains that are not represented in LLMs’ pretraining. Our work will hopefully drive future work on improving LLM as a planning formalizer,including experiments on partially-observed environments that require exploration and interaction, more complex environments with a larger action space, and so on.
8 Limitation
While BlocksWorld is one of the most frequently used domains in LLM’s formal planning, it is the only domain considered in this work, due to the choice to optimize for the analysis’ depth instead of breadth. As future work, we are currently considering complex domains such as Barman in IPC. Even so, a common and valid criticism for using those simulations or text problems for evaluation is that these settings may be too contrived and removed from the reality. Nevertheless, it is likely that LLMs’ satisfactory performance on these datasets is a necessary condition to success in real life.
While we only consider zero-shot prompting without any attempt for prompt tuning, it is possible that the models’ performance significantly increases otherwise. Therefore, experimental results in all settings may be underestimated. Moreover, advanced prompting techniques such as chain-of-thought, self-refine, and voting can all potentially improve model performance. However, the study of those is out of the scope of this work.
While we advocate for the LLM-as-formalizer methodology over LLM-as-planner, the former’s success may be dependent on the task. Highly symbolic tasks which can be relatively easily described, like BlocksWorld, are likely to favor LLM-as-formalizer. However, LLM-as-planner might shine in tasks with a more complex action space requiring common-sense knowledge that is easily accessed by pretraining. Furthermore, while we only consider the most straightforward LLM-as-planner prompting method, more involved methods, like Kambhampati et al. (2024) that combines LLM-as-planner with symbolic validation, will likely lead to a stronger baseline.
Since this work uses only the BlocksWorld and Mystery BlocksWorld domains, it is a small toy example to the usage of LLMs as formalizers and are not representative to problems in the real world, which would be much more challenging. This may pose a risk to users using this code on real world problems.
The datasets we use and we propose are all under the MIT License.
Appendix A Data Examples
A.1 BlocksWorld-111 PDDL
The following are an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for BlocksWorld-111.
$\mathbb{DF}$:
(define (domain blocksworld) (:predicates (clear ?x) (on-table ?x) (arm-empty) (holding ?x) (on ?x ?y)) (:action pickup :parameters (?ob) :precondition (and (clear ?ob) (on-table ?ob) (arm-empty)) :effect (and (holding ?ob) (not (clear ?ob)) (not (on-table ?ob)) (not (arm-empty)))) (:action putdown :parameters (?ob) :precondition (holding ?ob) :effect (and (clear ?ob) (arm-empty) (on-table ?ob) (not (holding ?ob)))) (:action stack :parameters (?ob ?underob) :precondition (and (clear ?underob) (holding ?ob)) :effect (and (arm-empty) (clear ?ob) (on ?ob ?underob) (not (clear ?underob)) (not (holding ?ob)))) (:action unstack :parameters (?ob ?underob) :precondition (and (on ?ob ?underob) (clear ?ob) (arm-empty)) :effect (and (holding ?ob) (clear ?underob) (not (on ?ob ?underob)) (not (clear ?ob)) (not (arm-empty)))))
The $\mathbb{DF}$ contains all four actions (pickup, putdown, stack and unstack) and their pre-conditions and post-conditions, as well as predicates needed for the domain.
$\mathbb{PF}$:
(define (problem blocksworld-p99) (:domain blocksworld) (:objects red blue green yellow ) (:init (on-table red) (on blue red) (clear blue) (on-table green) (on yellow green) (clear yellow) (arm-empty) ) (:goal (and (on-table red) (on green red) (on yellow green) (on blue yellow) )) )
The $\mathbb{PF}$ contains the objects, initial state and goal state for the problem.
A.2 BlocksWorld-111 $\mathbb{DD}$ and $\mathbb{PD}$
The following display example $\mathbb{DD}$ and $\mathbb{PD}$ for all natural settings in the BlocksWorld-111 dataset. We can see that the descriptions have all the same components as the $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL, but written in different levels of naturalness.
For the Heavily Templated $\mathbb{DD}$, all preconditions and post-conditions are written out explicitly and sound similar to PDDL. The $\mathbb{PD}$ is similar, in that it lists all the predicates needed for to solve the task.
Heavily Templated $\mathbb{DD}$:
I am playing with a set of blocks. Here are the actions I can do Pickup block Unstack block from another block Putdown block Stack block on another block I have the following restrictions on my actions: To perform Pickup action, the following facts need to be true: clear block, block on table, arm-empty. Once Pickup action is performed the following facts will be true: holding block. Once Pickup action is performed the following facts will be false: clear block, block on table, arm-empty. To perform Putdown action, the following facts need to be true: holding block. Once Putdown action is performed the following facts will be true: clear block, block on table, arm-empty. Once Putdown action is performed the following facts will be false: holding block. To perform Stack action, the following needs to be true: clear block2, holding block1. Once Stack action is performed the following will be true: arm-empty, clear block1, block1 on block2. Once Stack action is performed the following will be false: clear block2, holding block1. To perform Unstack action, the following needs to be true: block1 on block2, clear block1, arm-empty. Once Unstack action is performed the following will be true: holding block1, clear block2. Once Unstack action is performed the following will be false:, block1 on block2, clear block1, arm-empty.
Heavily Templated $\mathbb{PD}$:
As initial conditions I have that, the blue block is clear, the yellow block is clear, arm-empty, the blue block is on top of the red block, the yellow block is on top of the green block, the red block is on the table, and the green block is on the table. My goal is to have that the blue block is on top of the yellow block, the green block is on top of the red block, the yellow block is on top of the green block, and the red block is on the table.
For the Moderately Templated data, the $\mathbb{DD}$ and $\mathbb{PD}$ are much more natural than the Heavily Templated data, but all predicates are still listed.
Moderately Templated $\mathbb{DD}$:
I am playing with a set of blocks where I need to arrange the blocks into stacks. Here are the actions I can do Pick up a block Unstack a block from on top of another block Put down a block Stack a block on top of another block I have the following restrictions on my actions: I can only pick up or unstack one block at a time. I can only pick up or unstack a block if my hand is empty. I can only pick up a block if the block is on the table and the block is clear. A block is clear if the block has no other blocks on top of it and if the block is not picked up. I can only unstack a block from on top of another block if the block I am unstacking was really on top of the other block. I can only unstack a block from on top of another block if the block I am unstacking is clear. Once I pick up or unstack a block, I am holding the block. I can only put down a block that I am holding. I can only stack a block on top of another block if I am holding the block being stacked. I can only stack a block on top of another block if the block onto which I am stacking the block is clear. Once I put down or stack a block, my hand becomes empty. Once you stack a block on top of a second block, the second block is no longer clear.
Moderately Templated $\mathbb{PD}$:
As initial conditions I have that, the blue block is clear, the yellow block is clear, the hand is empty, the blue block is on top of the red block, the yellow block is on top of the green block, the red block is on the table, and the green block is on the table. My goal is to have that the blue block is on top of the yellow block, the green block is on top of the red block, the yellow block is on top of the green block, and the red block is on the table.
Finally for the natural data, we can see that the $\mathbb{DD}$ and $\mathbb{PD}$ give all necessary information to complete the task, but does not sound like PDDL, and does not describe all predicates needed to perform the task.
Natural $\mathbb{DD}$:
The Blocksworld game involves a set of blocks of different colors, which can be stacked on top of each other or placed on the table. The objective is to move the blocks from an initial configuration to a goal configuration using a series of legal moves. Legal moves in Blocksworld include: picking up a block from the table or from the top of another block, stacking a block onto the table, or stacking a block onto another block.
Natural $\mathbb{PD}$:
In this particular game, there are 4 blocks: a red block, a blue block, a green block, and a yellow block. At the start, the red block is on the table, the blue block is on top of the red block, the green block is on the table, and the yellow block is on top of the green block. The goal is to have the red block on the table, the green block on top of the red block, the yellow block on top of the green block, and the blue block on top of the yellow block.
A.3 MysteryBlocksWorld-100 PDDL
This section displays an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for MysteryBlocksWorld-100.
$\mathbb{DF}$:
(define (domain mystery_blocksworld) (:predicates (province ?x) (planet ?x) (harmony) (pain ?x) (craves ?x ?y)) (:action attack :parameters (?ob) :precondition (and (province ?ob) (planet ?ob) (harmony)) :effect (and (pain ?ob) (not (province ?ob)) (not (planet ?ob)) (not (harmony)))) (:action succumb :parameters (?ob) :precondition (pain ?ob) :effect (and (province ?ob) (harmony) (planet ?ob) (not (pain ?ob)))) (:action overcome :parameters (?ob ?underob) :precondition (and (province ?underob) (pain ?ob)) :effect (and (harmony) (province ?ob) (craves ?ob ?underob) (not (province ?underob)) (not (pain ?ob)))) (:action feast :parameters (?ob ?underob) :precondition (and (craves ?ob ?underob) (province ?ob) (harmony)) :effect (and (pain ?ob) (province ?underob) (not (craves ?ob ?underob)) (not (province ?ob)) (not (harmony)))))
$\mathbb{PF}$:
(define (problem mystery_blocksworld-p01) (:domain mystery_blocksworld) (:objects a b c d ) (:init (craves a b) (craves b c) (harmony) (planet c) (planet d) (province a) (province d) ) (:goal (and (craves a d) (craves c a) )) )
A.4 MysteryBlocksWorld-100 $\mathbb{DD}$ and $\mathbb{PD}$
The following are example $\mathbb{DD}$ and $\mathbb{PD}$ of the Heavily Templated MysteryBlocksWorld-100. Text written in green demonstrates natural sounding text while text written in red demonstrates text that sounds the most like PDDL.
$\mathbb{DD}$:
I am playing with a set of objects. Here are the actions I can do Attack object Feast object from another object Succumb object Overcome object from another object I have the following restrictions on my actions: To perform Attack action, the following facts need to be true: Province object, Planet object, Harmony. Once Attack action is performed the following facts will be true: Pain object. Once Attack action is performed the following facts will be false: Province object, Planet object, Harmony. To perform Succumb action, the following facts need to be true: Pain object. Once Succumb action is performed the following facts will be true: Province object, Planet object, Harmony. Once Succumb action is performed the following facts will be false: Pain object. To perform Overcome action, the following needs to be true: Province other object, Pain object. Once Overcome action is performed the following will be true: Harmony, Province object, Object Craves other object. Once Overcome action is performed the following will be false: Province other object, Pain object. To perform Feast action, the following needs to be true: Object Craves other object, Province object, Harmony. Once Feast action is performed the following will be true: Pain object, Province other object. Once Feast action is performed the following will be false:, Object Craves other object, Province object, Harmony.
$\mathbb{PD}$:
As initial conditions I have that, object a craves object b, object b craves object c, harmony, planet object c, planet object d, province object a and province object d. My goal is to have that object a craves object d and object c craves object a.
Appendix B Prompts
For the LLM-as-planner, we give all the models the following prompt for BlocksWorld-111:
Here is a game involving a table with blocks on it. {domain_description} {problem_description} Write the plan that would solve this problem. These are the available actions: (PICK-UP block): pick up a block from the table (PUT-DOWN block): put down a block on the table (STACK block1 block2): stack block1 onto block2 (UNSTACK block1 block2): unstack block1 from block2 Here is what the output should look like: (PICK-UP A) (STACK A B) (UNSTACK A B) (PUT-DOWN A)
For MysteryBlocksWorld-100, we use the following prompt:
Here is a game involving a table with blocks on it. {domain_description} {problem_description} Write the plan that would solve this problem. These are the available actions: (ATTACK object): attack object (SUCCUMB object): succumb (OVERCOME object1 object2): overcome object1 from object2 (FEAST object1 object2): feast object1 from object2 Here is what the output should look like: (ATTACK A) (OVERCOME A B) (FEAST A B) (SUCCUMB A)
Whenever possible, we asked the model to return the output in a JSON object for easier parsing.
Appendix C Results for Templated BlocksWorld-111
Metrics
Moderately Templated
Models Solvability Correctness
gemma-2-9b-it 0/111 0/111
gemma-2-9$\dagger$ - 5/111
gemma-2-27b-it 21/111 12/111
gemma-2-27$\dagger$ - 3/111
Llama-3.1-8B 0/111 0/111
[…]
Llama-3.1-70B 0/111 0/111
[…]
gpt-3.5-turbo 14/111 4/111
gpt-4o-mini 9/111 5/111
gpt-4$\dagger$ - 7/111
gpt-4o 87/111 76/111
gpt-4$\dagger$ - 37/111
Heavily Templated
Models Solvability Correctness
gemma-2-9b-it 71/111 10/111
gemma-2-9$\dagger$ - 5/111
gemma-2-27b-it 90/111 89/111
gemma-2-27$\dagger$ - 9/111
Llama-3.1-8B 0/111 0/111
[…]
Llama-3.1-70B 0/111 0/111
[…]
gpt-3.5-turbo 41/111 33/111
gpt-4o-mini 71/111 64/111
gpt-4$\dagger$ - 1/111
gpt-4o 99/111 99/111
gpt-4$\dagger$ - 32/111
Table 5: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on Moderately Templated and Heavily Templated BlocksWorld-111.
Table 5 displays results for Heavily Templated and Moderately Templated BlocksWorld-111.
Appendix D Results for Heavily Templated MysteryBlocksWorld-100
Table 6 displays the results for Heavily Templated MysteryBlocksWorld-100.
Models Solvability Correctness
Llama-3.1-8B 0/100 0/100
[…]
Llama-3.1-70B 0/100 0/100
[…]
gemma-2-9b-it 100/100 99/100
gemma-2-9$\dagger$ - 9/100
gemma-2-27b-it 99/100 98/100
gemma-2-27$\dagger$ - 0/100
gpt-3.5-turbo 4/100 0/100
gpt-4o-mini 36/100 5/100
gpt-4$\dagger$ - 0/100
gpt-4o 74/100 70/100
gpt-4$\dagger$ - 0/100
Table 6: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on the Heavily Templated MysteryBlocksWorld-100.
Appendix E Sample Model Output
The following is an example $\mathbb{DF}$ and $\mathbb{PF}$ that Llama-3.1-8B-Instruct gave. We can see that there are syntax errors, as well as semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$.
We give the model the following input $\mathbb{DD}$ and $\mathbb{PD}$:
Blocksworld is a puzzle domain where the aim is to move a set of blocks to achieve a specific goal configuration. The domain consists of a flat surface, often called the ’table’, and an arbitrary number of blocks. Every block can be stacked on top of another block or placed on the table. Blocks can be moved one at a time, and only the block at the top of a stack can be moved. The main actions available in Blocksworld are: 1. **Pick Up (Pick-Up B)**: Picks up block B from the table. 2. **Put Down (Put-Down B)**: Places block B onto the table. 3. **Stack (Stack B1 B2)**: Stacks block B1 on top of block B2. 4. **Unstack (Unstack B1 B2)**: Removes block B1 from the top of block B2. The goal is to transform an initial configuration of blocks into a desired target configuration using these actions. In this Blocksworld problem, we will consider a simple scenario with 2 blocks: Block1 and Block2. The initial state of the blocks is as follows: **Initial State**: - Block1 is on the table. - Block2 is on Block1. The goal is to achieve the following configuration: **Goal State**: - Block2 is on the table. - Block1 is on the table. To solve this problem, you will need to use the available actions to rearrange the blocks from the initial state to the goal state.
Which returns the following
Output $\mathbb{DF}$:
(define (domain blocksworld) (:requirements :strips :typing) (:types block) (:predicates (on-table?b - block) (on?b1 - block?b2 - block) (clear?b - block) ) (:action pick-up :parameters (?b - block) :preconditions (and (on-table?b) (clear?b)) :effects (and (not (on-table?b)) (not (clear?b)) (on-table?b)) ) (:action put-down :parameters (?b - block) :preconditions (and (not (on-table?b)) (clear?b)) :effects (and (on-table?b) (clear?b)) ) (:action stack :parameters (?b1 - block?b2 - block) :preconditions (and (on-table?b1) (clear?b2)) :effects (and (not (on-table?b1)) (not (clear?b2)) (on?b1?b2)) ) (:action unstack :parameters (?b1 - block?b2 - block) :preconditions (and (on?b1?b2) (clear?b1)) :effects (and (on-table?b1) (not (on?b1?b2)) (clear?b2)) ) )
Output $\mathbb{PF}$:
(define (problem blocksworld-problem) (:domain blocksworld) (:objects block1 block2 - block) (:init (on-table block1) (on block2 block1) (clear block1) (clear block2) ) (:goal (and (on-table block2) (on-table block1) (clear block1) (clear block2) ) ) )
Text in red displays errors outputted from the model. We can see that in the $\mathbb{DF}$ there are syntax errors (incorrect keyword “preconditions”) as well as semantic errors (incorrect predicates in preconditions and effects). For the output $\mathbb{PF}$ there are semantic errors, such as incorrect preconditions (a block cannot be clear and have another block on top of it) in the init section.
Environment LLM predicts? Natural Descriptions?
Zuo et al. (2024) fully-observed $\mathbb{PF}$ N
Zhang et al. (2024a) partially-observed $\mathbb{PF}$ N
Liu et al. (2023a) fully-observed $\mathbb{PF}$ N
Xie et al. (2023) fully-observed & partially-observed $\mathbb{PF}$ goal N
Lyu et al. (2023) fully-observed $\mathbb{PF}$ goal N
Zhang et al. (2024c) procedural texts $\mathbb{DF}$ action semantics N
Wong et al. (2023) partially-observed $\mathbb{DF}$ N
Guan et al. (2023) fully-observed $\mathbb{DF}$ $\mathbb{PF}$ goal N
Zhu et al. (2024) fully-observed $\mathbb{DF}$ action semantics N
Tang et al. (2024) partially-observed Python N/A
Silver et al. (2024) fully-observed plan N
Valmeekam et al. (2024) fully-observed plan N
Stechly et al. (2024) fully-observed plan N
This work fully-observed $\mathbb{DF}$ $\mathbb{PF}$ Y
Table 7: Comparison with related work.
Appendix F Related Works Comparison
Table 7 compares works related to this paper. We can see that other works as the LLM to predict either the plan, parts of PDDL files and other languages. We can also see that other works have mostly templated natural language descriptions, while this work uses both templated and natural descriptions.

### v2

@@
 We use LLMs to generate the entirety of a PDDL representation, including the domain file and the problem file, given a natural-sounding description of the environment and the task (see Figure 1).
-On one of the most widely used planning datasets, BlocksWorld PIC (1998) and its derivative, MysteryBlocksWorld Valmeekam et al.
-(2024), we benchmark a suite of LLMs on generating PDDL that is both solvable and correct.
+On 4 widely used planning simulations from the International Planning Competition, we benchmark a suite of LLMs on generating PDDL that is both solvable and correct.
 As the descriptions in these datasets are templated, we also construct model-generated, human-validated descriptions that are natural-sounding to different levels.
@@
 Discussed in detail in Section 5, our key findings are as follows.
-  • On fully-observed environments such as the BlocksWorld domain, larger GPT models can decently generate entire PDDL, while smaller open-source models cannot.
+  • On various planning simulations, closed-source models like GPT can decently generate entire PDDL, while open-source models like Llama up to 405B cannot.
   • When feasible, LLM-as-formalizer greatly outperforms LLM-as-planner.
@@
   • The performance of LLM-as-formalizer is robust to lexical perturbation, while that of LLM-as-planner is not.
-  • Errors in PDDL generation span syntax and semantics in both the domain and problem.
+  • Open-source models succumb to syntax errors unlike GPT models, while semantic errors are common for both types of models.
 2 Task: Formal Planning with PDDL
@@
 $\mathbb{DF}$ describes general properties in a planning domain that holds true across problems, while $\mathbb{PF}$ describes specific configurations of each problem instance.
-Concretely, the $\mathbb{DF}$ defines all available actions for a state-based environment as well as predicates that represent the properties of object types.
-Each action definition contains the name of the action, parameters and semantics.
-The semantics of an action include the preconditions which describe the necessary world states where the action is valid to execute, and effects which describe how the states change after the action is executed.
+Concretely, the $\mathbb{DF}$ defines all available actions (and their parameters and pre- and post-conditions) for a state-based environment as well as predicates that represent the properties of object types.
 The $\mathbb{PF}$ defines the involved objects, the initial states, and the goal states.
@@
 3 Methodology: LLM-as-Formalizer
-Figure 2: Methodologies for using LLMs in Formal Planning.
+Figure 2: Methodologies for using LLMs in planning.
 LLM-as-Planner generates the plan directly, while LLM-as-Formalizer translates the $\mathbb{DD}$ and $\mathbb{PD}$ into PDDL.
-Previous work such as Liu et al.
-(2023a) use the LLM to generate partial PDDL, while we generate the entire PDDL (dashed arrow).
+Previous work like Liu et al.
+(2023a) use the LLM to generate partial PDDL such as $\mathbb{PF}$ only, while we generate the entire PDDL including $\mathbb{PF}$ and $\mathbb{DF}$.
+Note the $\mathbb{DD}$ and $\mathbb{PD}$ are always provided and the $\mathbb{DF}$ and $\mathbb{PF}$ are always generated by the LLM.
 To tackle the task above, recent work involving LLMs diverged into two methodologies.
@@
 Therefore, we bridge this gap by asking the LLM to predict the entire PDDL – both the $\mathbb{DF}$ and $\mathbb{PF}$ [^2] It is however minimally necessary to provide the action space, the identifiers and parameters of the actions in $\mathbb{DF}$, so the agent knows what actions are possible.
-4 Evaluation: Datasets, Models, Metrics
+4 Evaluation: Datasets, Metrics, Models
 To evaluate both approaches above, we work with fully-observed textual environments.
 Here, the provided $\mathbb{DD}$ and $\mathbb{PF}$ contain all necessary information for the model to make a complete plan.
-4.1 Data
-We consider two datasets, BlocksWorld PIC (1998) and MysteryBlocksWorld Valmeekam et al.
-(2024), one of the most widely used domains in related work.
-Here, the model must rearrange stacks of blocks on a table from an initial configuration to a goal configuration using a single arm.
-While each instance has ground-truth $\mathbb{DF}$ and $\mathbb{PF}$, they are invisible to the models and for evaluation only.
-Notwithstanding existing datasets, we generate the problem configurations ourselves to maximize control over ablation studies (see Section 5).
-Problem configurations were generated by randomly sampling the number of blocks between 2 and 15 and number of stacks in the initial state and goal between 1 and the number of blocks.
-Then, the natural language description $\mathbb{DD}$s and $\mathbb{PD}$s were created using 3 methods for 3 different levels of ‘naturalness’.
+4.1 Datasets
+We consider four simulated planning domains, BlocksWorld, Logistics, Barman from the International Planning Competition IPC (1998), and MysteryBlocksWorld Valmeekam et al. (2024).
+BlocksWorld, also used in Liu et al. (2023a), is a domain to rearrange stacks of blocks on a table using a robotic arm. There is 1 type of entities, 5 predicates, and 4 actions.
+Mystery BlocksWorld obfuscates the original BlocksWorld domain by replacing all the names of the types, predicates, actions, and objects with nonsensical words, akin to a wug test Berko (1958).
+This dataset as an control group can effectively test whether models create plans via lexical pattern-matching and memorization.
+Logistics, also used in Guan et al.
+(2023), is a domain to transport packages across different locations using both trucks and airplanes.
+In this domain, there are 6 types of entities, 3 predicates, and 6 actions.
+Barman, also used in Zhu et al.
+(2024), is a domain to to create cocktails from ingredients using different containers and two robotic arms.
+In this domain, there are 7 types of entities, 13 predicates, and 12 actions.
+Each dataset comes with ground-truth PDDL describing domains ($\mathbb{DF}$) and problems ($\mathbb{PF}$).
+The input to the model is a natural language description of the domain ($\mathbb{DD}$) and the problem ($\mathbb{DD}$).
+The output of the model is a plan, namely a sequence of instantiated actions defined in $\mathbb{DF}$.
+For each of these datasets, the natural language description $\mathbb{DD}$s and $\mathbb{PD}$s were created in 3 different levels of naturalness.
 Heavily Templated.
-The $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as MysteryBlocksworld Valmeekam et al.
+For BlocksWorld, Logistics and Barman, the heavily templated $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as MysteryBlocksworld Valmeekam et al.
 (2024).
 This description is almost a word-by-word translation of PDDL.
-For example, for the ‘pick-up’ action, the ground-truth PDDL $\mathbb{DF}$ would be the following:
+For example, for the ‘pick-up’ action in BlocksWorld, the ground-truth PDDL $\mathbb{DF}$ would be the following:
 (:action pick-up :parameters (?b - block) :precondition (and (clear ?b) (on-table ?b) (arm-empty)) :effect (and (not (on-table ?b)) (not (clear ?b)) (not (arm-empty)) (holding ?b)) )
@@
 While more natural-sounding than the Heavily Templated version, the description still explicitly discusses the preconditions and effects as well as predicates like ‘clear’.
+Moderately Templated datasets are available only for BlocksWorld and Logistics due to the size of Barman.
 Natural.
@@
 To create such descriptions, we use a human-in-the-loop, model-assisted data generation approach.
-To generate $\mathbb{DD}$, we ask GPT-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$ for BlocksWorld, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique.
+To generate $\mathbb{DD}$, we ask GPT-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique.
 We next verify the naturalness of the generated text by making there were variations in language throughout all descriptions generated, but was not giving out unnecessary information.
-To generate $\mathbb{PD}$, we first algorithmically and randomly permutate BlocksWorld problems, given various total numbers of blocks and stacks.
-We then provide the model with a symbolic configuration that contains the number blocks, the initial stack configuration and the goal stack configuration.
+To generate $\mathbb{PD}$, we provide the model with a symbolic configuration that contains the number blocks, the initial stack configuration and the goal stack configuration.
 The model then ‘humanize’ the problem by making it sound natural, given a couple of seed exemplars.
@@
 We then verify the naturalness of the $\mathbb{PD}$ by making sure there is variation but no ambiguity in its language.
+Figure 3: Performance of both usages of LLMs on Heavily Templated BlocksWorld-100, Logistics-100, and MysteryBlocksWorld-100. Detailed results are shown in Appendix C.
 The robot arm can pick up and move one block at a time from one position to another.
@@
 Therefore, the $\mathbb{DD}$ can be much more concise, requires less algorithmic awareness, and more realistic.
-In total, we construct 111 problems each varying in the number of objects in the unified domain of BlocksWorld.
-For each of the two Templated descriptions, there is 1 $\mathbb{DD}$ paired with each of 111 $\mathbb{PD}$s.
-For the Natural description, there are 111 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s.
-We refer to this dataset as BlocksWorld-111.
-Mystery BlocksWorld is a dataset created in Valmeekam et al.
-(2024) that obfuscates the original BlocksWorld domain by replacing all the names of the types, predicates, actions, and objects with nonsensical words, akin to a wug test (see an example in Appendix A).
-This dataset as an control group can effectively test whether models create plans via lexical pattern-matching and memorization.
-The original Mystery BlocksWorld dataset is Heavily Templated.
-Since the dataset itself is an artificial perturbation, we do not provide more natural descriptions as that would defeat its purpose.
-We sampled 100 random problems for our experiment, resulting in 1 $\mathbb{DF}$, 1 $\mathbb{DD}$, 100 $\mathbb{PF}$s, and 100 $\mathbb{PD}$s.
-We refer to this dataset as MysteryBlocksWorld-100.
+In total, we construct 100 problems varying in complexity for all domains.
+For each of the two Templated descriptions, there is 1 $\mathbb{DD}$ paired with each of 100 $\mathbb{PD}$s.
+For the Natural description, there are 100 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s.
+We refer to these dataset as BlocksWorld-100, MysteryBlocksWorld-100, Logistics-100 and Barman-100.
 Data examples can be found in Appendix A.
@@
 (2024): solvability and correctness.
-Solvability indicates the percentage of plans that were able to be found based on the predicted $\mathbb{DF}$ and $\mathbb{PF}$, regardless of whether it can correctly work with the actual environment (in our case, described by the gold $\mathbb{DF}$ and $\mathbb{PF}$).
-In contrast, correctness is a subset that indicates the percentage of actually correct plans.
+Solvability only applies to LLM-as-formalizer and indicates the percentage solvable predicted $\mathbb{DF}$ and $\mathbb{PF}$, regardless of whether the resulting plan can be executed bsaed on the the gold $\mathbb{DF}$ and $\mathbb{PF}$.
+Correctness indicates the percentage of actually correct plans.
 Solvability was determined using the planner dual-bfws-ffparser implemented by Muise (2016) and Correctness was evaluated using VAL333nms.kcl.ac.uk/planning/software/val.html.
+Figure 4: Performance of LLM-as-planner (as-P) and LLM-as-formalzier (as-F) across different naturalness level of description on BlocksWorld-100 and Logistics-100.
+Detailed results are shown in Appendix C.
 4.3 Models
-For both of the LLM-as-planner and LLM-as-formalizer approach, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b-it, gemma-2-27b-it Team et al.
-(2024), llama-3.1-8B-Instruct, llama-3.1-70B-Instruct Dubey et al.
-(2024), gpt-3.5-turbo-0125, gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o1-preview-2024-09-12444platform.openai.com/docs/models o1 is only reported in some experiments due to prohibitive cost..
+For both of the LLM-as-planner and LLM-as-formalizer approach, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b|27b-it Team et al.
+(2024), llama-3.1-8B|70B|405B-Instruct Dubey et al.
+(2024), DeepSeek-R1-Distill-Llama-8B|70B Guo et al.
+(2025) [^4] Due to the cost and zero solvability on our easiest dataset, results of Llama-405B and DeepSeek-8B—70B are ommitted., gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o3-mini-2025-01-31555platform.openai.com/docs/models.
 We query these models using KANI Zhu et al.
 (2023) with default hyper-parameters.
-The open-source models are run using 4 RTX A6000 GPUs, averaging about 1062 input and output tokens for the LLM-as-formalizer approach in BlocksWorld-111.
+The open-source models are run using 4 RTX A6000 GPUs, averaging about 1062 input and output tokens for the LLM-as-formalizer approach in BlocksWorld-100.
 To emulate real-life application with minimal user interference, we use zero-shot prompts for all naturalness levels across all datasets (see prompts in Appendix B).
@@
 5.1 Can LLMs formalize?
-Models Solvability Correctness
-Llama-3.1-8B 0/111 0/111
-[…]
-Llama-3.1-70B 0/111 0/111
-[…]
-gemma-2-9b-it 3/111 3/111
-gemma-2-9$\dagger$ - 10/111
-gemma-2-27b-it 0/111 0/111
-gemma-2-27$\dagger$ - 12/111
-gpt-3.5-turbo 2/111 1/111
-gpt-4o-mini 22/111 3/111
-gpt-4$\dagger$ - 7/111
-gpt-4o 73/111 66/111
-gpt-4$\dagger$ - 37/111
-o1-preview 101/111 101/111
-o1$\dagger$ - 93/111
-Table 1: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on the Natural BlocksWorld-111.
-Similarly trends hold for the Heavily and Moderately Templated versions (results are shown in Appendix C).
-First, we seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, instead of partial components in previous work.
-Table 1 displays the results on our experiment using the most natural sounding descriptions on the BlockWorld.
-These results demonstrate that big enough LLMs can decently generate PDDL on fully observed environments.
-We can see that for BlocksWorld, the performance for gpt-4o surpasses the performance for gpt-4o-mini, which surpasses the performance for gpt-3.5-turbo.
-As the model gets larger, the number of syntax errors decreases, indicating larger models’ stronger ability of code generation.
-This conclusion also holds for the other levels of naturalness (Figure 3).
-We also observe that open-sourced models with size up to 70B almost cannot generate PDDL.
-Llama models cannot able to find a single plan across all three datasets while gemma models show poor though non-zero performance across all three datasets.
-Detailed analysis of errors will be discussed in Section 5.5.
+We seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, instead of partial components in previous work.
+Figure 3 displays the results on our experiments using the most natural sounding descriptions on Heavily Templated BlocksWorld-100, Logistics-100 and MysteryBlocksWorld-100.
+Results on the most complex domain Barman-100 is omitted due to close-to-zero performance for all models.
+These results demonstrate that GPT-family LLMs can decently generate PDDL, while open-source models even up to 405B parameters struggle.
+As formalizer, gpt-4o-mini, gpt-4o, o3-mini demonstrate non-trivial and increasing performance on BlocksWorld-100.
+On the more complex Logistics-100, gpt-4o-mini succumbs to zero performance whereas the other two show decreased performance.
+The solvability of gpt-4o-mini is often much higher than its correctness, suggesting a good grasp of the PDDL syntax but a lack of semantic understanding.
+In contrast, the solvability of gpt-4o and o3-mini is often 80% to 100% of their correctness.
+On the other hand, open-sourced models can rarely generate PDDL, a low-resource language, despite them being reportedly strong at generating high-resource languages like Python Cassano et al.
+(2022).
+All Llama models up to 405B cannot generate any solvable PDDL across all three datasets, while gemma models show poor though non-zero performance on BlocksWorld-100 and Logistics-100, and strong performance on MysteryBlocksWorld-100.
 5.2 Should LLMs formalize?
-Between LLM-as-planner and LLM-as-formalizer, which is the preferred methodology for symbolic planning tasks like BlocksWorld?
-From Table 1, we can see that the best performing gpt-4o is able to generate solvable PDDL 73/111 times, and of those 73 plans, 66 of them are correct.
-This far surpasses the LLM-as-planner baseline, which only found correct plans 37/111 times.
-This conclusion holds for the Templated BlocksWorld-111 data (Figure 3).
-From Figure 3, for MysteryBlocksWorld-100, we can see that LLM-as-formalizer can generate 70/100 correct plans, which far surpassed LLM-as-planner which did not find a single correct plan as the description becomes unorthodox.
-For gpt-4o-mini, LLM-as-formalizer still outperforms LLM-as-planner for all cases (all naturalness level across two datasets) except Natural BlocksWorld-111.
-These results demonstrate that LLMs as formalizers greatly outperforms those as planners in most cases, whenever these LLMs can foramlize PDDL at all.
-However, these results also show that models that might not be able to formalize (eg.
-Llama models) but might still be able to plan, though they will have low performance.
+Between LLM-as-planner and LLM-as-formalizer, which is the preferred methodology?
+Figure 3 shows that on BlocksWorld-100, gpt-4o is able to generate solvable PDDL 64/100 times, and of those 64 plans, 60 of them are correct.
+This far surpasses the LLM-as-planner baseline, which only found correct plans 33/100 times.
+This trend holds for Logistics-100 as well as the Moderately Templated and Natural BlocksWorld-100 data (Figure 4).
+On MysteryBlocksWorld-100, we can see that LLM-as-formalizer can generate 70/100 correct plans, which far surpassed LLM-as-planner which did not find a single correct plan as the description becomes unorthodox.
+The superiority of LLM-as-formalizer also extends to gpt-4o-mini but not o3-mini, who shows strong performance as a planner.
+These results demonstrate that LLM-as-formalizer greatly outperforms LLM-as-planner in most cases, whenever these LLMs can foramlize PDDL at all.
+However, these results also show that models that cannot formalize (e.g., Llama models) can still plan, though with close-to-zero performance.
 5.3 The more natural, the harder?
-Figure 3: Performance across different naturalness level of description on BlocksWorld-111.
-In this section, we discuss the question of whether using humanized descriptions makes the problem more difficult.
-Results from Figure 3 show that as the problem sounds more similar to PDDL, and less natural, the performance of all the models improves.
-Also, from our results in the Mystery BlocksWorld domain, LLM-as-formalizer performs quite well for a few of the models.
-However, looking at the domain and problem descriptions, they are very templated and sound like PDDL.
+We now examine whether using humanized descriptions makes the problem more difficult.
+Results from Figure 4 show that on BlocksWorld-100 as the problem sounds more similar to PDDL and less natural, the performance of all the models improves.
+Similar results hold for the Logistics domain (see results in Appendix C).
 This suggests that a more natural-sounding domain and problem description is much more challenging than templated, less natural sounding descriptions.
 One potential explanation is that pattern matching a template back to PDDL is much easier than having to first parse all the predicates and objects from a passage.
-Another reason is a more-natural sounding description may leave out implicit common-sense.
-For example, the Natural BlocksWorld-111 does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved.
+Another reason is a more natural sounding description may leave out implicit common-sense.
+For example, the Natural BlocksWorld-100 does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved.
 However, models often fail to invoke this knowledge and will leave out the ‘clear’ predicate, leading to unsolvable PDDL or incorrect plans.
@@
 5.5 What kind of errors?
-Natural BlocksWorld-111
+In this section, we discuss the kind of errors in PDDL generation.
+We perform an error analysis on a random 20 sample subset of problems where a plan was not found, or the found plan was not correct.
+From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and in the $\mathbb{PF}$.
+Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect action preconditions and effects, incorrect or missing predicates, and missing or incorrect action parameters.
+The error analysis for Natural BlocksWorld-100, Logistics-100 and Heavily Templated MysteryBlocksWorld-100 and can be found in Table 1.
+For the open-source models, the most common error is syntax errors on BlocksWorld-100 and Logistics-100.
+For example, models repeatedly use the keyword ‘preconditions’ instead of ‘precondition’ which might suggest a lack of grasp of the PDDL syntax.
+Gemma models like gemma-2-27b make significantly less syntax errors (3 out of 20 on BlocksWorld-100) than Llama models like Llama-3.1-70B (20 out of 20), despite being smaller.
+Despite the syntax errors, there are still many semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$, which include missing predicates.
+As shown in Table 3, there is a significant gap between the number of plans that were found, and the number of found that were correct.
+We find that the most common error made was swapping the parameters in the preconditions of the ‘stack’ action, leading to incorrect plans.
+While making much less syntax errors, GPT models frequently suffer from semantic errors.
+Interestingly, the most common error made for gpt-4o come from the $\mathbb{PF}$, which is intuitively easier to generate than the $\mathbb{DF}$.
+Common errors in the $\mathbb{PF}$ include incorrect predicates in the initial state and goal state.
+Common errors in the $\mathbb{DF}$ is incorrect effects in an action.
+For example, in the ‘unstack’ action, the model does not make the next block ‘clear’ when the top block has been placed in the hand.
+For MysteryBlocksWorld-100, there are barely any syntax errors or semantic errors in the $\mathbb{PF}$ but rather the most common errors come from the $\mathbb{DF}$.
+Since this domain is a result of lexical perturbation, formalizing in PDDL is akin to symbolic information extraction and translation, devoid of much use of commonsense knowledge.
+Due to the heavily templated descriptions, all the predicates would be listed out in the $\mathbb{PD}$ and the model would just need to match them to PDDL syntax in the $\mathbb{PF}$.
+While a similar essence, this is more of a challenge for $\mathbb{DF}$ since the clauses of preconditions and effects are more involved.
+From Table 2, a similar trend between BlocksWorld-100 and MysteryBlocksWorld-100 also suggests that the LLM-as-formalizer methodology is robust to such perturbation.
 Models Syntax Error $\mathbb{DF}$ Error $\mathbb{PF}$ Error
+Natural BlocksWorld-100
 gemma-2-9b-it 15/20 20/20 20/20
@@
 Llama-3.1-70B 20/20 20/20 17/20
-gpt-3.5-turbo 10/20 20/20 20/20
 gpt-4o-mini 2/20 20/20 19/20
 gpt-4o 2/20 2/20 18/20
-Table 2: Different model errors in generating PDDL for all the natural BlocksWorld-111, manually annotated on a 20–example subset.
+Natural Logistics-100
+gemma-2-9b-it 7/20 20/20 15/20
+gemma-2-27b-it 8/20 20/20 20/20
+Llama-3.1-8B 20/20 20/20 20/20
+Llama-3.1-70B 20/20 20/20 10/20
+gpt-4o-mini 2/20 20/20 19/20
+gpt-4o 5/20 20/20 19/20
 MysteryBlocksWorld-100
-Models Syntax Error $\mathbb{DF}$ Error $\mathbb{PF}$ Error
-gpt-3.5-turbo 16/20 19/20 2/20
 gpt-4o-mini 6/20 20/20 1/20
 gpt-4o 5/20 16/20 0/20
-Table 3: Different errors of gpt-3.5-turbo, gpt-4o-mini and gpt-4o on Heavily Templated MysteryBlocksWorld-100, manually annotated on a 20–example subset.
-Natural BlocksWorld-111
+Table 1: Error analysis of LLM-as-formalizer on various datasets, manually annotated on a 20–example subset.
 Models Wrong Precondition Wrong Effect Missing Predicate Missing Action Missing Parameters
+Natural BlocksWorld-100
 gpt-4o-mini 11/20 18/20 19/20 1/20 2/20
 gpt-4o 0/20 2/20 0/20 0/20 0/20
+Natural Logistics-100
+gpt-4o-mini 20/20 16/20 20/20 5/20 17/20
+gpt-4o 17/20 9/20 19/20 1/20 9/20
 Heavily Templated MysteryBlocksWorld-100
-Models Wrong Precondition Wrong Effect Missing Predicate Missing Action Missing Parameters
 gpt-4o-mini 14/20 17/20 17/20 0/20 5/20
 gpt-4o 13/20 14/20 0/20 0/20 2/20
-Table 4: Analysis of errors found in $\mathbb{DF}$ for Natural BlocksWorld-111 and Heavily Templated MysteryBlocksWorld-100 out of 20 randomly sampled instances.
-In this section, we discuss the kind of errors in PDDL generation.
-We perform an error analysis on a random 20 sample subset of problems where a plan was not found, or the found plan was not correct.
-From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and in the $\mathbb{PF}$.
-Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect action preconditions and effects, incorrect or missing predicates, and missing or incorrect action parameters.
-The error analysis can be found for the Natural BlocksWorld-111 in Table 2 and Table 4.
-As mentioned before, as the models get larger, the amount of syntax errors in the generated PDDL decreases.
-Smaller models consistently make mistakes in generating the $\mathbb{DF}$.
-Interestingly, the most common error made for gpt-4o came from the $\mathbb{PF}$, which is intuitively easier to generate than the $\mathbb{DF}$.
-Common errors in the $\mathbb{PF}$ included incorrect predicates in the initial state and goal state, possibly because the descriptions are natural rather than templated, so that the implicit information in the $\mathbb{PD}$ may be nontrivial to infer.
-Regarding the $\mathbb{DF}$, the most common error is an incorrect effect in an action.
-For example, in the ‘unstack’ action, the model does not make the next block ‘clear’ when the top block has been placed in the hand.
-For the open-source models, the most common error is syntax errors.
-For examples, models repeatedly use the keyword ‘preconditions’ instead of ‘precondition’ which might suggest a lack of grasp of the PDDL language.
-The smallest gemma model also made many mistakes in syntax, though fewer than the Llama models, However, despite the syntax errors, there are still many semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$, which include missing predicates and incorrect effects in the actions.
-For the Heavily Templated BlocksWorld-111 results, there is a significant gap between the number of plans that were found, and the number of found that were correct.
-Upon looking at the generated PDDL, we found that the most common error made was swapping the parameters in the preconditions of the ‘stack’ action, leading to incorrect plans.
-For the MysteryBlocksWorld-100 dataset (Table 3), we see a similar trend of the syntax errors decreasing as the models get larger.
-For all three GPT models analyzed, there are barely any errors in the $\mathbb{PF}$ but rather the most common errors came from the $\mathbb{DF}$.
-Since the Mystery BlocksWorld domain is a result of lexical perturbation, the task of PDDL generation is akin to symbolic information extraction or translation, devoid of much use of commonsense knowledge.
-As the descriptions for MysteryBlocksWorld-100 were heavily templated, sounding the most similar to PDDL, all the predicates would be listed out in the $\mathbb{PD}$ and the model would just need to match them to PDDL syntax in the $\mathbb{PF}$.
-While a similar essence, this is more of a challenge for $\mathbb{DF}$ since the clauses of preconditions and effects are more involved.
-From Table 4, a similar trend between BlocksWorld-111 and MysteryBlocksWorld-100 also suggests that the LLM-as-formalizer methodology is robust to such perturbation.
+Table 2: Analysis of errors found in $\mathbb{DF}$ for Natural BlocksWorld-100 and Heavily Templated MysteryBlocksWorld-100 out of 20 randomly sampled instances.
 6 Related Work
@@
 While the above discussions pertain to LLMs generating PDDL, many work on embodied agents outside the NLP community tackle similar problems with different focus Li et al. (2024).
-LLMs as Formalizer Our work hinges on modern LLMs’ ability to generate code Chen et al.
+LLM code generation Our work hinges on modern LLMs’ ability to generate code Chen et al.
 (2021).
@@
 (2023).
-A table comparing a couple of these works can be seen in Table 7 in the Appendix (Section F).
+A table comparing a couple of these works can be seen in Table 7 in the Appendix (Section E).
 7 Conclusion
-We explore the limit of state-of-the-art LLMs to be used as a formalizer by generating the entire $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL given natural language descriptions of different naturalness levels.
-We find that the performance depends on a number of factors, including the size of the model and the naturalness of the descriptions.
-While the LLM-as-formalizer methodology greatly outperforms the LLM-as-planner methodology in domains exemplified by BlocksWorld, we conclude that with zero-shot prompting only large GPT models are currently capable for the task.
+We explore the limit of state-of-the-art LLMs to be used as a PDDL formalizer for planning with natural language descriptions of different naturalness levels.
+While the LLM-as-formalizer methodology greatly outperforms the LLM-as-planner baseline in various planning domains, we conclude that with zero-shot prompting, only GPT models are sufficiently capable for the task.
 Therefore, future work should attempt to equip open-source models with similar ability to democratize the ability of making executable plans.
-We also find that LLM-as-formalizer is robust to lexical perturbation, demonstrating strong performance in long-tail domains that are not represented in LLMs’ pretraining.
-Our work will hopefully drive future work on improving LLM as a planning formalizer,including experiments on partially-observed environments that require exploration and interaction, more complex environments with a larger action space, and so on.
+We also find that LLM-as-formalizer is robust to lexical perturbation, demonstrating strong performance in long-tail domains that are underrepresented in pretraining.
+Our work will inform future efforts of using LLM as a planning formalizer, including experiments on partially-observed environments that require exploration and interaction, more complex environments with a larger action space, and so on.
 8 Limitation
-While BlocksWorld is one of the most frequently used domains in LLM’s formal planning, it is the only domain considered in this work, due to the choice to optimize for the analysis’ depth instead of breadth.
-As future work, we are currently considering complex domains such as Barman in IPC.
-Even so, a common and valid criticism for using those simulations or text problems for evaluation is that these settings may be too contrived and removed from the reality.
+A common and valid criticism for using those simulations or text problems for evaluation is that these settings may be too contrived and removed from the reality.
 Nevertheless, it is likely that LLMs’ satisfactory performance on these datasets is a necessary condition to success in real life.
@@
 (2024) that combines LLM-as-planner with symbolic validation, will likely lead to a stronger baseline.
-Since this work uses only the BlocksWorld and Mystery BlocksWorld domains, it is a small toy example to the usage of LLMs as formalizers and are not representative to problems in the real world, which would be much more challenging.
+Since this work uses only the BlocksWorld, Mystery BlocksWorld, Logistics and Barman domains, it is a small toy example to the usage of LLMs as formalizers and are not representative to problems in the real world, which would be much more challenging.
 This may pose a risk to users using this code on real world problems.
 The datasets we use and we propose are all under the MIT License.
+Acknowledgment
+We thank Peter Clark for providing invaluable input throughout this work and Andrew Zhu for providing technical support on LLM inference.
 Appendix A Data Examples
-A.1 BlocksWorld-111 PDDL
-The following are an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for BlocksWorld-111.
+As discussed above, each dataset comes with ground-truth PDDL describing domains ($\mathbb{DF}$) and problems ($\mathbb{PF}$).
+To maximize flexibility when performing analysis, we construct problem instances ourselves for some datasets, so that we can measure complexity with metrics like the number of Blocks in BlocksWorld, for which we ensure a uniform distribution to avoid biases.
+Instances for BlocksWorld were randomly generated by varying the number of blocks and number of stacks in the initial and goal states from 2 to 15.
+Instances for Mystery BlocksWorld were randomly sampled from Valmeekam et al.
+(2024).
+Instances of Logistics were taken directly from IPC (1998) and IPC (2000).
+Instances of Barman were generated by using Seipp et al.
+(2022) and varying the number of shot-glasses, ingredients and cocktails from 1 to 9.
+Below are the example PDDL and descriptions for all datasets.
+A.1 BlocksWorld-100 PDDL
+The following are an example of the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$ for BlocksWorld-100.
 $\mathbb{DF}$:
@@
 The $\mathbb{PF}$ contains the objects, initial state and goal state for the problem.
-A.2 BlocksWorld-111 $\mathbb{DD}$ and $\mathbb{PD}$
-The following display example $\mathbb{DD}$ and $\mathbb{PD}$ for all natural settings in the BlocksWorld-111 dataset.
+A.2 BlocksWorld-100 $\mathbb{DD}$ and $\mathbb{PD}$
+The following display example $\mathbb{DD}$ and $\mathbb{PD}$ for all natural settings in the BlocksWorld-100 dataset.
 We can see that the descriptions have all the same components as the $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL, but written in different levels of naturalness.
@@
 My goal is to have that object a craves object d and object c craves object a.
+A.5 Logistics-100 PDDL
+The following are an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for Logistics-100.
+$\mathbb{DF}$:
+(define (domain logistics) (:requirements :strips) (:predicates (package ?obj) (truck ?truck) (airplane ?airplane) (airport ?airport) (location ?loc) (in-city ?obj ?city) (city ?city) (at ?obj ?loc) (in ?obj ?obj)) (:action load-truck :parameters (?obj ?truck ?loc) :precondition (and (package ?obj) (truck ?truck) (location ?loc) (at ?truck ?loc) (at ?obj ?loc)) :effect (and (not (at ?obj ?loc)) (in ?obj ?truck))) (:action load-airplane :parameters (?obj ?airplane ?loc) :precondition (and (package ?obj) (airplane ?airplane) (location ?loc) (at ?obj ?loc) (at ?airplane ?loc)) :effect (and (not (at ?obj ?loc)) (in ?obj ?airplane))) (:action unload-truck :parameters (?obj ?truck ?loc) :precondition (and (package ?obj) (truck ?truck) (location ?loc) (at ?truck ?loc) (in ?obj ?truck)) :effect (and (not (in ?obj ?truck)) (at ?obj ?loc)))
+$\mathbb{DF}$ (cont’d)
+(:action unload-airplane :parameters (?obj ?airplane ?loc) :precondition (and (package ?obj) (airplane ?airplane) (location ?loc) (in ?obj ?airplane) (at ?airplane ?loc)) :effect (and (not (in ?obj ?airplane)) (at ?obj ?loc))) (:action drive-truck :parameters (?truck ?loc-from ?loc-to ?city) :precondition (and (truck ?truck) (location ?loc-from) (location ?loc-to) (city ?city) (at ?truck ?loc-from) (in-city ?loc-from ?city) (in-city ?loc-to ?city)) :effect (and (not (at ?truck ?loc-from)) (at ?truck ?loc-to))) (:action fly-airplane :parameters (?airplane ?loc-from ?loc-to) :precondition (and (airplane ?airplane) (airport ?loc-from) (airport ?loc-to) (at ?airplane ?loc-from)) :effect (and (not (at ?airplane ?loc-from)) (at ?airplane ?loc-to))) )
+The $\mathbb{DF}$ contains all six actions (load-truck, load-airplane, unload-truck, unload-airplane, drive-truck, drive-airplane) and their pre-conditions and post-conditions, as well as predicates needed for the domain.
+$\mathbb{PF}$:
+(define (problem logistics-4-0) (:domain logistics) (:objects apn1 apt2 pos2 apt1 pos1 cit2 cit1 tru2 tru1 obj23 obj22 obj21 obj13 obj12 obj11 ) (:init (package obj11) (package obj12) (package obj13) (package obj21) (package obj22) (package obj23) (truck tru1) (truck tru2) (city cit1) (city cit2) (location pos1) (location apt1) (location pos2) (location apt2) (airport apt1) (airport apt2) (airplane apn1) (at apn1 apt2) (at tru1 pos1) (at obj11 pos1) (at obj12 pos1) (at obj13 pos1) (at tru2 pos2) (at obj21 pos2) (at obj22 pos2) (at obj23 pos2) (in-city pos1 cit1) (in-city apt1 cit1) (in-city pos2 cit2) (in-city apt2 cit2)) (:goal (and (at obj11 apt1) (at obj23 pos1) (at obj13 apt1) (at obj21 pos1))) )
+The $\mathbb{PF}$ contains the objects, initial state and goal state for the problem.
+A.6 Logistics-100 $\mathbb{DD}$ and $\mathbb{PD}$
+The following display example $\mathbb{DD}$ and $\mathbb{PD}$ for all natural settings in the Logistics-100 dataset.
+We can see that the descriptions have all the same components as the $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL, but written in different levels of naturalness.
+For the Heavily Templated $\mathbb{DD}$, all preconditions and post-conditions are written out explicitly and sound similar to PDDL.
+The $\mathbb{PD}$ is similar, in that it lists all the predicates needed for to solve the task.
+Heavily Templated $\mathbb{DD}$:
+I need to move packages between locations.
+Here are the actions I can do Load an package onto a truck at a location (load-truck package truck location) Load an package onto an airplane at a location (load-airplane package airplane location) Unload an package from a truck at a location (unload-truck package truck location) Unload an package from an airplane at a location (unload-airplane package airplane location) Drive a truck from location1 to location2 in a city (drive-truck truck location1 location2 city) Fly an airplane from airport1 to airport2 (fly-airplane airplane airport1 airport2) I have the following restrictions on my actions: To perform load-truck action, the following facts need to be true: o is an package, t is a truck, l is a location, the truck is at the location, the package is at the location.
+Once load-truck action is performed the following facts will be true: the package is in the truck.
+Once load-truck action is performed the following facts will be false: the package is at the location.
+To perform load-airplane action, the following facts need to be true: o is an package, a is an airplane, l is a location, the airplane is at the location, the package is at the location.
+Once load-airplane action is performed the following facts will be true: the package is in the airplane.
+Once load-airplane action is performed the following facts will be false: the package is at the location.
+To perform unload-truck action, the following facts need to be true: o is an package, t is a truck, l is a location, the truck is at the location, the package is in the truck.
+Once unload-truck action is performed the following facts will be true: the package is at the location.
+Once unload-truck action is performed the following facts will be false: the package is in the truck.
+$\mathbb{DD}$ (cont’d)
+To perform unload-airplane action, the following facts need to be true: o is an package, a is a airplane, l is a location, the airplane is at the location, the package is in the airplane.
+Once unload-airplane action is performed the following facts will be true: the package is at the location.
+Once unload-airplane action is performed the following facts will be false: the package is in the airplane.
+To perform drive-truck action, the following need to be true: t is a truck, l1 is a location, l2 is a location, c is a city, the truck is at l1, l1 is in the city, l2 is in the city.
+Once drive-truck action is performed the following facts will be true: the truck is at l2.
+Once drive-truck action is performed the following facts will be false: the truck is at l1.
+To perform fly-airplane action, the following must be true: p is an airplane, a1 is an airport, a2 is an airport, the airplane is at a1.
+Once fly-airplane is performed the following facts will be true: the airplane is at a2.
+Once fly-airplane is performed the following facts will be false: the airplane is at a1.
+Heavily Templated $\mathbb{PD}$:
+As initial conditions, I have that, obj11 is a package, obj12 is a package, obj13 is a package, obj21 is a package, obj22 is a package, obj23 is a package, tru1 is a truck, tru2 is a truck, cit1 is a city, cit2 is a city, pos1 is a location, apt1 is a location, pos2 is a location, apt2 is a location, apt1 is an airport, apt2 is an airport, apn1 is an airplane, apn1 is at apt2, tru1 is at pos1, obj11 is at pos1, obj12 is at pos1, obj13 is at pos1, tru2 is at pos2, obj21 is at pos2, obj22 is at pos2, obj23 is at pos2, pos1 is in cit1, apt1 is in cit1, pos2 is in cit2, and apt2 is in cit2.
+My goal is to have that obj11 is at apt1, obj23 is at pos1, obj13 is at apt1, and obj21 is at pos1.
+For the Moderately Templated data, the $\mathbb{DD}$ and $\mathbb{PD}$ are much more natural than the Heavily Templated data, but all predicates are still listed.
+Moderately Templated $\mathbb{DD}$:
+I need to move packages between locations.
+Here are the actions I can do Load an package onto a truck at a location (load-truck package truck location) Load an package onto an airplane at a location (load-airplane package airplane location) Unload an package from a truck at a location (unload-truck package truck location) Unload an package from an airplane at a location (unload-airplane package airplane location) Drive a truck from location1 to location2 in a city (drive-truck truck location1 location2 city) Fly an airplane from airport1 to airport2 (fly-airplane airplane airport1 airport2) I have the following restrictions on my actions: I can only load a package onto a truck or airplane if both the package and airplane are at the location.
+Once I load the package in the truck or airplane, it is no longer at the location.
+I can only unload a package from a truck or airplane if the truck or airplane is at the location and the package is in the truck or airplane.
+Once I unload the truck or airplane, the object is at the location and no longer in the truck or airplane.
+I can only drive a truck between locations if the truck is at the first location and both the first and second locations are in the same city.
+Once I drive a truck, the truck is in the second city and no longer in the first city.
+I can only fly an airplane between two airports and the airplane is at the first airport.
+Once I fly an airplane, the airplane is at the second airport and no longer at the first airport.
+The $\mathbb{PD}$ for the Moderately Templated data is the same as the $\mathbb{PD}$ for Heavily Templated data.
+Finally for the natural data, we can see that the $\mathbb{DD}$ and $\mathbb{PD}$ give all necessary information to complete the task, but does not sound like PDDL, and does not describe all predicates needed to perform the task.
+Natural $\mathbb{DD}$:
+In the Logistics game, your goal is to transport packages between different locations using trucks and airplanes.
+Here are the actions you can perform: You can load a package onto a truck at a particular location if both the package and the truck are present there.
+Similarly, loading a package onto an airplane requires both the package and the airplane to be at the same location.
+Once a package is loaded onto a truck or airplane, it leaves its original location.
+To unload a package, the truck or airplane must be at the same location where you want to unload, and the package should be inside the vehicle.
+When you unload, the package arrives at the new location and exits the vehicle.
+If you want to drive a truck from one location to another within a city, the truck needs to begin its journey at the starting point, and both locations must be within the city boundaries.
+After driving, the truck will find itself at the destination, leaving the starting point behind.
+As for flying, an airplane can travel between two airports, but it must be ready for takeoff from the initial airport.
+After the flight, the airplane lands at the destination airport, departing from the origin airport in the process.
+Natural $\mathbb{PD}$:
+In this logistics scenario, we begin with several objects and locations.
+Package obj11, obj12, and obj13 are initially at location pos1.
+Similarly, package obj21, obj22, and obj23 start at location pos2.
+We have two trucks: truck tru1 is stationed at pos1, and truck tru2 is at pos2.
+Additionally, we have a single airplane, apn1, which is located at airport apt2.
+Our map comprises two cities: cit1 and cit2.
+City cit1 contains location pos1 and airport apt1, while city cit2 includes location pos2 and airport apt2.
+The ultimate objective is to relocate package obj11 and obj13 to airport apt1 and to move package obj21 and obj23 to location pos1.
+A.7 Barman-100 PDDL
+The following are an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for Logistics-100.
+$\mathbb{DF}$:
+(define (domain barman) (:requirements :strips :typing) (:types hand level beverage dispenser container - object ingredient cocktail - beverage shot shaker - container) (:predicates (ontable ?c - container) (holding ?h - hand ?c - container) (handempty ?h - hand) (empty ?c - container) (contains ?c - container ?b - beverage) (clean ?c - container) (used ?c - container ?b - beverage) (dispenses ?d - dispenser ?i - ingredient) (shaker-empty-level ?s - shaker ?l - level) (shaker-level ?s - shaker ?l - level) (next ?l1 ?l2 - level) (unshaked ?s - shaker) (shaked ?s - shaker) (cocktail-part1 ?c - cocktail ?i - ingredient) (cocktail-part2 ?c - cocktail ?i - ingredient)) (:action grasp :parameters (?h - hand ?c - container) :precondition (and (ontable ?c) (handempty ?h)) :effect (and (not (ontable ?c)) (not (handempty ?h)) (holding ?h ?c))) (:action leave :parameters (?h - hand ?c - container) :precondition (holding ?h ?c) :effect (and (not (holding ?h ?c)) (handempty ?h) (ontable ?c))) (:action fill-shot :parameters (?s - shot ?i - ingredient ?h1 ?h2 - hand ?d - dispenser) :precondition (and (holding ?h1 ?s) (handempty ?h2) (dispenses ?d ?i) (empty ?s) (clean ?s)) :effect (and (not (empty ?s)) (contains ?s ?i) (not (clean ?s)) (used ?s ?i)))
+$\mathbb{DF}$ (cont’d)
+(:action refill-shot :parameters (?s - shot ?i - ingredient ?h1 ?h2 - hand ?d - dispenser) :precondition (and (holding ?h1 ?s) (handempty ?h2) (dispenses ?d ?i) (empty ?s) (used ?s ?i)) :effect (and (not (empty ?s)) (contains ?s ?i))) (:action empty-shot :parameters (?h - hand ?p - shot ?b - beverage) :precondition (and (holding ?h ?p) (contains ?p ?b)) :effect (and (not (contains ?p ?b)) (empty ?p))) (:action clean-shot :parameters (?s - shot ?b - beverage ?h1 ?h2 - hand) :precondition (and (holding ?h1 ?s) (handempty ?h2) (empty ?s) (used ?s ?b)) :effect (and (not (used ?s ?b)) (clean ?s))) (:action pour-shot-to-clean-shaker :parameters (?s - shot ?i - ingredient ?d - shaker ?h1 - hand ?l ?l1 - level) :precondition (and (holding ?h1 ?s) (contains ?s ?i) (empty ?d) (clean ?d) (shaker-level ?d ?l) (next ?l ?l1)) :effect (and (not (contains ?s ?i)) (empty ?s) (contains ?d ?i) (not (empty ?d)) (not (clean ?d)) (unshaked ?d) (not (shaker-level ?d ?l)) (shaker-level ?d ?l1))) (:action pour-shot-to-used-shaker :parameters (?s - shot ?i - ingredient ?d - shaker ?h1 - hand ?l ?l1 - level) :precondition (and (holding ?h1 ?s) (contains ?s ?i) (unshaked ?d) (shaker-level ?d ?l) (next ?l ?l1)) :effect (and (not (contains ?s ?i)) (contains ?d ?i) (empty ?s) (not (shaker-level ?d ?l)) (shaker-level ?d ?l1)))
+$\mathbb{DF}$ (cont’d)
+(:action empty-shaker :parameters (?h - hand ?s - shaker ?b - cocktail ?l ?l1 - level) :precondition (and (holding ?h ?s) (contains ?s ?b) (shaked ?s) (shaker-level ?s ?l) (shaker-empty-level ?s ?l1)) :effect (and (not (shaked ?s)) (not (shaker-level ?s ?l)) (shaker-level ?s ?l1) (not (contains ?s ?b)) (empty ?s))) (:action clean-shaker :parameters (?h1 ?h2 - hand ?s - shaker) :precondition (and (holding ?h1 ?s) (handempty ?h2) (empty ?s)) :effect (and (clean ?s))) (:action shake :parameters (?b - cocktail ?d1 ?d2 - ingredient ?s - shaker ?h1 ?h2 - hand) :precondition (and (holding ?h1 ?s) (handempty ?h2) (contains ?s ?d1) (contains ?s ?d2) (cocktail-part1 ?b ?d1) (cocktail-part2 ?b ?d2) (unshaked ?s)) :effect (and (not (unshaked ?s)) (not (contains ?s ?d1)) (not (contains ?s ?d2)) (shaked ?s) (contains ?s ?b))) (:action pour-shaker-to-shot :parameters (?b - beverage ?d - shot ?h - hand ?s - shaker ?l ?l1 - level) :precondition (and (holding ?h ?s) (shaked ?s) (empty ?d) (clean ?d) (contains ?s ?b) (shaker-level ?s ?l) (next ?l1 ?l)) :effect (and (not (clean ?d)) (not (empty ?d)) (contains ?d ?b) (shaker-level ?s ?l1) (not (shaker-level ?s ?l)))) )
+The $\mathbb{DF}$ contains all twelve actions and their pre-conditions and post-conditions, as well as predicates needed for the domain.
+$\mathbb{PF}$:
+(define (problem prob) (:domain barman) (:objects shaker1 - shaker left right - hand shot1 - shot ingredient1 ingredient2 - ingredient cocktail1 - cocktail dispenser1 dispenser2 - dispenser l0 l1 l2 - level ) (:init (ontable shaker1) (ontable shot1) (dispenses dispenser1 ingredient1) (dispenses dispenser2 ingredient2) (clean shaker1) (clean shot1) (empty shaker1) (empty shot1) (handempty left) (handempty right) (shaker-empty-level shaker1 l0) (shaker-level shaker1 l0) (next l0 l1) (next l1 l2) (cocktail-part1 cocktail1 ingredient1) (cocktail-part2 cocktail1 ingredient2) ) (:goal (and (contains shot1 cocktail1) )))
+The $\mathbb{PF}$ contains the objects, initial state and goal state for the problem.
+A.8 Barman-100 $\mathbb{DD}$ and $\mathbb{PD}$
+The following display example $\mathbb{DD}$ and $\mathbb{PD}$ for the heavily templated data for Barman-100.
+We can see that the descriptions have all the same components as the $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL.
+For the Heavily Templated $\mathbb{DD}$, all preconditions and post-conditions are written out explicitly and sound similar to PDDL.
+The $\mathbb{PD}$ is similar, in that it lists all the predicates needed for to solve the task.
+Heavily Templated $\mathbb{DD}$:
+I am creating a cocktail from a set of ingredients.
+Here are the actions I can do Grasp a container (grasp hand container) Leave a container (leave hand container) Fill a shot glass with with an ingredient (fill-shot shot ingredient hand1 hand2 dispenser) Re-fill a shot glass with an ingredient (refill-shot shot ingredient hand1 hand2 dispenser) Empty a shot glass (empty-shot hand shot beverage) Clean a shot glass (clean-shot shot beverage hand hand2) Pour an ingredient from a shot glass to a clean shaker (pour-shot-to-clean-shaker shot ingredient shaker hand level level1) Pour an ingredient from a shot glass to a used shaker (pour-shot-to-used-shaker shot ingredient shaker hand level level1) Empty a shaker (empty-shaker hand shaker cocktail level level1) Clean a shaker (clean-shaker hand1 hand2 shaker) Shake a shaker (shaker cocktil ingredient1 ingredient2 shaker hand1 hand2) Pour a cocktail from a shaker to a shot glass (pour-shaker-to-shot beverage shot hand shaker level level1) I have the following restrictions on my actions: To perform Grasp action, the following facts need to be true: container on table, hand empty.
+Once Grasp action is performed the following facts will be true: hand holding container.
+Once Grasp action is performed the following facts will be false: container on table, hand empty.
+To perform Leave action, the following facts need to be true: hand holding container.
+Once Leave action is performed the following facts will be true: hand empty, container on table.
+Once Leave action is performed the following facts will be false: hand holding container.
+To perform Fill-shot action, the following needs to be true: hand1 holding shot glass, hand2 empty, dispenser dispenses ingredient, empty shot glass, clean shot glass.
+Once Fill-shot action is performed the following will be true: shot glass contains ingredient, shot glass used with ingredient.
+Once Fill-shot action is performed the following will be false: empty shot glass, clean shot glass.
+To perform Refill-shot action, the following needs to be true: hand1 holding shot glass, hand2 empty, dispenser dispenses ingredient, empty shot glass, shot glass used with ingredient.
+Once Refill-shot action is performed the following will be true: shot glass contains ingredient.
+Once Refill action is performed the following will be false: empty shot glass.
+To perform Empty-shot action, the following needs to be true: hand holding shot glass, shot glass contains beverage.
+Once Empty-shot action is performed the following will be true: empty shot glass.
+Once Empty-shot action is performed the following will be false: shot glass contains beverage.
+To perform Clean-shot action, the following needs to be true: hand1 holding shot glass, hand2 empty, empty shot glass, shot glass used with beverage.
+Once Clean-shot action is performed the following will be true: clean shot glass.
+Once Clean-shot action is performed the following will be false: shot glass used with beverage To perform Pour-shot-to-clean-shaker action, the following needs to be true: hand1 holding shot glass, shot glass contains ingredient, empty shaker, clean shaker, shaker-level of shaker is l, next level from l is l1.
+Once Pour-shot-to-clean-shaker action is performed the following will be true: empty shot glass, shaker contains ingredient, unshaked shaker, shaker-level of shaker is l1.
+Once Pour-shot-to-clean-shaker action is performed the following will be false: shot glass contains ingredient, empty shaker, clean shaker, shaker-level of shaker is l.
+To perform Pour-shot-to-used-shaker action, the following needs to be true: hand1 holding shot glass, shot glass contains ingredient, unshaked shaker, shaker-level of shaker is l, next level from l is l1.
+Once Pour-shot-to-used-shaker action is performed the following will be true: shaker contains ingredient, empty shot glass, shaker-level of shaker is l1.
+Once Pour-shot-to-used-shaker action is performed the following will be false: shot glass contains ingredient, shaker-level of shaker is l.
+$\mathbb{DD}$ (cont’d)
+To perform Empty-shaker action, the following needs to be true: hand holding shaker, shaker contains cocktail, shaked shaker, shaker-level of shaker is l, shaker-empty-level of shaker is l1.
+Once Empty-shaker action is performed the following will be true: shaker-level of shaker is l1, empty shaker.
+Once Empty-shaker action is performed the following will be false: shaked shaker, shaker-level of shaker is l, shaker contains cocktail.
+To perform Clean-shaker action, the following needs to be true: hand1 holding shaker, hand2 empty, empty shaker.
+Once Clean-shaker action is performed the following will be true: clean shaker.
+To perform Shake action, the following needs to be true: hand1 holding shaker, empty hand2, shaker contains ingredient1, shaker contains ingredient2, part 1 of cocktail is ingredient1, part 2 of cocktail is ingredient2, unshaked shaker.
+Once Shake action is performed the following will be true: shaked shaker, shaker contains cocktail.
+Once Shake action is performed the following will be false: unshaked shaker, shaker contains ingredient1, shaker contains ingredient2.
+To perform Pour-shaker-to-shot action, the following needs to be true: hand holding shaker, shaked shaker, empty shot glass, clean shot glass, shaker contains cocktail, shaker level of shaker is l, next level from l1 is l.
+Once Pour-shaker-to-shot action is performed the following will be true: shot glass contains cocktail, shot glass used with cocktail, shaker-level of shaker is l1.
+Once Pour-shaker-to-shot action is performed the following will be false: clean shot glass, empty shot glass, shaker-level of shaker is l.
+Heavily Templated $\mathbb{PD}$:
+For this cocktail, I have the following: shaker shaker1, my left hand, my right hand, shot glass shot1, ingredient ingredient1, ingredient ingredient2, dispenser dispenser1, and dispenser dispenser2.
+The shaker has the following levels: l0, l1, and l2.
+I want to make the following cocktails: cocktail1.
+As initial conditions, I have that shaker1 is on the table, shot1 is on the table, dispenser1 dispenses ingredient1, dispenser2 dispenses ingredient2, shaker1 is clean, shot1 is clean, shaker1 is empty, shot1 is empty, handempty left, handempty right, shaker-empty-level shaker1 l0, shaker-level shaker1 l0, next l0 l1, next l1 l2, cocktail-part1 cocktail1 ingredient1, and cocktail-part2 cocktail1 ingredient2.
+My goal is to have that shot1 contains cocktail1.
 Appendix B Prompts
-For the LLM-as-planner, we give all the models the following prompt for BlocksWorld-111:
+For the LLM-as-planner, we give all the models the following prompt for BlocksWorld-100:
 Here is a game involving a table with blocks on it.
@@
 Whenever possible, we asked the model to return the output in a JSON object for easier parsing.
-Appendix C Results for Templated BlocksWorld-111
+Appendix C Detailed Results
+Beyond the visualizations above, we show the detailed results of all models on all simulations of all naturalness levels.
+C.1 Results for BlocksWorld-100
 Metrics
+Natural
+Models Solvability Correctness
+gemma-2-9b-it 3/100 3/100
+gemma-2-9b-itp - 9/100
+gemma-2-27b-it 0/100 0/100
+gemma-2-27b-itp - 11/100
+Llama-3.1-8B 0/100 0/100
+Llama-3.1-8Bp - 1/100
+Llama-3.1-70B 0/100 0/100
+Llama-3.1-70Bp - 13/100
+gpt-3.5-turbo 2/100 1/100
+gpt-4o-mini 19/100 3/100
+gpt-4o-minip - 7/100
+gpt-4o 64/100 60/100
+gpt-4op - 33/100
+o1-preview 91/100 91/100
+o1-previewp - 82/100
+o3-mini 79/100 68/100
+o3-minip - 87/100
 Moderately Templated
 Models Solvability Correctness
-gemma-2-9b-it 0/111 0/111
-gemma-2-9$\dagger$ - 5/111
-gemma-2-27b-it 21/111 12/111
-gemma-2-27$\dagger$ - 3/111
-Llama-3.1-8B 0/111 0/111
-[…]
-Llama-3.1-70B 0/111 0/111
-[…]
-gpt-3.5-turbo 14/111 4/111
-gpt-4o-mini 9/111 5/111
-gpt-4$\dagger$ - 7/111
-gpt-4o 87/111 76/111
-gpt-4$\dagger$ - 37/111
+gemma-2-9b-it 0/100 0/100
+gemma-2-9b-itp - 5/100
+gemma-2-27b-it 17/100 10/100
+gemma-2-27b-itp - 3/100
+Llama-3.1-8B 0/100 0/100
+Llama-3.1-8Bp - 1/100
+Llama-3.1-70B 0/100 0/100
+Llama-3.1-70Bp - 10/100
+gpt-3.5-turbo 14/100 4/100
+gpt-4o-mini 9/100 5/100
+gpt-4o-minip - 7/100
+gpt-4o 77/100 67/100
+gpt-4op - 35/100
+o3-mini 82/100 70/100
+o3-minip - 87/100
 Heavily Templated
 Models Solvability Correctness
-gemma-2-9b-it 71/111 10/111
-gemma-2-9$\dagger$ - 5/111
-gemma-2-27b-it 90/111 89/111
-gemma-2-27$\dagger$ - 9/111
-Llama-3.1-8B 0/111 0/111
-[…]
-Llama-3.1-70B 0/111 0/111
-[…]
-gpt-3.5-turbo 41/111 33/111
-gpt-4o-mini 71/111 64/111
-gpt-4$\dagger$ - 1/111
-gpt-4o 99/111 99/111
-gpt-4$\dagger$ - 32/111
-Table 5: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on Moderately Templated and Heavily Templated BlocksWorld-111.
-Table 5 displays results for Heavily Templated and Moderately Templated BlocksWorld-111.
-Appendix D Results for Heavily Templated MysteryBlocksWorld-100
+gemma-2-9b-it 61/100 10/100
+gemma-2-9b-itp - 5/100
+gemma-2-27b-it 81/100 80/100
+gemma-2-27b-itp - 7/100
+Llama-3.1-8B 0/100 0/100
+Llama-3.1-8Bp - 0/100
+Llama-3.1-70B 0/100 0/100
+Llama-3.1-70Bp - 10/100
+gpt-3.5-turbo 39/100 29/100
+gpt-4o-mini 66/100 59/100
+gpt-4o-minip - 1/100
+gpt-4o 89/100 89/100
+gpt-4op - 29/100
+o3-mini 94/100 94/100
+o3-minip - 96/100
+Table 3: Performance of LLM-as-formalizer and LLM-as-planner (p) all BlocksWorld-100 data.
+Table 3 displays results for all results for BlocksWorld-100.
+C.2 Results for MysteryBlocksWorld-100
 Table 4 displays the results for Heavily Templated MysteryBlocksWorld-100.
@@
 Llama-3.1-8B 0/100 0/100
-[…]
+Llama-3.1-8Bp - 0/100
 Llama-3.1-70B 0/100 0/100
-[…]
+Llama-3.1-70Bp - 0/100
 gemma-2-9b-it 100/100 99/100
-gemma-2-9$\dagger$ - 9/100
+gemma-2-9b-itp - 9/100
 gemma-2-27b-it 99/100 98/100
-gemma-2-27$\dagger$ - 0/100
+gemma-2-27b-itp - 0/100
 gpt-3.5-turbo 4/100 0/100
 gpt-4o-mini 36/100 5/100
-gpt-4$\dagger$ - 0/100
+gpt-4o-minip - 0/100
 gpt-4o 74/100 70/100
-gpt-4$\dagger$ - 0/100
-Table 6: Performance of LLM-as-formalizer and LLM-as-planner ($\dagger$) on the Heavily Templated MysteryBlocksWorld-100.
-Appendix E Sample Model Output
+gpt-4op - 0/100
+o3-mini 95/100 95/100
+o3-minip - 74/100
+Table 4: Performance of LLM-as-formalizer and LLM-as-planner (p) on the Heavily Templated MysteryBlocksWorld-100.
+C.3 Results for Logistics-100
+Table 5 displays results for all naturalness settings for Logistics-100
+Metrics
+Natural
+Models Solvability Correctness
+gemma-2-9b-it 1/100 0/100
+gemma-2-9b-itp - 0/100
+gemma-2-27b-it 1/100 0/100
+gemma-2-27b-itp - 0/100
+Llama-3.1-8B 0/100 0/100
+Llama-3.1-8Bp - 1/100
+Llama-3.1-70B 0/100 0/100
+Llama-3.1-70Bp - 0/100
+gpt-3.5-turbo 1/100 0/100
+gpt-3.5-turbop - 0/100
+gpt-4o-mini 13/100 0/100
+gpt-4o-minip - 0/100
+gpt-4o 20/100 2/100
+gpt-4op - 1/100
+o3-mini 43/100 7/100
+o3-minip - 14/100
+Moderately Templated
+Models Solvability Correctness
+gemma-2-9b-it 0/100 0/100
+gemma-2-9b-itp - 0/100
+gemma-2-27b-it 3/100 0/100
+gemma-2-27b-itp - 0/100
+Llama-3.1-8B 0/100 0/100
+Llama-3.1-8Bp - 0/100
+Llama-3.1-70B 0/100 0/100
+Llama-3.1-70Bp - 0/100
+gpt-3.5-turbo 0/100 0/100
+gpt-3.5-turbop - 0/100
+gpt-4o-mini 29/100 0/100
+gpt-4o-minip - 0/100
+gpt-4o 34/100 13/100
+gpt-4op - 0/100
+o3-mini 51/100 39/100
+o3-minip - 21/100
+Heavily Templated
+Models Solvability Correctness
+gemma-2-9b-it 1/100 0/100
+gemma-2-9b-itp - 0/100
+gemma-2-27b-it 0/100 0/100
+gemma-2-27b-itp - 1/100
+Llama-3.1-8B 0/100 0/100
+Llama-3.1-8Bp - 0/100
+Llama-3.1-70B 0/100 0/100
+Llama-3.1-70Bp - 0/100
+gpt-3.5-turbo 1/100 1/100
+gpt-3.5-turbop - 0/100
+gpt-4o-mini 26/100 0/100
+gpt-4o-minip - 0/100
+gpt-4o 33/100 13/100
+gpt-4op - 0/100
+o3-mini 55/100 47/100
+o3-minip - 21/100
+Table 5: Performance of LLM-as-formalizer and LLM-as-planner (p) all Logistics-100 data.
+C.4 Results for Barman-100
+Figure 5: Performance for Barman-100.
+Table 6 displays results for Barman-100.
+Models Solvability Correctness
+Llama-3.1-8B 0/100 0/100
+Llama-3.1-8Bp - 0/100
+Llama-3.1-70B 0/100 0/100
+Llama-3.1-70Bp - 15/100
+gemma-2-9b-it 0/100 0/100
+gemma-2-9b-itp - 2/100
+gemma-2-27b-it 0/100 0/100
+gemma-2-27b-itp - 0/100
+gpt-3.5-turbo 0/100 0/100
+gpt-3.5-turbop - 4/100
+gpt-4o-mini 0/100 0/100
+gpt-4o-minip - 8/100
+gpt-4o 1/100 0/100
+gpt-4op - 3/100
+o3-mini 1/100 0/100
+o3-minip - 19/100
+Table 6: Performance of LLM-as-formalizer and LLM-as-planner (p) on the Heavily Templated Barman-100.
+Appendix D Sample Model Output
 The following is an example $\mathbb{DF}$ and $\mathbb{PF}$ that Llama-3.1-8B-Instruct gave. We can see that there are syntax errors, as well as semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$.
@@
 Table 7: Comparison with related work.
-Appendix F Related Works Comparison
+Appendix E Related Works Comparison
 Table 7 compares works related to this paper.
@@
 We can also see that other works have mostly templated natural language descriptions, while this work uses both templated and natural descriptions.
+Appendix F Methodology Comparison
+Both methodologies used in this paper incorporate LLMs into planning.
+There are pros and cons to each methodology.
+When using LLM-as-planner, we have a lightweight solution that returns results very quickly.
+However, due to the lack of reasoning skills in LLMs, they often struggle to create formal plans.
+Meanwhile using LLM-as-formalizer provides better executability and interpretability, though it uses a solver, which may results in getting results slower.
+We believe that for performance reasons that LLM-as-formalizer is the superior methodology.

### v3

@@
 Abstract
-Large Language Models have been shown to fail to create executable and verifiable plans in grounded environments.
-An emerging line of work shows success in using LLM as a formalizer to generate a formal representation (e.g., PDDL) of the planning domain, which can be deterministically solved to find a plan.
+Large Language Models have been found to create plans that are neither executable nor verifiable in grounded environments.
+An emerging line of work demonstrates success in using the LLM as a formalizer to generate a formal representation of the planning domain in some language, such as Planning Domain Definition Language (PDDL).
+This formal representation can be deterministically solved to find a plan.
 We systematically evaluate this methodology while bridging some major gaps.
-While previous work only generates a partial PDDL representation given templated and thus unrealistic environment descriptions, we generate the complete representation given descriptions of various naturalness levels.
-Among an array of observations critical to improve LLMs’ formal planning ability, we note that large enough models can effectively formalize descriptions as PDDL, outperforming those directly generating plans, while being robust to lexical perturbation.
+While previous work only generates a partial PDDL representation, given templated, and therefore unrealistic environment descriptions, we generate the complete representation given descriptions of various naturalness levels.
+Among an array of observations critical to improve LLMs’ formal planning abilities, we note that most large enough models can effectively formalize descriptions as PDDL, outperforming those directly generating plans, while being robust to lexical perturbation.
 As the descriptions become more natural-sounding, we observe a decrease in performance and provide detailed error analysis.
@@
 (2024); Valmeekam et al.
-(2024); Stechly et al.
+(2024a); Stechly et al.
 (2024).
@@
 Instead of using the LLM as a planner to generate the plan directly, an alternative line of work uses the LLM as a formalizer.
-Here, the LLM generates a formal representation of a planning domain, for example in the planning domain definition language (PDDL), based on some natural language descriptions of the environment.
+Here, the LLM generates a formal representation of a planning domain, for example in PDDL, based on some natural language descriptions of the environment.
 This representation can then be fed into a solver to find the plan deterministically (see Figure 1).
@@
 However, the above work has two major shortcomings.
-First, to simplify the task and evaluation, most have only attempted to generate a partial PDDL representation while assuming the rest is provided, often unrealistic in real life.
+First, to simplify the task and evaluation, most have only attempted to generate a partial PDDL representation while assuming the rest is provided, which is often unrealistic in real life.
 Second, the language used to describe the environments is often artificially templated and structured, leading to potential overestimation of models’ ability.
@@
 We use LLMs to generate the entirety of a PDDL representation, including the domain file and the problem file, given a natural-sounding description of the environment and the task (see Figure 1).
-On 4 widely used planning simulations from the International Planning Competition, we benchmark a suite of LLMs on generating PDDL that is both solvable and correct.
+On four widely used planning simulations from the International Planning Competition, we benchmark a suite of LLMs on generating PDDL that is both solvable and correct.
 As the descriptions in these datasets are templated, we also construct model-generated, human-validated descriptions that are natural-sounding to different levels.
-Our work is the first to systematically analyze state-of-the-art LLMs’ ability of the trending methodology of LLM-as-formalizer on the highly challenging task of formal planning.
+Our work is the first to systematically analyze state-of-the-art LLMs’ ability to use the trending methodology of LLM-as-formalizer on the highly challenging task of formal planning.
 We put forward an array of observations that will benefit future efforts.
 Discussed in detail in Section 5, our key findings are as follows.
-  • On various planning simulations, closed-source models like GPT can decently generate entire PDDL, while open-source models like Llama up to 405B cannot.
-  • When feasible, LLM-as-formalizer greatly outperforms LLM-as-planner.
-  • As the environment descriptions sound more human-like, the models are more challenged.
-  • The performance of LLM-as-formalizer is robust to lexical perturbation, while that of LLM-as-planner is not.
-  • Open-source models succumb to syntax errors unlike GPT models, while semantic errors are common for both types of models.
+  • Strongest models like gpt-4o, o3-mini, or DeepSeek-R1 can not only plan but also formalize entire PDDL, while other models like Gemma-2 or Llama-3.1 (up to 405B) cannot do either.
+  • LLM-as-formalizer often outperforms LLM-as-planner for many model-data combinations, though not always.
+  • A more human-like, natural description of the planning environment creates more challenges for models.
+  • LLM-as-formalizer is much more robust to long-tail lexical distribution than LLM-as-planner.
+  • Weaker models commonly succumb to syntax errors, while all models face semantic errors.
 2 Task: Formal Planning with PDDL
 Formal planning (or classical planning) with PDDL involves a domain file ($\mathbb{DF}$) and problem file ($\mathbb{PF}$) (Figure 1).
-$\mathbb{DF}$ describes general properties in a planning domain that holds true across problems, while $\mathbb{PF}$ describes specific configurations of each problem instance.
-Concretely, the $\mathbb{DF}$ defines all available actions (and their parameters and pre- and post-conditions) for a state-based environment as well as predicates that represent the properties of object types.
-The $\mathbb{PF}$ defines the involved objects, the initial states, and the goal states.
+The $\mathbb{DF}$ describes general properties in a planning domain that hold true across problems, while the $\mathbb{PF}$ describes specific configurations of each problem instance.
+Concretely, the $\mathbb{DF}$ defines all available actions for a state-based environment as well as predicates that represent the properties of object types.
+Each action definition contains the name of the action, parameters and semantics.
+The semantics of an action includes the preconditions which describe the necessary world states where the action is valid to execute, and effects which describe how the states change after the action is executed.
+The $\mathbb{PF}$ defines the involved objects, initial states, and goal states.
 These two files are then given to a deterministic planner which will algorithmically search for a plan.
@@
 In the AI community, classical planning has been deemed an effective approach to solve real-world users’ problems, as the process is precise, explainable, verifiable, and deterministic.
-However, formal planning demands a well-crafted pair of $\mathbb{DF}$ and $\mathbb{PF}$.
-In a real-world planning scenario, an average user would not describe their environment and problem with PDDL, but more likely with a textual description of the domain ($\mathbb{DD}$) and the problem ($\mathbb{PD}$), which can be specific or loose.
-Hence, we focus on the textual flavor of formal planning: given $\mathbb{DD}$ and $\mathbb{PD}$, the model outputs a successful plan with regard to the $\mathbb{DF}$ and $\mathbb{PF}$ that are withheld from the model.
+However, formal planning demands a well-crafted $\mathbb{DF}$ and $\mathbb{PF}$ pair.
+In a real-world planning scenario, an average user would not describe their environment and problem with PDDL, but more likely with a textual domain description ($\mathbb{DD}$) and problem description ($\mathbb{PD}$).
+While the $\mathbb{DD}$ and $\mathbb{PD}$ technically describe all the necessary information from the $\mathbb{DF}$ and $\mathbb{PF}$ in natural language, information that can be easily inferred by humans is omitted as the language in the descriptions gets more natural.
+For example, when describing the action “picking up a block”, a very natural $\mathbb{DD}$ would omit the effect that our hand is no longer empty.
+These descriptions are then translated into PDDL, which can then be used to find a plan when given to a solver.
+Hence, we focus on the textual flavor of formal planning: given a $\mathbb{DD}$ and $\mathbb{PD}$, the model outputs a successful plan with regard to the $\mathbb{DF}$ and $\mathbb{PF}$ that are withheld from the model.
 3 Methodology: LLM-as-Formalizer
@@
 To tackle the task above, recent work involving LLMs diverged into two methodologies.
-The first, LLM-as-planner, directly uses LLMs to generate a plan based on the $\mathbb{DD}$ and a $\mathbb{PD}$.
+The first, LLM-as-planner, directly uses LLMs to generate a plan based on the $\mathbb{DD}$ and $\mathbb{PD}$.
 The second, LLM-as-formalizer, uses LLMs to recover the $\mathbb{DF}$ and $\mathbb{PF}$, before using a deterministic planner to arrive at the plan (Figure 2).
 Our work will focus on the second while using the first as a baseline.
-LLM-as-formalizer is in essence neurosymbolic, where LLMs help define the structured representation that is otherwise prohibitively costly to annotate and brittle to generalize.
-Existing works in this line demonstrated success but only generated a partial PDDL representation, while assuming the rest, including $\mathbb{PF}$ goals Lyu et al.
+LLM-as-formalizer is in essence neuro-symbolic, where LLMs help define the structured representation that is otherwise prohibitively costly to annotate and brittle to generalize.
+Existing works in this line demonstrated success but only generate a partial PDDL representation, while assuming the rest.
+For example, some works generate only the $\mathbb{PF}$ goals while the $\mathbb{DF}$ and remaining $\mathbb{PF}$ are given Lyu et al.
 (2023); Xie et al.
-(2023), the $\mathbb{PF}$ Liu et al.
+(2023).
+Others generate the entire $\mathbb{PF}$ while the $\mathbb{DF}$ is given Liu et al.
 (2023a); Zhang et al.
 (2024a); Zuo et al.
-(2024), the action semantics in the $\mathbb{DF}$ Zhang et al.
+(2024).
+There have also been works that generate the action semantics in the $\mathbb{DF}$ while the remaining parts of the $\mathbb{DF}$ and entire $\mathbb{PF}$ are provided Zhang et al.
 (2024c); Zhu et al.
-(2024), and the domain file Wong et al.
+(2024), and work that generates the $\mathbb{DF}$ while the $\mathbb{PF}$ is given Wong et al.
 (2023); Guan et al.
 (2023).
-While this simplifies the task and evaluation, the assumption of provided PDDL components is often unrealistic.
+While these assumptions of given PDDL components simplify the task and evaluation, they are often unrealistic.
 Therefore, we bridge this gap by asking the LLM to predict the entire PDDL – both the $\mathbb{DF}$ and $\mathbb{PF}$ [^2] It is however minimally necessary to provide the action space, the identifiers and parameters of the actions in $\mathbb{DF}$, so the agent knows what actions are possible.
@@
 To evaluate both approaches above, we work with fully-observed textual environments.
-Here, the provided $\mathbb{DD}$ and $\mathbb{PF}$ contain all necessary information for the model to make a complete plan.
+Here, the provided $\mathbb{DD}$ and $\mathbb{PD}$ contain all necessary information for the model to make a complete plan.
 4.1 Datasets
-We consider four simulated planning domains, BlocksWorld, Logistics, Barman from the International Planning Competition IPC (1998), and MysteryBlocksWorld Valmeekam et al. (2024).
-BlocksWorld, also used in Liu et al. (2023a), is a domain to rearrange stacks of blocks on a table using a robotic arm. There is 1 type of entities, 5 predicates, and 4 actions.
+We consider four simulated planning domains, BlocksWorld, Logistics, Barman from the International Planning Competition IPC (1998), and Mystery BlocksWorld Valmeekam et al. (2024a).
+BlocksWorld, also used in Liu et al. (2023a), is a domain to rearrange stacks of blocks on a table using a robotic arm. The domain includes 1 type of entity, 5 predicates, and 4 actions.
 Mystery BlocksWorld obfuscates the original BlocksWorld domain by replacing all the names of the types, predicates, actions, and objects with nonsensical words, akin to a wug test Berko (1958).
-This dataset as an control group can effectively test whether models create plans via lexical pattern-matching and memorization.
+This dataset as a control group can effectively test whether models create plans via lexical pattern-matching and memorization.
 Logistics, also used in Guan et al.
@@
 Barman, also used in Zhu et al.
-(2024), is a domain to to create cocktails from ingredients using different containers and two robotic arms.
+(2024), is a domain to create cocktails from ingredients using different containers and two robotic arms.
 In this domain, there are 7 types of entities, 13 predicates, and 12 actions.
-Each dataset comes with ground-truth PDDL describing domains ($\mathbb{DF}$) and problems ($\mathbb{PF}$).
-The input to the model is a natural language description of the domain ($\mathbb{DD}$) and the problem ($\mathbb{DD}$).
-The output of the model is a plan, namely a sequence of instantiated actions defined in $\mathbb{DF}$.
-For each of these datasets, the natural language description $\mathbb{DD}$s and $\mathbb{PD}$s were created in 3 different levels of naturalness.
+Each dataset comes with ground-truth PDDL files describing the domain ($\mathbb{DF}$) and problem ($\mathbb{PF}$).
+The input to the model is a natural language description of the domain ($\mathbb{DD}$) and problem ($\mathbb{PD}$).
+The output of the model is a plan, namely a sequence of instantiated actions defined in the $\mathbb{DF}$.
+For each of these datasets, the natural language description ($\mathbb{DD}$s and $\mathbb{PD}$s) were created in 3 different levels of naturalness.
 Heavily Templated.
 For BlocksWorld, Logistics and Barman, the Heavily Templated $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as Mystery BlocksWorld Valmeekam et al.
-(2024).
+(2024a).
 This description is almost a word-by-word translation of PDDL.
@@
 Once Pickup action is performed the following facts will be false: clear block, block on table, arm-empty.
-From an application point of view, spelling out all preconditions and effects in terms of the predicates is paradoxical, as it assumes the user already have the algorithmic awareness of PDDL.
+From an application point of view, spelling out all preconditions and effects in terms of the predicates is paradoxical, as it assumes the user already has the algorithmic awareness of PDDL.
 Moderately Templated.
 The $\mathbb{DD}$ and $\mathbb{PD}$ are generated using the same template as the original BlocksWorld dataset, following Valmeekam et al.
-(2024).
-For example.
-the Moderately Templated description of the ‘pick-up’ action is:
+(2024a).
+For example, the Moderately Templated description of the ‘pick-up’ action is:
 I can only pick up or unstack one block at a time.
@@
 While more natural-sounding than the Heavily Templated version, the description still explicitly discusses the preconditions and effects as well as predicates like ‘clear’.
-Moderately Templated datasets are available only for BlocksWorld and Logistics due to the size of Barman.
 Natural.
@@
 To create such descriptions, we use a human-in-the-loop, model-assisted data generation approach.
-To generate $\mathbb{DD}$, we ask GPT-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique.
-We next verify the naturalness of the generated text by making there were variations in language throughout all descriptions generated, but was not giving out unnecessary information.
-To generate $\mathbb{PD}$, we provide the model with a symbolic configuration that contains the number blocks, the initial stack configuration and the goal stack configuration.
-The model then ‘humanize’ the problem by making it sound natural, given a couple of seed exemplars.
-We manually verify the correctness of the dataset of the non-templated problems by hand by comparing them against the problem configurations.
+To generate $\mathbb{DD}$s, we ask gpt-4o with high temperature to generate and paraphrase a seed annotated $\mathbb{DD}$, and then manually verify the correctness by making sure it lists the correct predicates, preconditions and effects, which are not unique.
+We next verify the naturalness of the generated text by making sure there were variations in language throughout all descriptions generated, but unnecessary information was not given out.
+To generate $\mathbb{PD}$s, we provide the model with a symbolic configuration containing the initial and goal states of the problem.
+The model then ‘humanizes’ the problem by making it sound natural, given a couple of seed exemplars.
+We manually verify the correctness of the datasets of the non-templated problems by hand by comparing them against the problem configurations.
 We then verify the naturalness of the $\mathbb{PD}$ by making sure there is variation but no ambiguity in its language.
-Figure 3: Performance of both usages of LLMs on Heavily Templated BlocksWorld-100, Logistics-100, and MysteryBlocksWorld-100. Detailed results are shown in Appendix C.
+Figure 3: Performance of both usages of LLMs on Heavily Templated BlocksWorld-100, Logistics-100, and MysteryBlocksWorld-100.
+Detailed results are shown in Appendix D.
+Due to the cost and zero solvability on our easiest dataset, results of Llama-405B and DeepSeek-8B|70B are omitted.
+In the following example of a Natural description of BlocksWorld:
 The robot arm can pick up and move one block at a time from one position to another.
@@
 The main actions available are ’pick up’, …
-The above example of the Natural description no longer discusses the preconditions and effects of each actions one by one, but rather focuses on the general rules to the domain.
+The example no longer discusses the preconditions and effects of each action one by one, but rather focuses on the general rules to the domain.
 These rules apply to not only ‘pick-up’ but also other actions.
-Therefore, the $\mathbb{DD}$ can be much more concise, requires less algorithmic awareness, and more realistic.
+Therefore, the $\mathbb{DD}$ can be much more concise, requires less algorithmic awareness, and is more realistic.
 In total, we construct 100 problems varying in complexity for all domains.
-For each of the two Templated descriptions, there is 1 $\mathbb{DD}$ paired with each of 100 $\mathbb{PD}$s.
-For the Natural description, there are 100 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s.
-We refer to these dataset as BlocksWorld-100, MysteryBlocksWorld-100, Logistics-100 and Barman-100.
+For each of the two Templated descriptions, there is one $\mathbb{DD}$ paired with each of 100 $\mathbb{PD}$s.
+For the Natural descriptions, there are 100 different pairs of $\mathbb{DD}$s and $\mathbb{PD}$s.
+We refer to these datasets as BlocksWorld-100, MysteryBlocksWorld-100, Logistics-100 and Barman-100.
+Moderately Templated and Natural datasets are available only for BlocksWorld-100 and Logistics-100 due to the size of Barman-100.
 Data examples can be found in Appendix A.
@@
 (2023); Zhu et al.
-(2024), the model-predicted plan is validated using VAL Howey et al.
-(2004) against the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$ provided above, instead of being compared against “ground-truth” plans like some work Lyu et al.
+(2024), the plan produced from LLM-as-Planner is validated using VAL Howey et al.
+(2004) against the ground-truth $\mathbb{DF}$s and $\mathbb{PF}$s provided above.
+This is done instead of being compared against “ground-truth” plans like some work Lyu et al.
 (2023); Liu et al.
@@
 (2023) since there could be multiple correct plans.
-For the LLM-as-formalizer approach, the predicted $\mathbb{DF}$ and $\mathbb{PF}$ are similarly not compared against the ground-truth, as only the eventual plan is validated because there might be more than one way to formalize the planning domain and problem in PDDL.
+Similarly, the predicted $\mathbb{DF}$ and $\mathbb{PF}$ for the LLM-as-formalizer approach are not compared against the ground-truth, as only the eventual plan is validated because there might be more than one way to formalize the planning domain and problem in PDDL.
 We evaluate the predicted plans following Zuo et al.
-(2024): solvability and correctness.
-Solvability only applies to LLM-as-formalizer and indicates the percentage solvable predicted $\mathbb{DF}$ and $\mathbb{PF}$, regardless of whether the resulting plan can be executed bsaed on the the gold $\mathbb{DF}$ and $\mathbb{PF}$.
+(2024): using solvability and correctness.
+Solvability only applies to LLM-as-formalizer and indicates the percentage of solvable predicted $\mathbb{DF}$s and $\mathbb{PF}$s, regardless of whether the resulting plan can be executed based on the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$.
 Correctness indicates the percentage of actually correct plans.
 Solvability was determined using the planner dual-bfws-ffparser implemented by Muise (2016) and Correctness was evaluated using VAL333nms.kcl.ac.uk/planning/software/val.html.
-Figure 4: Performance of LLM-as-planner (as-P) and LLM-as-formalzier (as-F) across different naturalness level of description on BlocksWorld-100 and Logistics-100.
-Detailed results are shown in Appendix C.
+More information about the solver and validator can be found in the Appendix, Section C.
 4.3 Models
-For both of the LLM-as-planner and LLM-as-formalizer approach, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b|27b-it Team et al.
+For both of the LLM-as-planner and LLM-as-formalizer approaches, we consider a number of models, including open-source and closed-source LLMs varying in size, including gemma-2-9b|27b-it Team et al.
 (2024), llama-3.1-8B|70B|405B-Instruct Dubey et al.
-(2024), DeepSeek-R1-Distill-Llama-8B|70B Guo et al.
-(2025) [^4] Due to the cost and zero solvability on our easiest dataset, results of Llama-405B and DeepSeek-8B—70B are ommitted., gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o3-mini-2025-01-31555platform.openai.com/docs/models.
+(2024), DeepSeek-R1-Distill-Llama-8B|70B|671B Guo et al.
+(2025), gpt-4o-mini-2024-07-18, gpt-4o-2024-08-06, and o3-mini-2025-01-31444platform.openai.com/docs/models.
 We query these models using KANI Zhu et al.
@@
 To emulate real-life application with minimal user interference, we use zero-shot prompts for all naturalness levels across all datasets (see prompts in Appendix B).
+We would also like to note that since the LLM has to generate the entire $\mathbb{DF}$ and $\mathbb{PF}$, using few-shot prompting would not be meaningful.
+Each dataset corresponds to one unique domain and we cannot provide the LLM the $\mathbb{DF}$ in the prompt as the model is supposed to predict it.
+As for the $\mathbb{PF}$, we also cannot provide it to the model because the typing, predicates, and usage in the $\mathbb{PF}$ would also give away critical information from the file.
+Nevertheless, we performed a small set of experiments using different prompting methods, such as few-shot and chain-of-thought prompting on Heavily Templated and Natural BlocksWorld-100.
+Results and discussion of these experiments can be found in the Appendix Sections D.5 and D.6.
 5 Results and Observations
@@
 5.1 Can LLMs formalize?
-We seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, instead of partial components in previous work.
-Figure 3 displays the results on our experiments using the most natural sounding descriptions on Heavily Templated BlocksWorld-100, Logistics-100 and MysteryBlocksWorld-100.
+We seek to understand the extent to which LLMs can act as a formalizer to generate entire PDDL, rather than partial components seen in previous work.
+Figure 3 displays the results of our experiments on Heavily Templated BlocksWorld-100, Logistics-100 and MysteryBlocksWorld-100.
 Results on the most complex domain Barman-100 is omitted due to close-to-zero performance for all models.
-These results demonstrate that GPT-family LLMs can decently generate PDDL, while open-source models even up to 405B parameters struggle.
-As formalizer, gpt-4o-mini, gpt-4o, o3-mini demonstrate non-trivial and increasing performance on BlocksWorld-100.
-On the more complex Logistics-100, gpt-4o-mini succumbs to zero performance whereas the other two show decreased performance.
+These results demonstrate that GPT-family LLMs and DeepSeek-R1 can decently generate PDDL, while open-source models like Gemma and Llama (even up to 405B parameters) struggle.
+As a formalizer, gpt-4o-mini, gpt-4o and o3-mini demonstrate non-trivial and increasing performance on BlocksWorld-100.
+On the more complex Logistics-100, gpt-4o-mini succumbs to zero performance, whereas the other two show decreased performance.
 The solvability of gpt-4o-mini is often much higher than its correctness, suggesting a good grasp of the PDDL syntax but a lack of semantic understanding.
@@
 (2022).
-All Llama models up to 405B cannot generate any solvable PDDL across all three datasets, while gemma models show poor though non-zero performance on BlocksWorld-100 and Logistics-100, and strong performance on MysteryBlocksWorld-100.
+All Llama models up to 405B cannot generate any solvable PDDL across all three datasets, while Gemma models show poor, though non-zero, performance on BlocksWorld-100 and Logistics-100, and strong performance on MysteryBlocksWorld-100.
 5.2 Should LLMs formalize?
 Between LLM-as-planner and LLM-as-formalizer, which is the preferred methodology?
-Figure 3 shows that on BlocksWorld-100, gpt-4o is able to generate solvable PDDL 64/100 times, and of those 64 plans, 60 of them are correct.
+Figure 3 shows that on BlocksWorld-100, gpt-4o is able to generate solvable PDDL 64 times, out of 100, and of those 64 plans, 60 of them are correct.
 This far surpasses the LLM-as-planner baseline, which only found correct plans 33/100 times.
@@
 The superiority of LLM-as-formalizer also extends to gpt-4o-mini but not o3-mini, who shows strong performance as a planner.
-These results demonstrate that LLM-as-formalizer greatly outperforms LLM-as-planner in most cases, whenever these LLMs can foramlize PDDL at all.
+These results demonstrate that LLM-as-formalizer greatly outperforms LLM-as-planner in most cases, whenever these LLMs can formalize PDDL at all.
 However, these results also show that models that cannot formalize (e.g., Llama models) can still plan, though with close-to-zero performance.
 5.3 The more natural, the harder?
+Figure 4: Performance of LLM-as-planner (as-P) and LLM-as-formalizer (as-F) across different naturalness level of description on BlocksWorld-100 and Logistics-100.
+Detailed results are shown in Appendix D.
 We now examine whether using humanized descriptions makes the problem more difficult.
 Results from Figure 4 show that on BlocksWorld-100 as the problem sounds more similar to PDDL and less natural, the performance of all the models improves.
-Similar results hold for the Logistics domain (see results in Appendix C).
+Similar results hold for the Logistics domain (see results in Appendix D).
 This suggests that a more natural-sounding domain and problem description is much more challenging than templated, less natural sounding descriptions.
@@
 Another reason is a more natural sounding description may leave out implicit common-sense.
-For example, the Natural BlocksWorld-100 does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved.
+For example, the Natural BlocksWorld-100 data does not explicitly specify that a block is ‘clear’, because any human who reads that a block is “on top of a stack” can understand that there is no block on top of it and hence ‘clear’ to be moved.
 However, models often fail to invoke this knowledge and will leave out the ‘clear’ predicate, leading to unsolvable PDDL or incorrect plans.
@@
 However, gpt-4o-as-formalizer surpassed this baseline with a Correctness score of 70/100.
-This suggests that LLM-as-formalizer is robust to lexical perturbation, and its success is not due to memorization of the domain which is a part of the pretraining data.
+Surprisingly, Gemma models performed very well on the MysteryBlocksWorld-100 dataset with near perfect solvability and correctness for both models.
+We conclude that the results are mixed across models, which contradicts previous works cited involving PDDL generation which seem to claim that LLM-as-formalizer is clearly better than other methodologies.
+Despite the inconclusiveness among models, there are more clear-cut results between two groups of models: those that can formalize (eg.
+gpt-4o and o3-mini) and those that cannot formalize (eg.
+all Llama models and gpt-4o-mini).
+Among the models that can formalize, we find that the results are robust to lexical perturbations.
 5.5 What kind of errors?
-In this section, we discuss the kind of errors in PDDL generation.
+In this section, we discuss the kinds of errors made in PDDL generation.
 We perform an error analysis on a random 20 sample subset of problems where a plan was not found, or the found plan was not correct.
-From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and in the $\mathbb{PF}$.
-Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect action preconditions and effects, incorrect or missing predicates, and missing or incorrect action parameters.
-The error analysis for Natural BlocksWorld-100, Logistics-100 and Heavily Templated MysteryBlocksWorld-100 and can be found in Table 1.
+From there, we categorize the errors by syntax errors in either file, semantic errors in the $\mathbb{DF}$, and semantic errors in the $\mathbb{PF}$.
+Of the errors in the $\mathbb{DF}$, we determine finer-grained errors such as incorrect or missing actions, preconditions, effects, predicates, and parameters.
+The error analysis for both Natural BlocksWorld-100 and Logistics-100 can be found in Figure 5.
 For the open-source models, the most common error is syntax errors on BlocksWorld-100 and Logistics-100.
 For example, models repeatedly use the keyword ‘preconditions’ instead of ‘precondition’ which might suggest a lack of grasp of the PDDL syntax.
-Gemma models like gemma-2-27b make significantly less syntax errors (3 out of 20 on BlocksWorld-100) than Llama models like Llama-3.1-70B (20 out of 20), despite being smaller.
-Despite the syntax errors, there are still many semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$, which include missing predicates.
-As shown in Table 3, there is a significant gap between the number of plans that were found, and the number of found that were correct.
+Gemma models like gemma-2-27b make significantly fewer syntax errors (3 out of 20 on BlocksWorld-100) than Llama models like Llama-3.1-70B (20 out of 20), despite being smaller.
+Despite the syntax errors, open-sourced models still make many semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$, which include missing predicates and incorrect effects in the actions.
+We also observe a significant gap between the number of plans that were found, and the number of found plans that were correct.
 We find that the most common error made was swapping the parameters in the preconditions of the ‘stack’ action, leading to incorrect plans.
-While making much less syntax errors, GPT models frequently suffer from semantic errors.
+While making much fewer syntax errors, GPT models frequently suffer from semantic errors.
 Interestingly, the most common error made for gpt-4o come from the $\mathbb{PF}$, which is intuitively easier to generate than the $\mathbb{DF}$.
-Common errors in the $\mathbb{PF}$ include incorrect predicates in the initial state and goal state.
-Common errors in the $\mathbb{DF}$ is incorrect effects in an action.
+A common error in the $\mathbb{PF}$ is incorrect predicates in the initial state and goal state.
+A common error in the $\mathbb{DF}$ is incorrect effects in the actions.
 For example, in the ‘unstack’ action, the model does not make the next block ‘clear’ when the top block has been placed in the hand.
@@
 Due to the Heavily Templated descriptions, all the predicates would be listed out in the $\mathbb{PD}$ and the model would just need to match them to PDDL syntax in the $\mathbb{PF}$.
-While a similar essence, this is more of a challenge for $\mathbb{DF}$ since the clauses of preconditions and effects are more involved.
+While a similar essence, this is more challenging for the $\mathbb{DF}$ since the clauses of preconditions and effects are more involved.
 From Table 9, a similar trend between BlocksWorld-100 and MysteryBlocksWorld-100 also suggests that the LLM-as-formalizer methodology is robust to such perturbation.
+gemma-2bgemma-27bllama-8bllama-70bgpt-4o-minigpt-4o0010101010202020203030303040404040Has syntax err.No syntax error but has semantic err.
+Figure 5: Number of syntax errors (solver cannot parse PDDL) and semantic errors (solver outputs either no plan or a wrong plan) in 40 manually annotated, randomly sampled errors in Natural BlocksWorld-100 and Logistics-100.
+See full results in Table 8.
+Missing PredicateMissing ActionWrong ParametersWrong PreconditionsWrong Effects0010101010202020203030303040404040gpt-4o-minigpt-4o
+Figure 6: Number of fine-grained $\mathbb{DF}$ errors in 40 manually annotated, randomly sampled errors in Natural BlocksWorld-100 and Logistics-100. See full results in Table 9.
+5.6 The curious case of DeepSeek-R1
+The only exception to our main findings is DeepSeek-R1 for all datasets.
+The model shows strong performance as both a formalizer and planner for BlocksWorld-100, while showing mixed results for Logistics-100, and following a similar trend as GPT models for MysteryBlocksWorld-100.
+We can conclude that for simple datasets, using DeepSeek-R1 as a formalizer is preferred for some datasets, and when formalizing, the model is robust to lexical perturbations.
+However, for more complex datasets, using the model as a planner is preferred.
+This result may be due to how DeepSeek-R1 generates its response.
+The model outputs a reasoning string that thinks through every step of the plan and attempts to check its work by stepping through the plan.
+After examining several reasoning strings, we notice that when we use the model as a planner, it will think through the plan and when it gets stuck, it will attempt to pattern-match the example output given in the prompt.
+While the example output for each domain is not found in any of the datasets, we believe this may have caused the higher performance on more complex domains.
+6 Related Work
+Planning with LLMs There has been a large amount of research using LLMs for planning tasks.
+Some use LLMs for informal planning, also known as script or procedure learning Zhang et al.
+(2020); Lyu et al.
+(2021); Lal et al.
+(2024).
+While modern LLMs can make coherent and plausible informal plans, they are ungrounded and so lack executability and verifiability.
+Work that uses LLMs for formal planning in grounded environments generally concludes with the inability of such LLMs-as-planners Silver et al.
+(2024); Valmeekam et al.
+(2024a); Stechly et al.
+(2024); Valmeekam et al.
+(2024b).
+Follow-up work tackles this shortcoming by using the LLM as a heuristic, not just a planner, such as by proposing candidate plans that are iteratively verified Valmeekam et al.
+(2023); Kambhampati et al.
+(2024).
+While we consider the standard LLM-as-planner as a baseline, our focus is on LLM-as-formalizer, an alternative methodology for the same problem.
+LLMs as PDDL formalizer Here, LLMs do not provide plans but rather generate a PDDL representation of the domain and problem, which is then run through a solver to find the plan.
+This methodology has proven successful in a number of recent works, where the LLM generates different parts but not all of the PDDL for simplified evaluation.
+Zuo et al.
+(2024); Zhang et al.
+(2024a); Liu et al.
+(2023a) use the LLM to predict the entire $\mathbb{PF}$, while Xie et al.
+(2023); Lyu et al.
+(2023) predict just the goal for the $\mathbb{PF}$.
+Some predict parts of the $\mathbb{DF}$, such as Zhang et al.
+(2024c); Zhu et al.
+(2024) which generate the action semantics of the $\mathbb{DF}$ and Wong et al.
+(2023) who also predicts the predicates from a candidate list.
+Closest to our work is Guan et al.
+(2023) which predicts the $\mathbb{DF}$ as well as the $\mathbb{PF}$ goal.
+However, our work of holistically generating PDDL shows that coming up with the initial state in the $\mathbb{PF}$ is non-trivial (Section 5.5).
+Moreover, we vary the level of naturalness of descriptions in addition to the templated ones, which prove to be more challenging and insightful (Section 5.3).
+While the above discussions pertain to LLMs generating PDDL, many work on embodied agents outside the NLP community tackle similar problems with different focus Li et al. (2024).
+LLM code generation Our work hinges on modern LLMs’ ability to generate code Chen et al.
+(2021).
+In addition to writing or debugging programs Jiang et al.
+(2024), LLMs are also used to generate formal, interim representations that are not necessarily PDDL for problem solving.
+For example, Gao et al.
+(2023); Lyu et al.
+(2023); Tang et al.
+(2024) use the LLM to generate executable Python code for solving symbolic problems.
+In other work, the generated code may not be executable and is provided to another LLMs to facilitate reasoning Madaan et al.
+(2022); Zhang et al.
+(2023).
+A table comparing a couple of these works can be seen in Table 1 in the Appendix (Section F).
+7 Conclusion and Future Work
+Amidst the trend of using LLMs as formalizers, especially to generate PDDL for planning, our comprehensive analysis can be distilled into the following key findings and suggestions for future work.
+State-of-the-art LLMs can plan and formalize at a high level on simple domains like BlocksWorld, and to some degree on more complex domains like Logistics.
+Moreover, reasoning models trained to produce reasoning tokens before the answer (such as o3-mini and DeepSeek-R1) are superior planners than those who are not (such as gpt-4o).
+We suggest future work closely study these models (such as Valmeekam et al.
+(2024b)) from various angles such as the soundness of their reasoning chain, time and monetary cost, etc.
+Most small open LLMs cannot plan nor formalize even on simple domains, succumbing to syntax errors.
+To improve their performance on formalizing, we suggest future efforts on methods such as finetuning, modular prompting, constrained decoding, retrieval by language documentation, wrapper of a high-resource language, and so on to improve the performance of generating not just PDDL, but also low-resource languages in general.
+Formalizing is often superior to planning, but not always, with positive evidence for models like gpt-4o but negative evidence for models like DeepSeek-R1.
+Given the rising attention towards the LLM-as-formalizer methodology, we suggest future work validate this claim and extend to other environments such as more complex ones, partially observable ones, etc.
+Natural descriptions are challenging in planning tasks, and so is predicting complete PDDL.
+We acknowledge that translating a simple natural language instruction to a specific PDDL component might be sufficient in some confined domains and a fruitful endeavor.
+However, we argue that generating a planning representation as complete as possible given a natural language query as complex and realistic as possible is a prerequisite to open-domain planning.
+8 Limitation
+A common and valid criticism for using these simulations or text problems for evaluation is that their settings may be too contrived and removed from reality.
+Nevertheless, it is likely that LLMs’ satisfactory performance on these datasets is a necessary condition to success in real life.
+While we only consider zero-shot prompting, and a few experiments with few-shot and chain-of-thought prompts, without any attempt for prompt tuning, it is possible that the models’ performance significantly increases otherwise.
+Therefore, experimental results in all settings may be underestimated.
+Moreover, advanced prompting techniques such as self-refine, and voting can all potentially improve model performance.
+However, the study of those is out of the scope of this work.
+While we advocate for the LLM-as-formalizer methodology over LLM-as-planner, the former’s success may be dependent on the task.
+Highly symbolic tasks which can be relatively easily described, like BlocksWorld, are likely to favor LLM-as-formalizer.
+However, LLM-as-planner might shine in tasks with a more complex action space requiring common-sense knowledge that is easily accessed by pretraining.
+Furthermore, while we only consider the most straightforward LLM-as-planner prompting method, more involved methods, like Kambhampati et al.
+(2024) that combines LLM-as-planner with symbolic validation, will likely lead to a stronger baseline.
+Since this work uses only the BlocksWorld, Mystery BlocksWorld, Logistics and Barman domains, it is a small toy example to the usage of LLMs as formalizers and are not representative to problems in the real world, which would be much more challenging.
+This may pose a risk to users using this code on real world problems.
+The datasets we use and we propose are all under the MIT License.
+Acknowledgment
+We thank Peter Clark for providing invaluable input throughout this work and Andrew Zhu for providing technical support on LLM inference.
+Appendix A Data Examples
+As discussed above, each dataset comes with ground-truth PDDL describing domains ($\mathbb{DF}$) and problems ($\mathbb{PF}$).
+To maximize flexibility when performing analysis, we construct problem instances ourselves for some datasets, so that we can measure complexity with metrics like the number of blocks in BlocksWorld, for which we ensure a uniform distribution to avoid biases.
+Instances for BlocksWorld were randomly generated by varying the number of blocks and number of stacks in the initial and goal states from 2 to 15.
+Instances for Mystery BlocksWorld were randomly sampled from Valmeekam et al.
+(2024a).
+Instances of Logistics were taken directly from IPC (1998) and IPC (2000).
+Instances of Barman were generated by using Seipp et al.
+(2022) and varying the number of shot-glasses, ingredients and cocktails from 1 to 9.
+Below are the example PDDL and descriptions for BlocksWorld-100 and MysteryBlocksWorld-100.
+Due to the length of the examples, the PDDL and natural language descriptions for Logistics-100 and Barman-100 can be found at https://github.com/CassieHuang22/llm-as-pddl-formalizer/tree/main/examples.
+A.1 BlocksWorld-100 PDDL
+Listings 1 and 2 display examples of the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$ for BlocksWorld-100.
+The $\mathbb{DF}$ contains all four actions (pickup, putdown, stack and unstack) and their preconditions and post-conditions, as well as predicates needed for the domain.
+The $\mathbb{PF}$ contains the objects, initial state and goal state for the problem.
+A.2 BlocksWorld-100 $\mathbb{DD}$ and $\mathbb{PD}$
+Listings 3 and 4 display example $\mathbb{DD}$ and $\mathbb{PD}$ for the Heavily Templated setting in the BlocksWorld-100 dataset.
+In this setting, all preconditions and post-conditions are written out explicitly and sound similar to PDDL.
+The $\mathbb{PD}$ is similar, in that it lists all the predicates needed for to solve the task.
+Listings 5 and 6 display example $\mathbb{DD}$ and $\mathbb{PD}$ for Moderately Templated data in the BlocksWorld-100 dataset.
+In this setting, the $\mathbb{DD}$ and $\mathbb{PD}$ are much more natural than the Heavily Templated data, but all predicates are still listed.
+Finally Listings 7 and 8 display example $\mathbb{DD}$ and $\mathbb{PD}$ for Natural BlocksWorld-100 data.
+For the Natural data, we can see that the $\mathbb{DD}$ and $\mathbb{PD}$ give all necessary information to complete the task, but does not sound like PDDL, and does not describe all predicates needed to perform the task.
+We can see that the descriptions have all the same components as the $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL, but written in different levels of naturalness.
+A.3 MysteryBlocksWorld-100 PDDL
+Listings 9 and 10 display examples of the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$ for MysteryBlocksWorld-100.
+A.4 MysteryBlocksWorld-100 $\mathbb{DD}$ and $\mathbb{PD}$
+Listings 11 and 12 are example $\mathbb{DD}$ and $\mathbb{PD}$ of the Heavily Templated MysteryBlocksWorld-100.
+Text written in green demonstrates natural sounding text while text written in red demonstrates text that sounds the most like PDDL.
+Appendix B Prompts
+Listings 13, 14, 15 and 16 displays the prompts for all domains given to all models. Whenever possible, we asked the model to return the output in a JSON object for easier parsing.
+Appendix C Experimental Setup Details
+C.1 Planner
+We use a dual-bfws-ffparser planner to find plans from the generated PDDL. This is a best first width search planner used with a fast forward planner.
+Environment LLM predicts? Natural Descriptions?
+Zuo et al. (2024) fully-observed $\mathbb{PF}$ N
+Zhang et al. (2024a) partially-observed $\mathbb{PF}$ N
+Liu et al. (2023a) fully-observed $\mathbb{PF}$ N
+Xie et al. (2023) fully-observed & partially-observed $\mathbb{PF}$ goal N
+Lyu et al. (2023) fully-observed $\mathbb{PF}$ goal N
+Zhang et al. (2024c) procedural texts $\mathbb{DF}$ action semantics N
+Wong et al. (2023) partially-observed $\mathbb{DF}$ N
+Guan et al. (2023) fully-observed $\mathbb{DF}$ $\mathbb{PF}$ goal N
+Zhu et al. (2024) fully-observed $\mathbb{DF}$ action semantics N
+Tang et al. (2024) partially-observed Python N/A
+Silver et al. (2024) fully-observed plan N
+Valmeekam et al. (2024a) fully-observed plan N
+Stechly et al. (2024) fully-observed plan N
+This work fully-observed $\mathbb{DF}$ $\mathbb{PF}$ Y
+Table 1: Comparison with related work.
+C.2 VAL
+The VAL library takes in a ground-truth PDDL $\mathbb{DF}$, $\mathbb{PF}$ and a plan and tries to execute the plan in the environment by checking whether each action in the found plan can be executed based on the preconditions written in the ground-truth files.
+If the plan is executable, VAL then checks whether the final state after executing the plan matches the goal state found in the ground-truth $\mathbb{PF}$.
+If either the plan is not executable or the final state does not match the goal state, VAL will return an error, therefore the plan found by the planner is not correct.
+C.3 Evaluation Metrics
+We use solvability and correctness to evaluate the results from the LLMs.
+Solvability was calculated by running the planner and counting how many plans were returned.
+If a plan was not returned due to not finding a plan or an error occurred, it did not count towards solvability.
+Correctness was calculated by counting how many plans returned with no errors after running VAL, the plan was counted if VAL was able to execute the plan and there were no errors, and the final state is equal to the goal state in ground-truth problem file.
+Appendix D Detailed Results
+Beyond the visualizations above, we show the detailed results of all models on all simulations of all naturalness levels.
+D.1 Results for BlocksWorld-100
+Table 2 displays results for all results for BlocksWorld-100.
+D.2 Results for MysteryBlocksWorld-100
+Table 3 displays the results for Heavily Templated MysteryBlocksWorld-100.
+D.3 Results for Logistics-100
+Table 4 displays results for all naturalness settings for Logistics-100.
+D.4 Results for Barman-100
+Figure 7: Performance for Barman-100.
+Table 5 displays results for Barman-100.
+D.5 Few-Shot Prompting Results
+As mentioned in Section 4.3, we cannot provide the entire $\mathbb{DF}$ and $\mathbb{PF}$ in the prompt for few-shot prompting as we are asking the LLM to predict the entire $\mathbb{DF}$ and $\mathbb{PF}$.
+Therefore, we performed experiments where we provide a template for the $\mathbb{DF}$ and $\mathbb{PF}$ where portions of the PDDL are stubbed out.
+We performed experiments on Heavily Templated and Natural BlocksWorld-100 on a small subset of models.
+From Tables 2 and 6, we can see that results are mixed, with performance improving for some models like Llama-3.1-8B-Instruct and worse performance for others like gpt-4o on the Natural BlocksWorld-100 dataset.
+This suggests few-shot prompting does not improve performance for our experiments.
+D.6 Chain-of-Thought Prompting Results
+For chain-of-thought prompting, we ask the model to first write a domain description , then write the $\mathbb{DF}$, then write a problem description then write the $\mathbb{PF}$.
+These experiments were performed on Heavily Templated and Natural BlocksWorld-100 on a small subset of models.
+From Tables 2 and 7, we can see that results are mixed, with performance improvement on some models like gpt-3.5-turbo and performance decline on other models like gpt-4o on the Natural BlocksWorld-100 dataset.
+This suggests that chain-of-thought prompting does not improve performance for our experiments.
+Appendix E Detailed Error Analysis
+E.1 Syntax and Semantic Errors
+Table 8 displays syntax and semantic errors of models on Natural BlocksWorld-100 and Logistics-100. We do not analyze Barman-100 which is too challenging for models.
+Table 9 displays fine-grained $\mathbb{DF}$ errors of models on Natural BlocksWorld-100 and Logistics-100.
+E.2 Sample Model Output
+We display an example of the $\mathbb{DF}$ and $\mathbb{PF}$ that Llama-3.1-8B-Instruct generated.
+We can see that there are syntax errors, as well as semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$.
+We give the model the $\mathbb{DD}$ and $\mathbb{PD}$ in Listing 17 as input, which returned the $\mathbb{DF}$ and $\mathbb{PF}$ shown in Listings 18 and 19.
+Text in red displays errors outputted from the model.
+We can see that in the $\mathbb{DF}$ there are syntax errors (incorrect keyword “preconditions”) as well as semantic errors (incorrect predicates in preconditions and effects).
+For the output $\mathbb{PF}$ there are semantic errors, such as incorrect preconditions (a block cannot be clear and have another block on top of it) in the init section.
+Appendix F Related Works Comparison
+Table 1 compares works related to this paper.
+We can see that other works as the LLM to predict either the plan, parts of PDDL files and other languages.
+We can also see that other works have mostly templated natural language descriptions, while this work uses both templated and natural descriptions.
+Appendix G Methodology Comparison
+Both methodologies used in this paper incorporate LLMs into planning.
+There are pros and cons to each methodology.
+When using LLM-as-planner, we have a lightweight solution that returns results very quickly.
+However, due to the lack of reasoning skills in LLMs, they often struggle to create formal plans.
+Meanwhile using LLM-as-formalizer provides better executability and interpretability, though it uses a solver, which may results in getting results slower.
+We believe that for performance reasons that LLM-as-formalizer is the superior methodology.
+Listing 1: Domain File for BlocksWorld-100
+(define (domain blocksworld)
+(:predicates (clear ?x)
+(on-table ?x)
+(arm-empty)
+(holding ?x)
+(on ?x ?y))
+(:action pickup
+:parameters (?ob)
+:precondition (and (clear ?ob) (on-table ?ob) (arm-empty))
+:effect (and (holding ?ob) (not (clear ?ob)) (not (on-table ?ob))
+(not (arm-empty))))
+(:action putdown
+:parameters (?ob)
+:precondition (holding ?ob)
+:effect (and (clear ?ob) (arm-empty) (on-table ?ob)
+(not (holding ?ob))))
+(:action stack
+:parameters (?ob ?underob)
+:precondition (and (clear ?underob) (holding ?ob))
+:effect (and (arm-empty) (clear ?ob) (on ?ob ?underob)
+(not (clear ?underob)) (not (holding ?ob))))
+(:action unstack
+:parameters (?ob ?underob)
+:precondition (and (on ?ob ?underob) (clear ?ob) (arm-empty))
+:effect (and (holding ?ob) (clear ?underob)
+(not (on ?ob ?underob)) (not (clear ?ob)) (not (arm-empty)))))
+Listing 2: Problem File for BlocksWorld-100
+(define (problem blocksworld-p99)
+(:domain blocksworld)
+(:objects red blue green yellow )
+(:init
+(on-table red)
+(on blue red)
+(clear blue)
+(on-table green)
+(on yellow green)
+(clear yellow)
+(arm-empty)
+(:goal (and
+(on-table red)
+(on green red)
+(on yellow green)
+(on blue yellow)
+Listing 3: Domain Description for Heavily Templated BlocksWorld-100
+I am playing with a set of blocks. Here are the actions I can do
+Pickup block
+Unstack block from another block
+Putdown block
+Stack block on another block
+I have the following restrictions on my actions:
+To perform Pickup action, the following facts need to be true: clear block, block on table, arm-empty.
+Once Pickup action is performed the following facts will be true: holding block.
+Once Pickup action is performed the following facts will be false: clear block, block on table, arm-empty.
+To perform Putdown action, the following facts need to be true: holding block.
+Once Putdown action is performed the following facts will be true: clear block, block on table, arm-empty.
+Once Putdown action is performed the following facts will be false: holding block.
+To perform Stack action, the following needs to be true: clear block2, holding block1.
+Once Stack action is performed the following will be true: arm-empty, clear block1, block1 on block2.
+Once Stack action is performed the following will be false: clear block2, holding block1.
+To perform Unstack action, the following needs to be true: block1 on block2, clear block1, arm-empty.
+Once Unstack action is performed the following will be true: holding block1, clear block2.
+Once Unstack action is performed the following will be false:, block1 on block2, clear block1, arm-empty.
+Listing 4: Problem Description for Heavily Templated BlocksWorld-100
+As initial conditions I have that, the blue block is clear, the yellow block is clear, arm-empty, the blue block is on top of the red block, the yellow block is on top of the green block, the red block is on the table, and the green block is on the table.
+My goal is to have that the blue block is on top of the yellow block, the green block is on top of the red block, the yellow block is on top of the green block, and the red block is on the table.
+Listing 5: Domain Description for Moderately Templated BlocksWorld-100
+I am playing with a set of blocks where I need to arrange the blocks into stacks. Here are the actions I can do
+Pick up a block
+Unstack a block from on top of another block
+Put down a block
+Stack a block on top of another block
+I have the following restrictions on my actions:
+I can only pick up or unstack one block at a time.
+I can only pick up or unstack a block if my hand is empty.
+I can only pick up a block if the block is on the table and the block is clear. A block is clear if the block has no other blocks on top of it and if the block is not picked up.
+I can only unstack a block from on top of another block if the block I am unstacking was really on top of the other block.
+I can only unstack a block from on top of another block if the block I am unstacking is clear.
+Once I pick up or unstack a block, I am holding the block.
+I can only put down a block that I am holding.
+I can only stack a block on top of another block if I am holding the block being stacked.
+I can only stack a block on top of another block if the block onto which I am stacking the block is clear.
+Once I put down or stack a block, my hand becomes empty.
+Once you stack a block on top of a second block, the second block is no longer clear.
+Listing 6: Problem Description for Moderately Templated BlocksWorld-100
+As initial conditions I have that, the blue block is clear, the yellow block is clear, the hand is empty, the blue block is on top of the red block, the yellow block is on top of the green block, the red block is on the table, and the green block is on the table.
+My goal is to have that the blue block is on top of the yellow block, the green block is on top of the red block, the yellow block is on top of the green block, and the red block is on the table.
+Listing 7: Domain Description for Natural BlocksWorld-100
+The Blocksworld game involves a set of blocks of different colors, which can be stacked on top of each other or placed on the table.
+The objective is to move the blocks from an initial configuration to a goal configuration using a series of legal moves.
+Legal moves in Blocksworld include: picking up a block from the table or from the top of another block, stacking a block onto the table, or stacking a block onto another block.
+Listing 8: Problem Description for Natural BlocksWorld-100
+In this particular game, there are 4 blocks: a red block, a blue block, a green block, and a yellow block.
+At the start, the red block is on the table, the blue block is on top of the red block, the green block is on the table, and the yellow block is on top of the green block.
+The goal is to have the red block on the table, the green block on top of the red block, the yellow block on top of the green block, and the blue block on top of the yellow block.
+Listing 9: Domain File for Mystery_BlocksWorld-100
+(define (domain mystery_blocksworld)
+(:predicates (province ?x)
+(planet ?x)
+(pain ?x)
+(craves ?x ?y))
+(:action attack
+:parameters (?ob)
+:precondition (and (province ?ob) (planet ?ob) (harmony))
+:effect (and (pain ?ob) (not (province ?ob)) (not (planet ?ob))
+(not (harmony))))
+(:action succumb
+:parameters (?ob)
+:precondition (pain ?ob)
+:effect (and (province ?ob) (harmony) (planet ?ob)
+(not (pain ?ob))))
+(:action overcome
+:parameters (?ob ?underob)
+:precondition (and (province ?underob) (pain ?ob))
+:effect (and (harmony) (province ?ob) (craves ?ob ?underob)
+(not (province ?underob)) (not (pain ?ob))))
+(:action feast
+:parameters (?ob ?underob)
+:precondition (and (craves ?ob ?underob) (province ?ob) (harmony))
+:effect (and (pain ?ob) (province ?underob)
+(not (craves ?ob ?underob)) (not (province ?ob)) (not (harmony)))))
+Listing 10: Problem File for Mystery_BlocksWorld-100
+(define (problem mystery_blocksworld-p01)
+(:domain mystery_blocksworld)
+(:objects a b c d )
+(:init
+(craves a b)
+(craves b c)
+(planet c)
+(planet d)
+(province a)
+(province d)
+(:goal (and
+(craves a d)
+(craves c a)
+Listing 11: Domain Description for Mystery_BlocksWorld-100
+I am playing with a set of objects. Here are the actions I can do
+Attack object
+Feast object from another object
+Succumb object
+Overcome object from another object
+I have the following restrictions on my actions:
+To perform Attack action, the following facts need to be true: Province object, Planet object,
+Harmony.
+Once Attack action is performed the following facts will be true: Pain object.
+Once Attack action is performed the following facts will be false: Province object, Planet object,
+Harmony.
+To perform Succumb action, the following facts need to be true: Pain object.
+Once Succumb action is performed the following facts will be true: Province object, Planet object,
+Harmony.
+Once Succumb action is performed the following facts will be false: Pain object.
+To perform Overcome action, the following needs to be true: Province other object, Pain object.
+Once Overcome action is performed the following will be true: Harmony, Province object,
+Object Craves other object.
+Once Overcome action is performed the following will be false: Province other object, Pain object.
+To perform Feast action, the following needs to be true: Object Craves other object, Province
+object, Harmony.
+Once Feast action is performed the following will be true: Pain object, Province other object.
+Once Feast action is performed the following will be false:, Object Craves other object, Province
+object, Harmony.
+Listing 12: Problem Description for Mystery_BlocksWorld-100
+As initial conditions I have that, object a craves object b, object b craves object c, harmony,
+planet object c, planet object d, province object a and province object d.
+My goal is to have that object a craves object d and object c craves object a.
+Listing 13: Prompt for LLM-as-Planner on BlocksWorld-100
+Here is a game we are playing.
+{domain_description}
+{problem_description}
+Write the plan that would solve this
+problem.
+These are the available actions:
+(PICK-UP block): pick up a block from the
+table
+(PUT-DOWN block): put down a block on the
+table
+(STACK block1 block2): stack block1 onto
+block2
+(UNSTACK block1 block2): unstack block1
+from block2
+Here is what the output should look like:
+(PICK-UP A)
+(STACK A B)
+(UNSTACK A B)
+(PUT-DOWN A)
+Listing 14: Prompt for LLM-as-Planner on Mystery_BlocksWorld-100
+Here is a game we are playing.
+{domain_description}
+{problem_description}
+Write the plan that would solve this
+problem.
+These are the available actions:
+(ATTACK object): attack object
+(SUCCUMB object): succumb
+(OVERCOME object1 object2): overcome object1 from object2
+(FEAST object1 object2): feast object1 from object2
+Here is what the output should look like:
+(ATTACK A)
+(OVERCOME A B)
+(FEAST A B)
+(SUCCUMB A)
+Listing 15: Prompt for LLM-as-Planner on Logistics-100
+Here is a game we are playing.
+{domain_description}
+{problem_description}
+Write the plan that would solve this
+problem.
+These are the available actions:
+(load-truck package truck location): load a package onto a truck at a location
+(load-airplane object airplane location): load a package onto an airplane at a location
+(unload-truck package truck location): unload a package from a truck at a location
+(unload-airplane package airplane location): unload a package from an airplan at a location
+(drive-truck truck location1 location2 city): drive a truck from location1 to location2 in a city
+(fly-airplane airplane location1 location2): fly an airplane from location1 to location2
+Here is what the output should look like:
+(load-truck package truck city1-1)
+(drive-truck truck city1-1 city1-2 city1)
+(unload-truck package truck city1-2)
+(load-airplane package plane city1-2)
+(fly-airplane plane city1-2 city2-1)
+(unload-airplane package plane city2-1)
+Listing 16: Prompt for LLM-as-Planner on Barman-100
+Here is a game we are playing.
+{domain_description}
+{problem_description}
+Write the plan that would solve this
+problem.
+These are the available actions:
+(GRASP hand container): grasp container with hand
+(LEAVE hand container): leave container with hand
+(FILL-SHOT shot ingredient hand1 hand2 dispenser): fill shot with ingredient from dispenser using hand1 and hand2
+(REFILL-SHOT shot ingredient hand1 hand2 dispenser): re-fill shot with ingredient from dispenser using hand1 and hand2
+(EMPTY-SHOT hand shot beverage): empty shot containing beverage with hand
+(CLEAN-SHOT shot beverage hand1 hand2): clean shot containing beverage using hand1 and hand2
+(POUR-SHOT-TO-CLEAN-SHAKER shot ingredient shaker hand level level1): pour shot containing ingredient into clean shaker using hand so that the level in the shaker changes from level to level1
+(POUR-SHOT-TO-USED-SHAKER shot ingredient shaker hand level level1): pour shot containing ingredient into used shaker using hand so that the level in the shaker changes from level to level1
+(EMPTY-SHAKER hand shaker cocktail level level1): empty shaker containing cocktail with hand so that the level in the shaker changes from level to level1
+(CLEAN-SHAKER hand1 hand2 shaker): clean shaker using hand1 and hand2
+(SHAKE cocktail ingredient1 ingredient2 shaker hand1 hand2): shake shaker containing cocktail of ingredient1 and ingredient2 using hand1 and hand2
+(POUR-SHAKER-TO-SHOT beverage shot hand shaker level level1): pour shaker containing beverage into shot using hand so that the level in the shaker changes from level to level1
+Here is what the output should look like:
+(GRASP right shot1)
+(FILL-SHOT shot1 ingredient10 right left dispenser10)
+(POUR-SHOT-TO-CLEAN-SHAKER shot1 ingredient10 shaker1 right l0 l1)
+(CLEAN-SHOT shot1 ingredient10 right left)
+(FILL-SHOT shot1 ingredient5 right left dispenser5)
+(GRASP left shaker1)
+(POUR-SHOT-TO-USED-SHAKER shot1 ingredient5 shaker1 right l1 l2)
+(LEAVE right shot1)
+(SHAKE cocktail1 ingredient5 ingredient10 shaker1 left right)
+(LEAVE left shaker1)
+(GRASP left shot1)
+(CLEAN-SHOT shot1 ingredient5 left right)
+(GRASP right shaker1)
+(POUR-SHAKER-TO-SHOT cocktail1 shot1 right shaker1 l2 l1)
+Metrics
+Natural Moderately Templated Heavily Templated
+Models Solvability Correctness Solvability Correctness Solvability Correctness
+gemma-2-9b-it 3/100 3/100 0/100 0/100 61/100 10/100
+gemma-2-9b-itp - 9/100 - 5/100 - 5/100
+gemma-2-27b-it 0/100 0/100 17/100 10/100 81/100 80/100
+gemma-2-27b-itp - 11/100 - 3/100 - 7/100
+Llama-3.1-8B 0/100 0/100 0/100 0/100 0/100 0/100
+Llama-3.1-8Bp - 1/100 - 1/100 - 0/100
+Llama-3.1-70B 0/100 0/100 0/100 0/100 0/100 0/100
+Llama-3.1-70Bp - 13/100 - 10/100 - 10/100
+DeepSeek-R1-Distill-Llama-8B 0/100 0/100 0/100 0/100 0/100 0/100
+DeepSeek-R1-Distill-Llama-8Bp - 2/100 - 2/100 - 2/100
+DeepSeek-R1-Distill-Llama-70B 0/100 0/100 0/100 0/100 0/100 0/100
+DeepSeek-R1-Distill-Llama-70Bp - 18/100 - 25/100 - 22/100
+DeepSeek-R1 83/100 81/100 76/100 75/100 84/100 84/100
+DeepSeek-R1p - 91/100 - 85/100 - 86/100
+gpt-3.5-turbo 2/100 1/100 14/100 4/100 39/100 29/100
+gpt-4o-mini 19/100 3/100 9/100 5/100 66/100 59/100
+gpt-4o-minip - 7/100 - 7/100 - 1/100
+gpt-4o 64/100 60/100 77/100 67/100 89/100 89/100
+gpt-4op - 33/100 - 35/100 - 29/100
+o1-preview 91/100 91/100 - - - -
+o1-previewp - 82/100 - - - -
+o3-mini 79/100 68/100 82/100 70/100 94/100 94/100
+o3-minip - 87/100 - 87/100 - 96/100
+Table 2: Performance of LLM-as-formalizer and LLM-as-planner (p) on all naturalness settings of the BlocksWorld-100 dataset.
+Models Solvability Correctness
+Llama-3.1-8B 0/100 0/100
+Llama-3.1-8Bp - 0/100
+Llama-3.1-70B 0/100 0/100
+Llama-3.1-70Bp - 0/100
+gemma-2-9b-it 100/100 99/100
+gemma-2-9b-itp - 9/100
+gemma-2-27b-it 99/100 98/100
+gemma-2-27b-itp - 0/100
+DeepSeek-R1 56/100 54/100
+DeepSeek-R1p - 21/100
+gpt-3.5-turbo 4/100 0/100
+gpt-4o-mini 36/100 5/100
+gpt-4o-minip - 0/100
+gpt-4o 74/100 70/100
+gpt-4op - 0/100
+o3-mini 95/100 95/100
+o3-minip - 74/100
+Table 3: Performance of LLM-as-formalizer and LLM-as-planner (p) on the Heavily Templated MysteryBlocksWorld-100 dataset.
+Metrics
+Natural Moderately Templated Heavily Templated
+Models Solvability Correctness Solvability Correctness Solvability Correctness
+gemma-2-9b-it 1/100 0/100 0/100 0/100 1/100 0/100
+gemma-2-9b-itp - 0/100 - 0/100 - 0/100
+gemma-2-27b-it 1/100 0/100 3/100 0/100 0/100 0/100
+gemma-2-27b-itp - 0/100 - 0/100 - 1/100
+Llama-3.1-8B 0/100 0/100 0/100 0/100 0/100 0/100
+Llama-3.1-8Bp - 1/100 - 0/100 - 0/100
+Llama-3.1-70B 0/100 0/100 0/100 0/100 0/100 0/100
+Llama-3.1-70Bp - 0/100 - 0/100 - 0/100
+DeepSeek-R1 51/100 9/100 73/100 73/100 64/100 64/100
+DeepSeek-R1p - 12/100 - 22/100 - 22/100
+gpt-3.5-turbo 1/100 0/100 0/100 0/100 1/100 1/100
+gpt-3.5-turbop - 0/100 - 0/100 - 0/100
+gpt-4o-mini 13/100 0/100 29/100 0/100 26/100 0/100
+gpt-4o-minip - 0/100 - 0/100 - 0/100
+gpt-4o 20/100 2/100 34/100 13/100 33/100 13/100
+gpt-4op - 1/100 - 0/100 - 0/100
+o3-mini 43/100 7/100 51/100 39/100 55/100 47/100
+o3-minip - 14/100 - 21/100 - 21/100
+Table 4: Performance of LLM-as-formalizer and LLM-as-planner (p) on all Logistics-100 datasets.
+Models Solvability Correctness
+Llama-3.1-8B 0/100 0/100
+Llama-3.1-8Bp - 0/100
+Llama-3.1-70B 0/100 0/100
+Llama-3.1-70Bp - 15/100
+gemma-2-9b-it 0/100 0/100
+gemma-2-9b-itp - 2/100
+gemma-2-27b-it 0/100 0/100
+gemma-2-27b-itp - 0/100
+DeepSeek-R1 10/100 2/100
+DeepSeek-R1p - 22/100
+gpt-3.5-turbo 0/100 0/100
+gpt-3.5-turbop - 4/100
+gpt-4o-mini 0/100 0/100
+gpt-4o-minip - 8/100
+gpt-4o 1/100 0/100
+gpt-4op - 3/100
+o3-mini 1/100 0/100
+o3-minip - 19/100
+Table 5: Performance of LLM-as-formalizer and LLM-as-planner (p) on the Heavily Templated Barman-100 dataset.
+Metrics
+Natural Heavily Templated
+Models Solvability Correctness Solvability Correctness
+Llama-3.1-8B 1/100 1/100 49/100 45/100
+Llama-3.1-70B 0/100 0/100 0/100 0/100
+gpt-3.5-turbo 8/100 1/100 35/100 24/100
+gpt-4o-mini 13/100 3/100 75/100 61/100
+gpt-4o 60/100 48/100 93/100 92/100
+Table 6: Performance of LLM-as-formalizer all the Natural and Heavily Templated BlocksWorld-100 datasets using few-shot prompting.
+Metrics
+Natural Heavily Templated
+Models Solvability Correctness Solvability Correctness
+Llama-3.1-8B 0/100 0/100 0/100 0/100
+Llama-3.1-70B 0/100 0/100 0/100 0/100
+gpt-3.5-turbo 6/100 0/100 23/100 17/100
+gpt-4o-mini 19/100 4/100 49/100 45/100
+gpt-4o 61/100 48/100 91/100 90/100
+Table 7: Performance of LLM-as-formalizer all the Natural and Heavily Templated BlocksWorld-100 datasets using chain-of-thought prompting.
 Models Syntax Error $\mathbb{DF}$ Error $\mathbb{PF}$ Error
@@
 gpt-4o 5/20 16/20 0/20
-Table 1: Error analysis of LLM-as-formalizer on various datasets, manually annotated on a 20–example subset.
+Table 8: Annotated types of 20 randomly sampled errors of LLM-as-formalizer for various models and datasets.
 Models Wrong Precondition Wrong Effect Missing Predicate Missing Action Missing Parameters
@@
 Table 9: Analysis of errors found in $\mathbb{DF}$ for Natural BlocksWorld-100 and Heavily Templated MysteryBlocksWorld-100 out of 20 randomly sampled instances.
-6 Related Work
-Planning with LLMs There has been a large amount of research using LLMs for planning tasks.
-Some use LLMs for informal planning, also known as script or procedure learning Zhang et al.
-(2020); Lyu et al.
-(2021); Lal et al.
-(2024).
-While modern LLMs can make coherent and plausible informal plans, they are ungrounded and so lack executability and verifiability.
-Work that use LLMs for formal planning in grounded environments generally conclude the inability of such LLMs-as-planners Silver et al.
-(2024); Valmeekam et al.
-(2024); Stechly et al.
-(2024).
-Follow-up work tackles this shortcoming by using the LLM as a heuristic, not just a planner, such as by proposing candidate plans that are iteratively verified Valmeekam et al.
-(2023); Kambhampati et al.
-(2024).
-While we consider the standard LLM-as-planner as a baseline, our focus is on LLM-as-formalizer, an alternative methodology for the same problem.
-LLMs as PDDL formalizer Here, LLMs do not provide plans but rather generate the a PDDL representation of the domain and problem, which is then run through a solver to find the plan.
-This methodology has proven successful in a number of recent works, where the LLM generates different parts but not all of the PDDL for simplified evaluation.
-Zuo et al.
-(2024); Zhang et al.
-(2024a); Liu et al.
-(2023a) use the LLM to predict the entire $\mathbb{PF}$, while Xie et al.
-(2023); Lyu et al.
-(2023) predict just the goal for the $\mathbb{PF}$.
-Some predict the $\mathbb{DF}$, such as Zhang et al.
-(2024c); Zhu et al.
-(2024) that generate the action semantics of the $\mathbb{DF}$ and Wong et al.
-(2023) who also predicts the predicates from a candidate list.
-Closest to our work is Guan et al.
-(2023) which predicts the $\mathbb{DF}$ as well as the $\mathbb{PF}$ goal.
-However, our work of holistically generating PDDL shows that coming up with the initial state in the $\mathbb{PF}$ is non-trivial (Section 5.5).
-Moreover, we vary the level of naturalness of descriptions in addition to the templated ones, which prove to be more challenging and insightful (Section 5.3).
-While the above discussions pertain to LLMs generating PDDL, many work on embodied agents outside the NLP community tackle similar problems with different focus Li et al. (2024).
-LLM code generation Our work hinges on modern LLMs’ ability to generate code Chen et al.
-(2021).
-In addition to writing or debugging programs Jiang et al.
-(2024), LLMs are also used to generate formal, interim representations that are not necessarily PDDL for problem solving.
-For example, Gao et al.
-(2023); Lyu et al.
-(2023); Tang et al.
-(2024) use the LLM to generate executable Python code for solving symbolic problems.
-In other work, the generated code may not be executable and is provided to another LLMs to facilitate reasoning Madaan et al.
-(2022); Zhang et al.
-(2023).
-A table comparing a couple of these works can be seen in Table 7 in the Appendix (Section E).
-7 Conclusion
-We explore the limit of state-of-the-art LLMs to be used as a PDDL formalizer for planning with natural language descriptions of different naturalness levels.
-While the LLM-as-formalizer methodology greatly outperforms the LLM-as-planner baseline in various planning domains, we conclude that with zero-shot prompting, only GPT models are sufficiently capable for the task.
-Therefore, future work should attempt to equip open-source models with similar ability to democratize the ability of making executable plans.
-We also find that LLM-as-formalizer is robust to lexical perturbation, demonstrating strong performance in long-tail domains that are underrepresented in pretraining.
-Our work will inform future efforts of using LLM as a planning formalizer, including experiments on partially-observed environments that require exploration and interaction, more complex environments with a larger action space, and so on.
-8 Limitation
-A common and valid criticism for using those simulations or text problems for evaluation is that these settings may be too contrived and removed from the reality.
-Nevertheless, it is likely that LLMs’ satisfactory performance on these datasets is a necessary condition to success in real life.
-While we only consider zero-shot prompting without any attempt for prompt tuning, it is possible that the models’ performance significantly increases otherwise.
-Therefore, experimental results in all settings may be underestimated.
-Moreover, advanced prompting techniques such as chain-of-thought, self-refine, and voting can all potentially improve model performance.
-However, the study of those is out of the scope of this work.
-While we advocate for the LLM-as-formalizer methodology over LLM-as-planner, the former’s success may be dependent on the task.
-Highly symbolic tasks which can be relatively easily described, like BlocksWorld, are likely to favor LLM-as-formalizer.
-However, LLM-as-planner might shine in tasks with a more complex action space requiring common-sense knowledge that is easily accessed by pretraining.
-Furthermore, while we only consider the most straightforward LLM-as-planner prompting method, more involved methods, like Kambhampati et al.
-(2024) that combines LLM-as-planner with symbolic validation, will likely lead to a stronger baseline.
-Since this work uses only the BlocksWorld, Mystery BlocksWorld, Logistics and Barman domains, it is a small toy example to the usage of LLMs as formalizers and are not representative to problems in the real world, which would be much more challenging.
-This may pose a risk to users using this code on real world problems.
-The datasets we use and we propose are all under the MIT License.
-Acknowledgment
-We thank Peter Clark for providing invaluable input throughout this work and Andrew Zhu for providing technical support on LLM inference.
-Appendix A Data Examples
-As discussed above, each dataset comes with ground-truth PDDL describing domains ($\mathbb{DF}$) and problems ($\mathbb{PF}$).
-To maximize flexibility when performing analysis, we construct problem instances ourselves for some datasets, so that we can measure complexity with metrics like the number of Blocks in BlocksWorld, for which we ensure a uniform distribution to avoid biases.
-Instances for BlocksWorld were randomly generated by varying the number of blocks and number of stacks in the initial and goal states from 2 to 15.
-Instances for Mystery BlocksWorld were randomly sampled from Valmeekam et al.
-(2024).
-Instances of Logistics were taken directly from IPC (1998) and IPC (2000).
-Instances of Barman were generated by using Seipp et al.
-(2022) and varying the number of shot-glasses, ingredients and cocktails from 1 to 9.
-Below are the example PDDL and descriptions for all datasets.
-A.1 BlocksWorld-100 PDDL
-The following are an example of the ground-truth $\mathbb{DF}$ and $\mathbb{PF}$ for BlocksWorld-100.
-$\mathbb{DF}$:
-(define (domain blocksworld) (:predicates (clear ?x) (on-table ?x) (arm-empty) (holding ?x) (on ?x ?y)) (:action pickup :parameters (?ob) :precondition (and (clear ?ob) (on-table ?ob) (arm-empty)) :effect (and (holding ?ob) (not (clear ?ob)) (not (on-table ?ob)) (not (arm-empty)))) (:action putdown :parameters (?ob) :precondition (holding ?ob) :effect (and (clear ?ob) (arm-empty) (on-table ?ob) (not (holding ?ob)))) (:action stack :parameters (?ob ?underob) :precondition (and (clear ?underob) (holding ?ob)) :effect (and (arm-empty) (clear ?ob) (on ?ob ?underob) (not (clear ?underob)) (not (holding ?ob)))) (:action unstack :parameters (?ob ?underob) :precondition (and (on ?ob ?underob) (clear ?ob) (arm-empty)) :effect (and (holding ?ob) (clear ?underob) (not (on ?ob ?underob)) (not (clear ?ob)) (not (arm-empty)))))
-The $\mathbb{DF}$ contains all four actions (pickup, putdown, stack and unstack) and their pre-conditions and post-conditions, as well as predicates needed for the domain.
-$\mathbb{PF}$:
-(define (problem blocksworld-p99) (:domain blocksworld) (:objects red blue green yellow ) (:init (on-table red) (on blue red) (clear blue) (on-table green) (on yellow green) (clear yellow) (arm-empty) ) (:goal (and (on-table red) (on green red) (on yellow green) (on blue yellow) )) )
-The $\mathbb{PF}$ contains the objects, initial state and goal state for the problem.
-A.2 BlocksWorld-100 $\mathbb{DD}$ and $\mathbb{PD}$
-The following display example $\mathbb{DD}$ and $\mathbb{PD}$ for all natural settings in the BlocksWorld-100 dataset.
-We can see that the descriptions have all the same components as the $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL, but written in different levels of naturalness.
-For the Heavily Templated $\mathbb{DD}$, all preconditions and post-conditions are written out explicitly and sound similar to PDDL.
-The $\mathbb{PD}$ is similar, in that it lists all the predicates needed for to solve the task.
-Heavily Templated $\mathbb{DD}$:
-I am playing with a set of blocks.
-Here are the actions I can do Pickup block Unstack block from another block Putdown block Stack block on another block I have the following restrictions on my actions: To perform Pickup action, the following facts need to be true: clear block, block on table, arm-empty.
-Once Pickup action is performed the following facts will be true: holding block.
-Once Pickup action is performed the following facts will be false: clear block, block on table, arm-empty.
-To perform Putdown action, the following facts need to be true: holding block.
-Once Putdown action is performed the following facts will be true: clear block, block on table, arm-empty.
-Once Putdown action is performed the following facts will be false: holding block.
-To perform Stack action, the following needs to be true: clear block2, holding block1.
-Once Stack action is performed the following will be true: arm-empty, clear block1, block1 on block2.
-Once Stack action is performed the following will be false: clear block2, holding block1.
-To perform Unstack action, the following needs to be true: block1 on block2, clear block1, arm-empty.
-Once Unstack action is performed the following will be true: holding block1, clear block2.
-Once Unstack action is performed the following will be false:, block1 on block2, clear block1, arm-empty.
-Heavily Templated $\mathbb{PD}$:
-As initial conditions I have that, the blue block is clear, the yellow block is clear, arm-empty, the blue block is on top of the red block, the yellow block is on top of the green block, the red block is on the table, and the green block is on the table.
-My goal is to have that the blue block is on top of the yellow block, the green block is on top of the red block, the yellow block is on top of the green block, and the red block is on the table.
-For the Moderately Templated data, the $\mathbb{DD}$ and $\mathbb{PD}$ are much more natural than the Heavily Templated data, but all predicates are still listed.
-Moderately Templated $\mathbb{DD}$:
-I am playing with a set of blocks where I need to arrange the blocks into stacks.
-Here are the actions I can do Pick up a block Unstack a block from on top of another block Put down a block Stack a block on top of another block I have the following restrictions on my actions: I can only pick up or unstack one block at a time.
-I can only pick up or unstack a block if my hand is empty.
-I can only pick up a block if the block is on the table and the block is clear.
-A block is clear if the block has no other blocks on top of it and if the block is not picked up.
-I can only unstack a block from on top of another block if the block I am unstacking was really on top of the other block.
-I can only unstack a block from on top of another block if the block I am unstacking is clear.
-Once I pick up or unstack a block, I am holding the block.
-I can only put down a block that I am holding.
-I can only stack a block on top of another block if I am holding the block being stacked.
-I can only stack a block on top of another block if the block onto which I am stacking the block is clear.
-Once I put down or stack a block, my hand becomes empty.
-Once you stack a block on top of a second block, the second block is no longer clear.
-Moderately Templated $\mathbb{PD}$:
-As initial conditions I have that, the blue block is clear, the yellow block is clear, the hand is empty, the blue block is on top of the red block, the yellow block is on top of the green block, the red block is on the table, and the green block is on the table.
-My goal is to have that the blue block is on top of the yellow block, the green block is on top of the red block, the yellow block is on top of the green block, and the red block is on the table.
-Finally for the natural data, we can see that the $\mathbb{DD}$ and $\mathbb{PD}$ give all necessary information to complete the task, but does not sound like PDDL, and does not describe all predicates needed to perform the task.
-Natural $\mathbb{DD}$:
-The Blocksworld game involves a set of blocks of different colors, which can be stacked on top of each other or placed on the table.
-The objective is to move the blocks from an initial configuration to a goal configuration using a series of legal moves.
-Legal moves in Blocksworld include: picking up a block from the table or from the top of another block, stacking a block onto the table, or stacking a block onto another block.
-Natural $\mathbb{PD}$:
-In this particular game, there are 4 blocks: a red block, a blue block, a green block, and a yellow block.
-At the start, the red block is on the table, the blue block is on top of the red block, the green block is on the table, and the yellow block is on top of the green block.
-The goal is to have the red block on the table, the green block on top of the red block, the yellow block on top of the green block, and the blue block on top of the yellow block.
-A.3 MysteryBlocksWorld-100 PDDL
-This section displays an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for MysteryBlocksWorld-100.
-$\mathbb{DF}$:
-(define (domain mystery_blocksworld) (:predicates (province ?x) (planet ?x) (harmony) (pain ?x) (craves ?x ?y)) (:action attack :parameters (?ob) :precondition (and (province ?ob) (planet ?ob) (harmony)) :effect (and (pain ?ob) (not (province ?ob)) (not (planet ?ob)) (not (harmony)))) (:action succumb :parameters (?ob) :precondition (pain ?ob) :effect (and (province ?ob) (harmony) (planet ?ob) (not (pain ?ob)))) (:action overcome :parameters (?ob ?underob) :precondition (and (province ?underob) (pain ?ob)) :effect (and (harmony) (province ?ob) (craves ?ob ?underob) (not (province ?underob)) (not (pain ?ob)))) (:action feast :parameters (?ob ?underob) :precondition (and (craves ?ob ?underob) (province ?ob) (harmony)) :effect (and (pain ?ob) (province ?underob) (not (craves ?ob ?underob)) (not (province ?ob)) (not (harmony)))))
-$\mathbb{PF}$:
-(define (problem mystery_blocksworld-p01) (:domain mystery_blocksworld) (:objects a b c d ) (:init (craves a b) (craves b c) (harmony) (planet c) (planet d) (province a) (province d) ) (:goal (and (craves a d) (craves c a) )) )
-A.4 MysteryBlocksWorld-100 $\mathbb{DD}$ and $\mathbb{PD}$
-The following are example $\mathbb{DD}$ and $\mathbb{PD}$ of the Heavily Templated MysteryBlocksWorld-100.
-Text written in green demonstrates natural sounding text while text written in red demonstrates text that sounds the most like PDDL.
-$\mathbb{DD}$:
-I am playing with a set of objects.
-Here are the actions I can do Attack object Feast object from another object Succumb object Overcome object from another object I have the following restrictions on my actions: To perform Attack action, the following facts need to be true: Province object, Planet object, Harmony.
-Once Attack action is performed the following facts will be true: Pain object.
-Once Attack action is performed the following facts will be false: Province object, Planet object, Harmony.
-To perform Succumb action, the following facts need to be true: Pain object.
-Once Succumb action is performed the following facts will be true: Province object, Planet object, Harmony.
-Once Succumb action is performed the following facts will be false: Pain object.
-To perform Overcome action, the following needs to be true: Province other object, Pain object.
-Once Overcome action is performed the following will be true: Harmony, Province object, Object Craves other object.
-Once Overcome action is performed the following will be false: Province other object, Pain object.
-To perform Feast action, the following needs to be true: Object Craves other object, Province object, Harmony.
-Once Feast action is performed the following will be true: Pain object, Province other object.
-Once Feast action is performed the following will be false:, Object Craves other object, Province object, Harmony.
-$\mathbb{PD}$:
-As initial conditions I have that, object a craves object b, object b craves object c, harmony, planet object c, planet object d, province object a and province object d.
-My goal is to have that object a craves object d and object c craves object a.
-A.5 Logistics-100 PDDL
-The following are an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for Logistics-100.
-$\mathbb{DF}$:
-(define (domain logistics) (:requirements :strips) (:predicates (package ?obj) (truck ?truck) (airplane ?airplane) (airport ?airport) (location ?loc) (in-city ?obj ?city) (city ?city) (at ?obj ?loc) (in ?obj ?obj)) (:action load-truck :parameters (?obj ?truck ?loc) :precondition (and (package ?obj) (truck ?truck) (location ?loc) (at ?truck ?loc) (at ?obj ?loc)) :effect (and (not (at ?obj ?loc)) (in ?obj ?truck))) (:action load-airplane :parameters (?obj ?airplane ?loc) :precondition (and (package ?obj) (airplane ?airplane) (location ?loc) (at ?obj ?loc) (at ?airplane ?loc)) :effect (and (not (at ?obj ?loc)) (in ?obj ?airplane))) (:action unload-truck :parameters (?obj ?truck ?loc) :precondition (and (package ?obj) (truck ?truck) (location ?loc) (at ?truck ?loc) (in ?obj ?truck)) :effect (and (not (in ?obj ?truck)) (at ?obj ?loc)))
-$\mathbb{DF}$ (cont’d)
-(:action unload-airplane :parameters (?obj ?airplane ?loc) :precondition (and (package ?obj) (airplane ?airplane) (location ?loc) (in ?obj ?airplane) (at ?airplane ?loc)) :effect (and (not (in ?obj ?airplane)) (at ?obj ?loc))) (:action drive-truck :parameters (?truck ?loc-from ?loc-to ?city) :precondition (and (truck ?truck) (location ?loc-from) (location ?loc-to) (city ?city) (at ?truck ?loc-from) (in-city ?loc-from ?city) (in-city ?loc-to ?city)) :effect (and (not (at ?truck ?loc-from)) (at ?truck ?loc-to))) (:action fly-airplane :parameters (?airplane ?loc-from ?loc-to) :precondition (and (airplane ?airplane) (airport ?loc-from) (airport ?loc-to) (at ?airplane ?loc-from)) :effect (and (not (at ?airplane ?loc-from)) (at ?airplane ?loc-to))) )
-The $\mathbb{DF}$ contains all six actions (load-truck, load-airplane, unload-truck, unload-airplane, drive-truck, drive-airplane) and their pre-conditions and post-conditions, as well as predicates needed for the domain.
-$\mathbb{PF}$:
-(define (problem logistics-4-0) (:domain logistics) (:objects apn1 apt2 pos2 apt1 pos1 cit2 cit1 tru2 tru1 obj23 obj22 obj21 obj13 obj12 obj11 ) (:init (package obj11) (package obj12) (package obj13) (package obj21) (package obj22) (package obj23) (truck tru1) (truck tru2) (city cit1) (city cit2) (location pos1) (location apt1) (location pos2) (location apt2) (airport apt1) (airport apt2) (airplane apn1) (at apn1 apt2) (at tru1 pos1) (at obj11 pos1) (at obj12 pos1) (at obj13 pos1) (at tru2 pos2) (at obj21 pos2) (at obj22 pos2) (at obj23 pos2) (in-city pos1 cit1) (in-city apt1 cit1) (in-city pos2 cit2) (in-city apt2 cit2)) (:goal (and (at obj11 apt1) (at obj23 pos1) (at obj13 apt1) (at obj21 pos1))) )
-The $\mathbb{PF}$ contains the objects, initial state and goal state for the problem.
-A.6 Logistics-100 $\mathbb{DD}$ and $\mathbb{PD}$
-The following display example $\mathbb{DD}$ and $\mathbb{PD}$ for all natural settings in the Logistics-100 dataset.
-We can see that the descriptions have all the same components as the $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL, but written in different levels of naturalness.
-For the Heavily Templated $\mathbb{DD}$, all preconditions and post-conditions are written out explicitly and sound similar to PDDL.
-The $\mathbb{PD}$ is similar, in that it lists all the predicates needed for to solve the task.
-Heavily Templated $\mathbb{DD}$:
-I need to move packages between locations.
-Here are the actions I can do Load an package onto a truck at a location (load-truck package truck location) Load an package onto an airplane at a location (load-airplane package airplane location) Unload an package from a truck at a location (unload-truck package truck location) Unload an package from an airplane at a location (unload-airplane package airplane location) Drive a truck from location1 to location2 in a city (drive-truck truck location1 location2 city) Fly an airplane from airport1 to airport2 (fly-airplane airplane airport1 airport2) I have the following restrictions on my actions: To perform load-truck action, the following facts need to be true: o is an package, t is a truck, l is a location, the truck is at the location, the package is at the location.
-Once load-truck action is performed the following facts will be true: the package is in the truck.
-Once load-truck action is performed the following facts will be false: the package is at the location.
-To perform load-airplane action, the following facts need to be true: o is an package, a is an airplane, l is a location, the airplane is at the location, the package is at the location.
-Once load-airplane action is performed the following facts will be true: the package is in the airplane.
-Once load-airplane action is performed the following facts will be false: the package is at the location.
-To perform unload-truck action, the following facts need to be true: o is an package, t is a truck, l is a location, the truck is at the location, the package is in the truck.
-Once unload-truck action is performed the following facts will be true: the package is at the location.
-Once unload-truck action is performed the following facts will be false: the package is in the truck.
-$\mathbb{DD}$ (cont’d)
-To perform unload-airplane action, the following facts need to be true: o is an package, a is a airplane, l is a location, the airplane is at the location, the package is in the airplane.
-Once unload-airplane action is performed the following facts will be true: the package is at the location.
-Once unload-airplane action is performed the following facts will be false: the package is in the airplane.
-To perform drive-truck action, the following need to be true: t is a truck, l1 is a location, l2 is a location, c is a city, the truck is at l1, l1 is in the city, l2 is in the city.
-Once drive-truck action is performed the following facts will be true: the truck is at l2.
-Once drive-truck action is performed the following facts will be false: the truck is at l1.
-To perform fly-airplane action, the following must be true: p is an airplane, a1 is an airport, a2 is an airport, the airplane is at a1.
-Once fly-airplane is performed the following facts will be true: the airplane is at a2.
-Once fly-airplane is performed the following facts will be false: the airplane is at a1.
-Heavily Templated $\mathbb{PD}$:
-As initial conditions, I have that, obj11 is a package, obj12 is a package, obj13 is a package, obj21 is a package, obj22 is a package, obj23 is a package, tru1 is a truck, tru2 is a truck, cit1 is a city, cit2 is a city, pos1 is a location, apt1 is a location, pos2 is a location, apt2 is a location, apt1 is an airport, apt2 is an airport, apn1 is an airplane, apn1 is at apt2, tru1 is at pos1, obj11 is at pos1, obj12 is at pos1, obj13 is at pos1, tru2 is at pos2, obj21 is at pos2, obj22 is at pos2, obj23 is at pos2, pos1 is in cit1, apt1 is in cit1, pos2 is in cit2, and apt2 is in cit2.
-My goal is to have that obj11 is at apt1, obj23 is at pos1, obj13 is at apt1, and obj21 is at pos1.
-For the Moderately Templated data, the $\mathbb{DD}$ and $\mathbb{PD}$ are much more natural than the Heavily Templated data, but all predicates are still listed.
-Moderately Templated $\mathbb{DD}$:
-I need to move packages between locations.
-Here are the actions I can do Load an package onto a truck at a location (load-truck package truck location) Load an package onto an airplane at a location (load-airplane package airplane location) Unload an package from a truck at a location (unload-truck package truck location) Unload an package from an airplane at a location (unload-airplane package airplane location) Drive a truck from location1 to location2 in a city (drive-truck truck location1 location2 city) Fly an airplane from airport1 to airport2 (fly-airplane airplane airport1 airport2) I have the following restrictions on my actions: I can only load a package onto a truck or airplane if both the package and airplane are at the location.
-Once I load the package in the truck or airplane, it is no longer at the location.
-I can only unload a package from a truck or airplane if the truck or airplane is at the location and the package is in the truck or airplane.
-Once I unload the truck or airplane, the object is at the location and no longer in the truck or airplane.
-I can only drive a truck between locations if the truck is at the first location and both the first and second locations are in the same city.
-Once I drive a truck, the truck is in the second city and no longer in the first city.
-I can only fly an airplane between two airports and the airplane is at the first airport.
-Once I fly an airplane, the airplane is at the second airport and no longer at the first airport.
-The $\mathbb{PD}$ for the Moderately Templated data is the same as the $\mathbb{PD}$ for Heavily Templated data.
-Finally for the natural data, we can see that the $\mathbb{DD}$ and $\mathbb{PD}$ give all necessary information to complete the task, but does not sound like PDDL, and does not describe all predicates needed to perform the task.
-Natural $\mathbb{DD}$:
-In the Logistics game, your goal is to transport packages between different locations using trucks and airplanes.
-Here are the actions you can perform: You can load a package onto a truck at a particular location if both the package and the truck are present there.
-Similarly, loading a package onto an airplane requires both the package and the airplane to be at the same location.
-Once a package is loaded onto a truck or airplane, it leaves its original location.
-To unload a package, the truck or airplane must be at the same location where you want to unload, and the package should be inside the vehicle.
-When you unload, the package arrives at the new location and exits the vehicle.
-If you want to drive a truck from one location to another within a city, the truck needs to begin its journey at the starting point, and both locations must be within the city boundaries.
-After driving, the truck will find itself at the destination, leaving the starting point behind.
-As for flying, an airplane can travel between two airports, but it must be ready for takeoff from the initial airport.
-After the flight, the airplane lands at the destination airport, departing from the origin airport in the process.
-Natural $\mathbb{PD}$:
-In this logistics scenario, we begin with several objects and locations.
-Package obj11, obj12, and obj13 are initially at location pos1.
-Similarly, package obj21, obj22, and obj23 start at location pos2.
-We have two trucks: truck tru1 is stationed at pos1, and truck tru2 is at pos2.
-Additionally, we have a single airplane, apn1, which is located at airport apt2.
-Our map comprises two cities: cit1 and cit2.
-City cit1 contains location pos1 and airport apt1, while city cit2 includes location pos2 and airport apt2.
-The ultimate objective is to relocate package obj11 and obj13 to airport apt1 and to move package obj21 and obj23 to location pos1.
-A.7 Barman-100 PDDL
-The following are an example of the groundtruth $\mathbb{DF}$ and $\mathbb{PF}$ for Logistics-100.
-$\mathbb{DF}$:
-(define (domain barman) (:requirements :strips :typing) (:types hand level beverage dispenser container - object ingredient cocktail - beverage shot shaker - container) (:predicates (ontable ?c - container) (holding ?h - hand ?c - container) (handempty ?h - hand) (empty ?c - container) (contains ?c - container ?b - beverage) (clean ?c - container) (used ?c - container ?b - beverage) (dispenses ?d - dispenser ?i - ingredient) (shaker-empty-level ?s - shaker ?l - level) (shaker-level ?s - shaker ?l - level) (next ?l1 ?l2 - level) (unshaked ?s - shaker) (shaked ?s - shaker) (cocktail-part1 ?c - cocktail ?i - ingredient) (cocktail-part2 ?c - cocktail ?i - ingredient)) (:action grasp :parameters (?h - hand ?c - container) :precondition (and (ontable ?c) (handempty ?h)) :effect (and (not (ontable ?c)) (not (handempty ?h)) (holding ?h ?c))) (:action leave :parameters (?h - hand ?c - container) :precondition (holding ?h ?c) :effect (and (not (holding ?h ?c)) (handempty ?h) (ontable ?c))) (:action fill-shot :parameters (?s - shot ?i - ingredient ?h1 ?h2 - hand ?d - dispenser) :precondition (and (holding ?h1 ?s) (handempty ?h2) (dispenses ?d ?i) (empty ?s) (clean ?s)) :effect (and (not (empty ?s)) (contains ?s ?i) (not (clean ?s)) (used ?s ?i)))
-$\mathbb{DF}$ (cont’d)
-(:action refill-shot :parameters (?s - shot ?i - ingredient ?h1 ?h2 - hand ?d - dispenser) :precondition (and (holding ?h1 ?s) (handempty ?h2) (dispenses ?d ?i) (empty ?s) (used ?s ?i)) :effect (and (not (empty ?s)) (contains ?s ?i))) (:action empty-shot :parameters (?h - hand ?p - shot ?b - beverage) :precondition (and (holding ?h ?p) (contains ?p ?b)) :effect (and (not (contains ?p ?b)) (empty ?p))) (:action clean-shot :parameters (?s - shot ?b - beverage ?h1 ?h2 - hand) :precondition (and (holding ?h1 ?s) (handempty ?h2) (empty ?s) (used ?s ?b)) :effect (and (not (used ?s ?b)) (clean ?s))) (:action pour-shot-to-clean-shaker :parameters (?s - shot ?i - ingredient ?d - shaker ?h1 - hand ?l ?l1 - level) :precondition (and (holding ?h1 ?s) (contains ?s ?i) (empty ?d) (clean ?d) (shaker-level ?d ?l) (next ?l ?l1)) :effect (and (not (contains ?s ?i)) (empty ?s) (contains ?d ?i) (not (empty ?d)) (not (clean ?d)) (unshaked ?d) (not (shaker-level ?d ?l)) (shaker-level ?d ?l1))) (:action pour-shot-to-used-shaker :parameters (?s - shot ?i - ingredient ?d - shaker ?h1 - hand ?l ?l1 - level) :precondition (and (holding ?h1 ?s) (contains ?s ?i) (unshaked ?d) (shaker-level ?d ?l) (next ?l ?l1)) :effect (and (not (contains ?s ?i)) (contains ?d ?i) (empty ?s) (not (shaker-level ?d ?l)) (shaker-level ?d ?l1)))
-$\mathbb{DF}$ (cont’d)
-(:action empty-shaker :parameters (?h - hand ?s - shaker ?b - cocktail ?l ?l1 - level) :precondition (and (holding ?h ?s) (contains ?s ?b) (shaked ?s) (shaker-level ?s ?l) (shaker-empty-level ?s ?l1)) :effect (and (not (shaked ?s)) (not (shaker-level ?s ?l)) (shaker-level ?s ?l1) (not (contains ?s ?b)) (empty ?s))) (:action clean-shaker :parameters (?h1 ?h2 - hand ?s - shaker) :precondition (and (holding ?h1 ?s) (handempty ?h2) (empty ?s)) :effect (and (clean ?s))) (:action shake :parameters (?b - cocktail ?d1 ?d2 - ingredient ?s - shaker ?h1 ?h2 - hand) :precondition (and (holding ?h1 ?s) (handempty ?h2) (contains ?s ?d1) (contains ?s ?d2) (cocktail-part1 ?b ?d1) (cocktail-part2 ?b ?d2) (unshaked ?s)) :effect (and (not (unshaked ?s)) (not (contains ?s ?d1)) (not (contains ?s ?d2)) (shaked ?s) (contains ?s ?b))) (:action pour-shaker-to-shot :parameters (?b - beverage ?d - shot ?h - hand ?s - shaker ?l ?l1 - level) :precondition (and (holding ?h ?s) (shaked ?s) (empty ?d) (clean ?d) (contains ?s ?b) (shaker-level ?s ?l) (next ?l1 ?l)) :effect (and (not (clean ?d)) (not (empty ?d)) (contains ?d ?b) (shaker-level ?s ?l1) (not (shaker-level ?s ?l)))) )
-The $\mathbb{DF}$ contains all twelve actions and their pre-conditions and post-conditions, as well as predicates needed for the domain.
-$\mathbb{PF}$:
-(define (problem prob) (:domain barman) (:objects shaker1 - shaker left right - hand shot1 - shot ingredient1 ingredient2 - ingredient cocktail1 - cocktail dispenser1 dispenser2 - dispenser l0 l1 l2 - level ) (:init (ontable shaker1) (ontable shot1) (dispenses dispenser1 ingredient1) (dispenses dispenser2 ingredient2) (clean shaker1) (clean shot1) (empty shaker1) (empty shot1) (handempty left) (handempty right) (shaker-empty-level shaker1 l0) (shaker-level shaker1 l0) (next l0 l1) (next l1 l2) (cocktail-part1 cocktail1 ingredient1) (cocktail-part2 cocktail1 ingredient2) ) (:goal (and (contains shot1 cocktail1) )))
-The $\mathbb{PF}$ contains the objects, initial state and goal state for the problem.
-A.8 Barman-100 $\mathbb{DD}$ and $\mathbb{PD}$
-The following display example $\mathbb{DD}$ and $\mathbb{PD}$ for the heavily templated data for Barman-100.
-We can see that the descriptions have all the same components as the $\mathbb{DF}$ and $\mathbb{PF}$ in PDDL.
-For the Heavily Templated $\mathbb{DD}$, all preconditions and post-conditions are written out explicitly and sound similar to PDDL.
-The $\mathbb{PD}$ is similar, in that it lists all the predicates needed for to solve the task.
-Heavily Templated $\mathbb{DD}$:
-I am creating a cocktail from a set of ingredients.
-Here are the actions I can do Grasp a container (grasp hand container) Leave a container (leave hand container) Fill a shot glass with with an ingredient (fill-shot shot ingredient hand1 hand2 dispenser) Re-fill a shot glass with an ingredient (refill-shot shot ingredient hand1 hand2 dispenser) Empty a shot glass (empty-shot hand shot beverage) Clean a shot glass (clean-shot shot beverage hand hand2) Pour an ingredient from a shot glass to a clean shaker (pour-shot-to-clean-shaker shot ingredient shaker hand level level1) Pour an ingredient from a shot glass to a used shaker (pour-shot-to-used-shaker shot ingredient shaker hand level level1) Empty a shaker (empty-shaker hand shaker cocktail level level1) Clean a shaker (clean-shaker hand1 hand2 shaker) Shake a shaker (shaker cocktil ingredient1 ingredient2 shaker hand1 hand2) Pour a cocktail from a shaker to a shot glass (pour-shaker-to-shot beverage shot hand shaker level level1) I have the following restrictions on my actions: To perform Grasp action, the following facts need to be true: container on table, hand empty.
-Once Grasp action is performed the following facts will be true: hand holding container.
-Once Grasp action is performed the following facts will be false: container on table, hand empty.
-To perform Leave action, the following facts need to be true: hand holding container.
-Once Leave action is performed the following facts will be true: hand empty, container on table.
-Once Leave action is performed the following facts will be false: hand holding container.
-To perform Fill-shot action, the following needs to be true: hand1 holding shot glass, hand2 empty, dispenser dispenses ingredient, empty shot glass, clean shot glass.
-Once Fill-shot action is performed the following will be true: shot glass contains ingredient, shot glass used with ingredient.
-Once Fill-shot action is performed the following will be false: empty shot glass, clean shot glass.
-To perform Refill-shot action, the following needs to be true: hand1 holding shot glass, hand2 empty, dispenser dispenses ingredient, empty shot glass, shot glass used with ingredient.
-Once Refill-shot action is performed the following will be true: shot glass contains ingredient.
-Once Refill action is performed the following will be false: empty shot glass.
-To perform Empty-shot action, the following needs to be true: hand holding shot glass, shot glass contains beverage.
-Once Empty-shot action is performed the following will be true: empty shot glass.
-Once Empty-shot action is performed the following will be false: shot glass contains beverage.
-To perform Clean-shot action, the following needs to be true: hand1 holding shot glass, hand2 empty, empty shot glass, shot glass used with beverage.
-Once Clean-shot action is performed the following will be true: clean shot glass.
-Once Clean-shot action is performed the following will be false: shot glass used with beverage To perform Pour-shot-to-clean-shaker action, the following needs to be true: hand1 holding shot glass, shot glass contains ingredient, empty shaker, clean shaker, shaker-level of shaker is l, next level from l is l1.
-Once Pour-shot-to-clean-shaker action is performed the following will be true: empty shot glass, shaker contains ingredient, unshaked shaker, shaker-level of shaker is l1.
-Once Pour-shot-to-clean-shaker action is performed the following will be false: shot glass contains ingredient, empty shaker, clean shaker, shaker-level of shaker is l.
-To perform Pour-shot-to-used-shaker action, the following needs to be true: hand1 holding shot glass, shot glass contains ingredient, unshaked shaker, shaker-level of shaker is l, next level from l is l1.
-Once Pour-shot-to-used-shaker action is performed the following will be true: shaker contains ingredient, empty shot glass, shaker-level of shaker is l1.
-Once Pour-shot-to-used-shaker action is performed the following will be false: shot glass contains ingredient, shaker-level of shaker is l.
-$\mathbb{DD}$ (cont’d)
-To perform Empty-shaker action, the following needs to be true: hand holding shaker, shaker contains cocktail, shaked shaker, shaker-level of shaker is l, shaker-empty-level of shaker is l1.
-Once Empty-shaker action is performed the following will be true: shaker-level of shaker is l1, empty shaker.
-Once Empty-shaker action is performed the following will be false: shaked shaker, shaker-level of shaker is l, shaker contains cocktail.
-To perform Clean-shaker action, the following needs to be true: hand1 holding shaker, hand2 empty, empty shaker.
-Once Clean-shaker action is performed the following will be true: clean shaker.
-To perform Shake action, the following needs to be true: hand1 holding shaker, empty hand2, shaker contains ingredient1, shaker contains ingredient2, part 1 of cocktail is ingredient1, part 2 of cocktail is ingredient2, unshaked shaker.
-Once Shake action is performed the following will be true: shaked shaker, shaker contains cocktail.
-Once Shake action is performed the following will be false: unshaked shaker, shaker contains ingredient1, shaker contains ingredient2.
-To perform Pour-shaker-to-shot action, the following needs to be true: hand holding shaker, shaked shaker, empty shot glass, clean shot glass, shaker contains cocktail, shaker level of shaker is l, next level from l1 is l.
-Once Pour-shaker-to-shot action is performed the following will be true: shot glass contains cocktail, shot glass used with cocktail, shaker-level of shaker is l1.
-Once Pour-shaker-to-shot action is performed the following will be false: clean shot glass, empty shot glass, shaker-level of shaker is l.
-Heavily Templated $\mathbb{PD}$:
-For this cocktail, I have the following: shaker shaker1, my left hand, my right hand, shot glass shot1, ingredient ingredient1, ingredient ingredient2, dispenser dispenser1, and dispenser dispenser2.
-The shaker has the following levels: l0, l1, and l2.
-I want to make the following cocktails: cocktail1.
-As initial conditions, I have that shaker1 is on the table, shot1 is on the table, dispenser1 dispenses ingredient1, dispenser2 dispenses ingredient2, shaker1 is clean, shot1 is clean, shaker1 is empty, shot1 is empty, handempty left, handempty right, shaker-empty-level shaker1 l0, shaker-level shaker1 l0, next l0 l1, next l1 l2, cocktail-part1 cocktail1 ingredient1, and cocktail-part2 cocktail1 ingredient2.
-My goal is to have that shot1 contains cocktail1.
-Appendix B Prompts
-For the LLM-as-planner, we give all the models the following prompt for BlocksWorld-100:
-Here is a game involving a table with blocks on it.
-{domain_description} {problem_description} Write the plan that would solve this problem.
-These are the available actions: (PICK-UP block): pick up a block from the table (PUT-DOWN block): put down a block on the table (STACK block1 block2): stack block1 onto block2 (UNSTACK block1 block2): unstack block1 from block2 Here is what the output should look like: (PICK-UP A) (STACK A B) (UNSTACK A B) (PUT-DOWN A)
-For MysteryBlocksWorld-100, we use the following prompt:
-Here is a game involving a table with blocks on it.
-{domain_description} {problem_description} Write the plan that would solve this problem.
-These are the available actions: (ATTACK object): attack object (SUCCUMB object): succumb (OVERCOME object1 object2): overcome object1 from object2 (FEAST object1 object2): feast object1 from object2 Here is what the output should look like: (ATTACK A) (OVERCOME A B) (FEAST A B) (SUCCUMB A)
-Whenever possible, we asked the model to return the output in a JSON object for easier parsing.
-Appendix C Detailed Results
-Beyond the visualizations above, we show the detailed results of all models on all simulations of all naturalness levels.
-C.1 Results for BlocksWorld-100
-Metrics
-Natural
-Models Solvability Correctness
-gemma-2-9b-it 3/100 3/100
-gemma-2-9b-itp - 9/100
-gemma-2-27b-it 0/100 0/100
-gemma-2-27b-itp - 11/100
-Llama-3.1-8B 0/100 0/100
-Llama-3.1-8Bp - 1/100
-Llama-3.1-70B 0/100 0/100
-Llama-3.1-70Bp - 13/100
-gpt-3.5-turbo 2/100 1/100
-gpt-4o-mini 19/100 3/100
-gpt-4o-minip - 7/100
-gpt-4o 64/100 60/100
-gpt-4op - 33/100
-o1-preview 91/100 91/100
-o1-previewp - 82/100
-o3-mini 79/100 68/100
-o3-minip - 87/100
-Moderately Templated
-Models Solvability Correctness
-gemma-2-9b-it 0/100 0/100
-gemma-2-9b-itp - 5/100
-gemma-2-27b-it 17/100 10/100
-gemma-2-27b-itp - 3/100
-Llama-3.1-8B 0/100 0/100
-Llama-3.1-8Bp - 1/100
-Llama-3.1-70B 0/100 0/100
-Llama-3.1-70Bp - 10/100
-gpt-3.5-turbo 14/100 4/100
-gpt-4o-mini 9/100 5/100
-gpt-4o-minip - 7/100
-gpt-4o 77/100 67/100
-gpt-4op - 35/100
-o3-mini 82/100 70/100
-o3-minip - 87/100
-Heavily Templated
-Models Solvability Correctness
-gemma-2-9b-it 61/100 10/100
-gemma-2-9b-itp - 5/100
-gemma-2-27b-it 81/100 80/100
-gemma-2-27b-itp - 7/100
-Llama-3.1-8B 0/100 0/100
-Llama-3.1-8Bp - 0/100
-Llama-3.1-70B 0/100 0/100
-Llama-3.1-70Bp - 10/100
-gpt-3.5-turbo 39/100 29/100
-gpt-4o-mini 66/100 59/100
-gpt-4o-minip - 1/100
-gpt-4o 89/100 89/100
-gpt-4op - 29/100
-o3-mini 94/100 94/100
-o3-minip - 96/100
-Table 3: Performance of LLM-as-formalizer and LLM-as-planner (p) all BlocksWorld-100 data.
-Table 3 displays results for all results for BlocksWorld-100.
-C.2 Results for MysteryBlocksWorld-100
-Table 4 displays the results for Heavily Templated MysteryBlocksWorld-100.
-Models Solvability Correctness
-Llama-3.1-8B 0/100 0/100
-Llama-3.1-8Bp - 0/100
-Llama-3.1-70B 0/100 0/100
-Llama-3.1-70Bp - 0/100
-gemma-2-9b-it 100/100 99/100
-gemma-2-9b-itp - 9/100
-gemma-2-27b-it 99/100 98/100
-gemma-2-27b-itp - 0/100
-gpt-3.5-turbo 4/100 0/100
-gpt-4o-mini 36/100 5/100
-gpt-4o-minip - 0/100
-gpt-4o 74/100 70/100
-gpt-4op - 0/100
-o3-mini 95/100 95/100
-o3-minip - 74/100
-Table 4: Performance of LLM-as-formalizer and LLM-as-planner (p) on the Heavily Templated MysteryBlocksWorld-100.
-C.3 Results for Logistics-100
-Table 5 displays results for all naturalness settings for Logistics-100
-Metrics
-Natural
-Models Solvability Correctness
-gemma-2-9b-it 1/100 0/100
-gemma-2-9b-itp - 0/100
-gemma-2-27b-it 1/100 0/100
-gemma-2-27b-itp - 0/100
-Llama-3.1-8B 0/100 0/100
-Llama-3.1-8Bp - 1/100
-Llama-3.1-70B 0/100 0/100
-Llama-3.1-70Bp - 0/100
-gpt-3.5-turbo 1/100 0/100
-gpt-3.5-turbop - 0/100
-gpt-4o-mini 13/100 0/100
-gpt-4o-minip - 0/100
-gpt-4o 20/100 2/100
-gpt-4op - 1/100
-o3-mini 43/100 7/100
-o3-minip - 14/100
-Moderately Templated
-Models Solvability Correctness
-gemma-2-9b-it 0/100 0/100
-gemma-2-9b-itp - 0/100
-gemma-2-27b-it 3/100 0/100
-gemma-2-27b-itp - 0/100
-Llama-3.1-8B 0/100 0/100
-Llama-3.1-8Bp - 0/100
-Llama-3.1-70B 0/100 0/100
-Llama-3.1-70Bp - 0/100
-gpt-3.5-turbo 0/100 0/100
-gpt-3.5-turbop - 0/100
-gpt-4o-mini 29/100 0/100
-gpt-4o-minip - 0/100
-gpt-4o 34/100 13/100
-gpt-4op - 0/100
-o3-mini 51/100 39/100
-o3-minip - 21/100
-Heavily Templated
-Models Solvability Correctness
-gemma-2-9b-it 1/100 0/100
-gemma-2-9b-itp - 0/100
-gemma-2-27b-it 0/100 0/100
-gemma-2-27b-itp - 1/100
-Llama-3.1-8B 0/100 0/100
-Llama-3.1-8Bp - 0/100
-Llama-3.1-70B 0/100 0/100
-Llama-3.1-70Bp - 0/100
-gpt-3.5-turbo 1/100 1/100
-gpt-3.5-turbop - 0/100
-gpt-4o-mini 26/100 0/100
-gpt-4o-minip - 0/100
-gpt-4o 33/100 13/100
-gpt-4op - 0/100
-o3-mini 55/100 47/100
-o3-minip - 21/100
-Table 5: Performance of LLM-as-formalizer and LLM-as-planner (p) all Logistics-100 data.
-C.4 Results for Barman-100
-Figure 5: Performance for Barman-100.
-Table 6 displays results for Barman-100.
-Models Solvability Correctness
-Llama-3.1-8B 0/100 0/100
-Llama-3.1-8Bp - 0/100
-Llama-3.1-70B 0/100 0/100
-Llama-3.1-70Bp - 15/100
-gemma-2-9b-it 0/100 0/100
-gemma-2-9b-itp - 2/100
-gemma-2-27b-it 0/100 0/100
-gemma-2-27b-itp - 0/100
-gpt-3.5-turbo 0/100 0/100
-gpt-3.5-turbop - 4/100
-gpt-4o-mini 0/100 0/100
-gpt-4o-minip - 8/100
-gpt-4o 1/100 0/100
-gpt-4op - 3/100
-o3-mini 1/100 0/100
-o3-minip - 19/100
-Table 6: Performance of LLM-as-formalizer and LLM-as-planner (p) on the Heavily Templated Barman-100.
-Appendix D Sample Model Output
-The following is an example $\mathbb{DF}$ and $\mathbb{PF}$ that Llama-3.1-8B-Instruct gave. We can see that there are syntax errors, as well as semantic errors in the $\mathbb{DF}$ and $\mathbb{PF}$.
-We give the model the following input $\mathbb{DD}$ and $\mathbb{PD}$:
+Listing 17: Example Domain Description and Problem Description given to Llama-3.1-8B-Instruct
+;Domain Description:
 Blocksworld is a puzzle domain where the aim is to move a set of blocks to achieve a specific goal configuration.
-The domain consists of a flat surface, often called the ’table’, and an arbitrary number of blocks.
+The domain consists of a flat surface, often called the table’, and an arbitrary number of blocks.
 Every block can be stacked on top of another block or placed on the table.
 Blocks can be moved one at a time, and only the block at the top of a stack can be moved.
-The main actions available in Blocksworld are: 1.
-**Pick Up (Pick-Up B)**: Picks up block B from the table.
-2.
-**Put Down (Put-Down B)**: Places block B onto the table.
-3.
-**Stack (Stack B1 B2)**: Stacks block B1 on top of block B2.
-4.
-**Unstack (Unstack B1 B2)**: Removes block B1 from the top of block B2.
+The main actions available in Blocksworld are:
+1. **Pick Up (Pick-Up B)**: Picks up block B from the table.
+2. **Put Down (Put-Down B)**: Places block B onto the table.
+3. **Stack (Stack B1 B2)**: Stacks block B1 on top of block B2.
+4. **Unstack (Unstack B1 B2)**: Removes block B1 from the top of block B2.
 The goal is to transform an initial configuration of blocks into a desired target configuration using these actions.
-In this Blocksworld problem, we will consider a simple scenario with 2 blocks: Block1 and Block2.
-The initial state of the blocks is as follows: **Initial State**: - Block1 is on the table.
+;Problem Description:
+In this Blocksworld problem, we will consider a simple scenario with 2 blocks: Block1 and Block2. The initial state of the blocks is as follows:
+**Initial State**:
+- Block1 is on the table.
 - Block2 is on Block1.
-The goal is to achieve the following configuration: **Goal State**: - Block2 is on the table.
+The goal is to achieve the following configuration:
+**Goal State**:
+- Block2 is on the table.
 - Block1 is on the table.
 To solve this problem, you will need to use the available actions to rearrange the blocks from the initial state to the goal state.
-Which returns the following
-Output $\mathbb{DF}$:
-(define (domain blocksworld) (:requirements :strips :typing) (:types block) (:predicates (on-table?b - block) (on?b1 - block?b2 - block) (clear?b - block) ) (:action pick-up :parameters (?b - block) :preconditions (and (on-table?b) (clear?b)) :effects (and (not (on-table?b)) (not (clear?b)) (on-table?b)) ) (:action put-down :parameters (?b - block) :preconditions (and (not (on-table?b)) (clear?b)) :effects (and (on-table?b) (clear?b)) ) (:action stack :parameters (?b1 - block?b2 - block) :preconditions (and (on-table?b1) (clear?b2)) :effects (and (not (on-table?b1)) (not (clear?b2)) (on?b1?b2)) ) (:action unstack :parameters (?b1 - block?b2 - block) :preconditions (and (on?b1?b2) (clear?b1)) :effects (and (on-table?b1) (not (on?b1?b2)) (clear?b2)) ) )
-Output $\mathbb{PF}$:
-(define (problem blocksworld-problem) (:domain blocksworld) (:objects block1 block2 - block) (:init (on-table block1) (on block2 block1) (clear block1) (clear block2) ) (:goal (and (on-table block2) (on-table block1) (clear block1) (clear block2) ) ) )
-Text in red displays errors outputted from the model.
-We can see that in the $\mathbb{DF}$ there are syntax errors (incorrect keyword “preconditions”) as well as semantic errors (incorrect predicates in preconditions and effects).
-For the output $\mathbb{PF}$ there are semantic errors, such as incorrect preconditions (a block cannot be clear and have another block on top of it) in the init section.
-Environment LLM predicts? Natural Descriptions?
-Zuo et al. (2024) fully-observed $\mathbb{PF}$ N
-Zhang et al. (2024a) partially-observed $\mathbb{PF}$ N
-Liu et al. (2023a) fully-observed $\mathbb{PF}$ N
-Xie et al. (2023) fully-observed & partially-observed $\mathbb{PF}$ goal N
-Lyu et al. (2023) fully-observed $\mathbb{PF}$ goal N
-Zhang et al. (2024c) procedural texts $\mathbb{DF}$ action semantics N
-Wong et al. (2023) partially-observed $\mathbb{DF}$ N
-Guan et al. (2023) fully-observed $\mathbb{DF}$ $\mathbb{PF}$ goal N
-Zhu et al. (2024) fully-observed $\mathbb{DF}$ action semantics N
-Tang et al. (2024) partially-observed Python N/A
-Silver et al. (2024) fully-observed plan N
-Valmeekam et al. (2024) fully-observed plan N
-Stechly et al. (2024) fully-observed plan N
-This work fully-observed $\mathbb{DF}$ $\mathbb{PF}$ Y
-Table 7: Comparison with related work.
-Appendix E Related Works Comparison
-Table 7 compares works related to this paper.
-We can see that other works as the LLM to predict either the plan, parts of PDDL files and other languages.
-We can also see that other works have mostly templated natural language descriptions, while this work uses both templated and natural descriptions.
-Appendix F Methodology Comparison
-Both methodologies used in this paper incorporate LLMs into planning.
-There are pros and cons to each methodology.
-When using LLM-as-planner, we have a lightweight solution that returns results very quickly.
-However, due to the lack of reasoning skills in LLMs, they often struggle to create formal plans.
-Meanwhile using LLM-as-formalizer provides better executability and interpretability, though it uses a solver, which may results in getting results slower.
-We believe that for performance reasons that LLM-as-formalizer is the superior methodology.
+Listing 18: Generated Domain File from Llama-3.1-8B-Instruct
+(define (domain blocksworld)
+(:requirements :strips :typing)
+(:types block)
+(:predicates (on-table?b - block) (on?b1 - block?b2 - block) (clear?b - block) )
+(:action pick-up
+:parameters (?b - block)
+:preconditions (and (on-table?b) (clear?b))
+:effects (and (not (on-table?b)) (not (clear?b)) (on-table?b))
+(:action put-down
+:parameters (?b - block)
+:preconditions (and (not (on-table?b)) (clear?b))
+:effects (and (on-table?b) (clear?b))
+(:action stack
+:parameters (?b1 - block?b2 - block)
+:preconditions (and (on-table?b1) (clear?b2))
+:effects (and (not (on-table?b1)) (not (clear?b2)) (on?b1?b2))
+(:action unstack
+:parameters (?b1 - block?b2 - block)
+:preconditions (and (on?b1?b2) (clear?b1))
+:effects (and (on-table?b1) (not (on?b1?b2)) (clear?b2))
+Listing 19: Generated Problem File from Llama-3.1-8B-Instruct
+(define (problem blocksworld-problem)
+(:domain blocksworld)
+(:objects block1 block2 - block)
+(:init
+(on-table block1)
+(on block2 block1)
+(clear block1)
+(clear block2)
+(:goal
+(and (on-table block2)
+(on-table block1)
+(clear block1)
+(clear block2)